### v1

# Overcoming Reward Model Noise in Instruction-Guided Reinforcement Learning

Abstract
Vision-language models (VLMs) have gained traction as auxiliary reward models to provide more informative reward signals in sparse reward environments. However, our work reveals a critical vulnerability of this method: a small amount of noise in the reward signal can severely degrade agent performance. In challenging environments with sparse rewards, we show that reinforcement learning agents using VLM-based reward models without proper noise handling perform worse than agents relying solely on exploration-driven methods. We hypothesize that false positive rewards – where the reward model incorrectly assigns rewards to trajectories that do not fulfill the given instruction – are more detrimental to learning than false negatives. Our analysis confirms this hypothesis, revealing that the widely used cosine similarity metric, when applied to comparing agent trajectories and language instructions, is prone to generating false positive reward signals. To address this, we introduce BiMI (Binary Mutual Information), a novel noise-resilient reward function. Our experiments demonstrate that, BiMI significantly boosts the agent performance, with an average improvement ratio of 44.5% across diverse environments with learned, non-oracle VLMs, thereby making VLM-based reward models practical for real-world applications [^1] Code available at https://anonymous.4open.science/r/Official-Overcoming-Reward-Noise-in-Instruction-Guided-RL-E9B6.
1 Introduction
In real-world reinforcement learning (RL) applications, reward signals are often absent, sparse, or difficult to define explicitly (Clark 2016). To address this, natural language instructions have been adopted as an effective source of reward signals, providing a direct, goal-oriented guidance to agents. This approach, known as instruction-guided reinforcement learning, employs vision-language models (VLMs) to measure the semantic similarity between an agent’s behavior and the provided instructions and reward the agents based on the similarity scores (Kaplan et al. 2017; Goyal et al. 2019, 2020; Du et al. 2023). Specifically, the VLM generates vector embeddings for both the agent’s trajectory history (a sequence of visual frames) and the instructions within a shared embedding space. Notably, prior approaches have consistently adopted cosine similarity to quantify the semantic similarity between vision and language, a metric first demonstrated and proven effective by Frome et al. (2013).
Figure 1: Illustration of false positives in a VLM embedding space. The undesired agent’s trajectory (dashed line) may exhibit high cosine similarity to the instruction (solid line) in the embedding space, as indicated by the proximity of their endpoints in the embedding space. Despite this apparent similarity, these trajectories fundamentally fail to fulfill the intended instruction. Therefore, the rewards guides the agent towards incorrect behaviors. The figure shows three distinct cases of false positives. Please refer to Section 4 for more details.
However, a crucial challenge in VLM-based reward models has been largely overlooked: these models are learned from data rather than being explicitly defined. Consequently, they inherently introduce reward noise due to imperfections in the learning process. This noise manifests in two forms: false positive rewards, where the model incorrectly assigns high rewards to trajectories that do not fulfill the given instructions, and false negative rewards, where the model assigns inadequate rewards to the matched pairs.
Previous research has largely celebrated the potential of VLM reward models under controlled settings, neglecting the pervasive issue of noisy signals. Also, investigations into policy learning failures with learned reward models have typically attributed these failures to false negative rewards and poor training data quality (Du et al. 2023). Our study, however, shifts the focus to the equally critical but less discussed issue of false positive rewards. We argue that (1) false positive rewards are more prevalent and detrimental to learning than false negatives; and (2) using cosine similarity metric is inadequate for capturing the true semantic similarity between trajectories and instructions in RL’s sequential decision-making context, resulting in false positive rewards. As illustrated in Figure 1, this inadequacy stems from two fundamental issues: state entanglement, where the metric fails to account for the impact of actions on state transitions, often assigning high similarity based solely on lexical similarities; and composition insensitivity, where the metric fails to adequately penalize incorrect execution sequences and incomplete task execution.
Recognizing that false positive rewards in VLM-based reward models can significantly hinder performance, we introduce BiMI, a novel reward function designed to mitigate their impact. Specifically, BiMI leverages binary reward signals to directly reduce the occurrence of false positives and incorporates a mutual information term to prevent overfitting to potentially noisy signal sources. Our main contributions are as follows: (1) We provide a detailed analysis of how false positive rewards are generated and its impact on RL, challenging the prevailing assumption that the degeneracy of instruction-guided RL is due to false negative rewards. (2) We introduce BiMI, a noise-robust reward function that significantly improves instruction-guided RL, achieving an average 44.5% performance increase across diverse environments.
2 Related Work
Use of VLMs as Reward Models. VLMs have been pivotal in robotics studies by serving as reward models that guide agents to follow instructions (Wang et al. 2018; Shridhar et al. 2022; Mahmoudieh et al. 2022). These studies primarily focus on leveraging VLMs t overcome the challenge of manual reward design for complex tasks, without explicitly considering the impact of reward noise or its implications for policy convergence rates. VLM-based reward models have also been commonly used to solve sparse reward environments such as Montezuma’s Revenge in Atari games (Kaplan et al. 2017; Goyal et al. 2019). However, their approaches, which are tested on simplified sub-tasks or require extensive prior knowledge of the environment, inadvertently evade the impact of noisy reward signal. Simplified sub-tasks prevent the accumulation of noise over longer horizons, while prior knowledge about the environment allows agents to reduce their dependence on the noisy reward signal.
Some work sidesteps the noisy reward problem by constructing oracle VLMs, which use internal state information from the game engine rather than pixel information to calculate semantic similarity (Du et al. 2023; Wang et al. 2023). Despite these efforts, the issue of noisy rewards remains a significant challenge in VLM-based reward models. For example, Du et al. (2023) reported difficulties in policy learning with a vanilla learned VLM, attributing the failure to false negatives – instances where the VLM fails to recognize matched pairs. They claimed that an average false negative rate of 11% led to unsuccessful training. However, our research challenges this emphasis on false negatives as the primary obstacle in instruction-guided RL. Our findings indicate that false positive rewards are actually more prevalent and detrimental to the learning process. We provide a comprehensive analysis of noise in VLM-based reward models and examine the comparative impact of false positives and false negatives on the learning efficacy of agents.
Reward Signal from Human Preference. Recent work on Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al. 2022) also leverage human preference as a reward signal. However, our work differs in key aspects. Unlike RLHF’s focus on textual outputs, our approach involves evaluating cross-modal similarities between visual and textual data in environments that require long-horizon decision-making and frequent embodied interactions, a domain not typically covered by RLHF.
Mitigating Reward Model Failures. Research by Ghosal et al. (2022) and Fu et al. (2024) has introduced methods to counteract unreliable reward signals from learned VLMs. These strategies involve employing a parallel exploration-based policy alongside the reward-maximizing policy, thereby reducing reliance on potentially misspecified VLM rewards. Our work extends these efforts by providing a comprehensive analysis of noise in VLM-based reward models, with a particular focus on the distinct impacts of false positive and false negative rewards. Furthermore, instead of relying on exploration policies to escape local optima induced by noisy VLM rewards, we directly address the noise within the reward function itself, preventing these pitfalls from arising in the first place.
3 Background
In the context of using instructions as auxiliary guide for RL agents, we frame our task as a MDP defined by a tuple $\langle\mathcal{S},\mathcal{A},\mathcal{T},s_{0},r^{\mathit{e}},\gamma\rangle$, where $\mathcal{S}$ represents a set of states $s\in\mathcal{S}$, $\mathcal{A}$ represents a set of actions $a\in\mathcal{A}$, and $\mathcal{T}(s^{\prime}\mid s,a)$ describes the dynamics of the environment. $s_{0}\in\mathcal{S}$ is the initial state, and $\gamma\in[0,1]$ is the reward discount factor. $r^{\mathit{e}}(s,a,s^{\prime})$ is the environmental reward function. In this work, we focus on a sparse reward setting, where the agent receives a non-zero reward only when reaching goal states $S_{G}\subset\mathcal{S}$, and 0 otherwise, with $|S_{G}|\ll|\mathcal{S}|$. An agent’s trajectory is a sequence of states and actions $\tau=\langle s_{t},a_{t},\ldots,s_{t+T}\rangle$, where $t$ is the initial time step and $T$ is the length of the trajectory. The objective is to learn a policy $\pi$ that maximizes the expected cumulative reward of the trajectories $\tau^{\pi}=\{\tau_{1},...,\tau_{n}\mid\tau_{i}\text{ induced by }\pi\}$.
A sparse reward function only realizes a set of acceptable policies, rather than distinguishing between them more finely (as derived from Abel et al. (2021); see Appendix A.1 for details). Building on this, we define the following:
(1) Policy Universe, $\Pi$: the set of all possible policies for the MDP.
(2) Acceptable Policies, $\Pi_{G}$: the policies in $\Pi_{G}$ are those policies whose start-state values $V^{\pi}(s_{0})$ are within a small $\epsilon$ of the optimal value, while all other policies have strictly lower values.
(3) Unacceptable Policies, $\Pi_{B}$: formally, $\Pi_{B}=\Pi\setminus\Pi_{G}$, which refers to the set of policies that do not meet the near-optimal criteria and are therefore considered not acceptable.
A task instruction $L$ is a sequence of natural language sentences $L=\{l_{1},l_{2},\ldots,l_{n}\}$, where each $l_{k}$ induces a set of desired sub-trajectories $\tau^{l_{k}}$. A complete task execution is represented by a trajectory that concatenates instructions’ corresponding sub-trajectories in order, traversing from the initial state $s_{0}$ to a goal state in $S_{G}$.
An auxiliary VLM-based reward model provides more frequent rewards by evaluating the semantic similarity between the agent’s trajectory and the specified instruction sentence. This can be formally represented as $r^{v}(\tau,l_{k})=p(l_{k}|\tau)$ (see Figure 2), where $r^{v}$ denotes the VLM-based reward function. The use of non-Markovian reward functions in MDP environments has been well-established in the literature, particularly through the work on reward machines (Icarte et al. 2018; Corazza et al. 2022). Such a VLM-based reward function consists of two components: an embedding encoder $\phi(\cdot)$ and a similarity function. The reward is calculated as the cosine similarity between the embeddings, which can be expressed as $\frac{\phi(\tau)\cdot\phi(l_{k})}{\|\phi(\tau)\|\|\phi(l_{k})\|}$.
Figure 2: Illustration of reward signals from a vision-language reward model
Reduction of Convergence Rate. To understand the effects of reward noise on policy learning, it’s crucial to analyze its impact on the convergence rate. In sparse reward settings, the convergence rate – defined as the number of training iterations needed to learn an $\epsilon$-optimal policy – can be significantly affected by false positive rewards that are incorrectly given for policies that do not lead to the goal states.
To illustrate this, we analyze the convergence rate by measuring the probability of learning acceptable policies ($P(\pi\in\Pi_{G})$) over each update iteration. The following theorem demonstrates how false positive rewards impact the convergence rate:
Theorem 3.1 (Convergence rate reduction).
In Actor-Critic algorithm, gradient ascent on $Q(s,a)\pi_{\theta_{i}}(a|s)$ pushes the next updated policy $\pi_{\theta_{i+1}}$ in the direction provided by the $Q$ value function. In the direction of maximizing $P(\pi\in\Pi_{G})$, the gradient of $Q(s,a)\pi(a|s)$ can be expressed as follows:
$\nabla_{\theta}Q_{\phi}(s,a)\pi_{\theta}(a|s)=\textit{const}\cdot\mathbb{E}[G^{\pi\in\Pi_{B}}]\nabla_{\theta}\pi_{\theta}(a|s)$
$\qquad+(\mathbb{E}[G^{\pi\in\Pi_{G}}]-\mathbb{E}[G^{\pi\in\Pi_{B}}])P(\pi\in\Pi_{G})\nabla_{\theta}\pi_{\theta}(a|s)$
where $\Pi_{G}$ is the set of acceptable policies; $\Pi_{B}$ is the set of unacceptable policies that follow instructions but fail to reach the goal; $\mathbb{E}[G^{\pi}]$ is expectation of the cumulative rewards by executing policy $\pi$ in one episode.
The proof is in Appendix A.2. In this context, false positive rewards correspond to rewards assigned to policies within the set $\Pi_{B}$, which represents undesirable policies. Therefore, $\mathbb{E}[G^{\pi\in\Pi_{B}}]$ can be interpreted as the expected cumulative false positive rewards. This theorem highlights how the presence of $\mathbb{E}[G^{\pi\in\Pi_{B}}]$ reduce the convergence rate by: (1) introducing deviation in the policy update directions (the first term), and (2) reducing the gradient of the desired policy update (the second term). Both effects can significantly hinder the convergence process.
The issue of noisy rewards is not merely a theoretical concern; it has practical implications for the effectiveness of instruction-guided RL in real-world applications. In the next section, we will investigate how cosine similarity metric used in VLM-based reward models inevitably introduce false positive rewards. Additionally, we will empirically evaluate the impact of both false positives and false negatives on the learning efficacy of agents in sparse reward environments.
4 Approximation Error of Cosine Similarity
Cosine similarity is insufficient for evaluating the alignment between trajectories and instructions. In this section, we will discuss two fundamental issues associated with cosine similarity scores in RL contexts: state entanglement and composition insensitivity. The former issue, state entanglement, refers to the metric’s inability to recognize trajectories $\{\tau_{1},...,\tau_{n}\mid\tau_{i}\text{ induced by }\pi\in\Pi_{B}\}$ that, while being cosine similar to the target instruction in the embedding space, fail to reach the goal states in $S_{G}$. The latter issue refers to the metric’s tendency to reward trajectories that execute the sub-tasks (corresponding to individual instruction sentences $l_{k}$) in an incorrect order. This is problematic in RL tasks where the specific sequence of executing sub-tasks, as defined by the order of instructions in $L=\{l_{1},l_{2},\ldots,l_{n}\}$, is crucial for reaching the goal states.
4.1 The Issue of State Entanglement
State entanglement refers to the issue where the cosine similarity metric erroneously pays more attention to lexical-level similarity while lacking comprehension of the underlying state transitions. Consequently, rewards are given to trajectory-instruction pairs that are cosine similar in embedding space but in fact result in distinct state transitions. For instance, consider the significant contrast between “take the red pill” and “take the blue pill”. Despite their lexical similarity, they lead to vastly different states. However, the cosine similarity metric may represent them as similar due to the shared action “take” and shared object “pill”, disregarding the critical difference in state outcomes. Understanding state transitions is crucial in sequential decision-making scenarios. Otherwise, rewards may be given to trajectories that lead to unintended states, resulting in the learning of unacceptable policies.
4.2 The Issue of Composition Insensitivity
Composition insensitivity in cosine similarity metrics gives rise to two issues: (1) rewarding incomplete task execution – cosine similarity may incorrectly reward partial task completion, as even incomplete trajectories can receive high similarity score in the embedding space. For instance, in a task to “pick up a tool, then fix a machine,” the model might prematurely reward the agent for merely picking up the tool, neglecting the crucial repair action. This can lead to suboptimal learning, with the agent potentially getting stuck in easier, incomplete subtasks rather than fully accomplishing the task. (2) insensitivity to the ordering of execution – cosine similarity often fails to adequately penalize incorrect execution sequences. In a safety protocol requiring an agent to “turn off the machinery, then perform maintenance,” the metric might assign high rewards based merely on the presence of relevant actions, disregarding their order. While some large language models have become sensitive to the order of tokens (Su et al. 2024), the compact visual and sentence embeddings from multimodal models remains largely insensitive to sequential information (Pham et al. 2021). This limitation can lead to potentially dangerous outcomes in scenarios where execution sequence is critical, such as in safety protocols or complex manufacturing processes.
Figure 1 illustrates various scenarios where high similarity scores are erroneously assigned to false positive cases. To empirically demonstrate the issue, Section 4.3 presents experiments on false positive rewards and their impact on agent learning in sparse reward environments.
Figure 3: Learned VLM models differentiate between matched and not-matched pairs, but struggle with O.O.D. cases. They incorrectly assign high scores to manipulated pairs, which should be low as the trajectories in the manipulated pairs do not fulfill the instruction
Figure 4: The false positive vs. false negative oracle model. The false positive model get a more severe drop in the final training score.
4.3 Experiments on Reward Noise Impact
Our experiments test the following hypothesis: (H1) The two issues of state entanglement and composition insensitivity exist; (H2) false positive rewards are prevalent during training; (H3) similarity-based reward models lacking noise handling mechanisms underperform against strong intrinsic reward models in sparse reward environments; (H4) false negatives may not be as harmful as false positives.
Setup. We evaluate these hypotheses through various challenging sparse-reward environments: (1) Crafter, an open-ended 2D Minecraft environment (Hafner 2021); (2) Montezuma’s Revenge, a classic hard adventure game in Atari (Bellemare et al. 2013); and (3) Minigrid ‘Go To Seq’, a hard task involving long-horizon navigation and object interactions (Chevalier-Boisvert et al. 2018). Two cosine similarity-based reward models were tested: (1) Pixl2R by Goyal et al. (2020), which uses only the current video frame to determine if the goal state specified in the instruction has been reached; and (2) ELLM-, a variant of ELLM by Du et al. (2023). Unlike ELLM, which queries instructions from LLMs in real-time, ELLM- directly uses preset expert instructions and compares them with the transition differences of the agent’s trajectory. The VLM backbones used are: (1) CLIP (Radford et al. 2021), pretrained by image-text pairs; and (2) X-CLIP (Ma et al. 2022), pretrained by video-text pairs. To ensure high-quality finetuning data, we used internal information from the game engine to annotate expert trajectories from expert agents. To demonstrate how noisy reward signals hinder learning, we selected a strong intrinsic reward model DEIR (Zhang et al. 2021) for comparison. It provides auxiliary rewards based on observation novelty to encourage exploration. See Appendix A.3 for detailed implementation of the experiments.
Evaluation Metric. We used the score metric adapted from the Crafter benchmark (Hafner 2021) for performance evaluation, as it better captures an agent’s consistent performance across multiple subtasks. This metric follows the intuitive ‘higher is better’ principle. Unlike maximum total rewards metric, which fail to capture consistent performance in sparse reward settings, the score metric provides a more reliable measure of learning progress.
Reward Noise Issue. To investigate H1, we evaluated the models’ sensitivity by examining how cosine similarity scores change for manipulated trajectory-instruction pairs. The state entanglement test involved reversing trajectories and negating instructions (i.e., “do not do $l_{k}$”). The composition insensitivity test examined concatenated pairs of trajectory-instruction data. For example, given $(\tau_{1},l_{1})$ and $(\tau_{2},l_{2})$, we create a concatenated pair $(\tau_{1}+\tau_{2},l_{1}+l_{2})$. We then test two types of manipulations – (1) swapping the order within one modality: e.g., $(\tau_{2}+\tau_{1},l_{1}+l_{2})$; and (2) truncating one modality: e.g., $(\tau_{1},l_{1}+l_{2})$. Overall, we categorized the trajectory-instruction pairs for evaluation into three types: (1) matched pairs – these are positive examples where the trajectory correctly corresponds to the given instruction. (2) not-matched pairs – these are negative examples where the trajectory does not correspond to the given instruction. (3) manipulated pairs: derived from matched pairs by altering either the trajectory or the instruction. Ideally, manipulated pairs should receive low similarity scores as essentially the trajectory are not fulfilling the instruction. However, our results reveal that the reward model assigns high scores to these manipulated pairs, particularly in the “do not do $l_{k}$” case. This finding highlights the noise issue in cosine similarity-based reward models (see Figure 4). It’s worth noting that the poor performance in the negation case aligns with broader challenges in natural language processing. Recent studies (Hossain et al. 2022; Truong et al. 2023) have highlighted that negation is central to language understanding but is not properly captured by modern language models. This limitation extends to VLMs and significantly impacts their ability to provide accurate rewards in complex scenarios and instructions.
Figure 5: The heatmap shows the cumulative rewards received at various locations, with larger circle sizes indicating higher rewards. The later figures shows the offsets between the state where rewards are given and the actual goal-reaching state. Agents are getting both issues of false positives and false negatives during training
Figure 6: The ratio of false positive rewards is significantly reduced after applying BiMI
Prevalence of False Positives. To address H2, we analyzed reward distribution heatmap from VLM-based reward models during training. The heatmap revealed a concerning trend: RL agents engage in reward hacking, receiving rewards across vast areas of the environment rather than just at goal states. For instance, in Montezuma environment where the goal is to grab the key and escape the room, we observed that agents received rewards even for falling off cliffs, which undoubtedly contribute to the detrimental $\mathbb{E}[G^{\pi\in\Pi_{B}}]$ specified in Theorem 3.1. For environments without fixed camera views, we calculated the step offset between the current rewarded state and the actual goal state. A positive offset indicates that a reward was given before reaching the goal state, signifying a false positive reward, while a negative offset suggests that the agent reached the goal state but the reward model failed to recognize it, indicating a false negative reward (see Figure 6). Interestingly, we observed a large amount of negative step offsets in Minigrid environments. We attribute this to Minigrid’s abstract shape-based visual representations, which fall outside the VLM’s pretraining distribution.
Impact on Learning. We trained agents using learned VLM reward models and compared their learning efficacy against intrinsic reward models. As shown in Table 1, our results confirmed H3: instruction-guided RL agents using learned VLM reward models without noise handling consistently underperform compared to DEIR, the intrinsic reward-based RL agent. To investigate the impact of false negatives versus false positives (H4), we designed an oracle Pixl2R model with two variants: a false negative model and a false positive model. The false negative model only rewards the agent for reaching subgoal states described in the instruction, with a probability of xthat some rewarding states in the map are removed. In contrast, the false positive model rewards the agent for reaching every subgoal, but also introduces a 0.1 one-off reward for certain locations, covering xof the map. The results indicate that false negatives may be less detrimental to agent performance than false positives (see Figure 4). This performance difference can be explained through our theoretical framework:
Table 1: Score metric across environments (equivalent to total rewards, higher is better). $\star$ denotes baseline intrinsic reward model. VLM reward models without noise handling underperformed. All models are based on PPO.
Models Type Monte. Minigrid Crafter % vs. DEIR
PPO Pure $0.151$ $24.993$ $16.863$ $-28\%$
DEIR $\star$ Intrinsic $0.174$ $55.556$ $19.758$
Pixl2R VLM $0.142$ $12.422$ $9.409$ $\mathbf{-49\%}$
ELLM- VLM $0.150$ $19.406$ $10.826$ $-41\%$
Pixl2R + DEIR VLM + intr. $0.176$ $17.372$ $10.440$ $-38\%$
ELLM- + DEIR VLM + intr. $0.178$ $30.985$ $11.857$ $-27\%$
  1. 1.
    False negative models maintain $\mathbb{E}[G^{\pi\in\Pi_{G}}]-\mathbb{E}[G^{\pi\in\Pi_{B}}]$, ensuring steady gradient ascent towards maximizing $P(\pi\in\Pi_{G})$.
  2. 2.
    By decreasing $\mathbb{E}[G^{\pi\in\Pi_{B}}]$, false negative models minimize deviations from the target direction, leading to more stable learning.
In contrast, scenarios with high $\mathbb{E}[G^{\pi\in\Pi_{B}}]$ (i.e., the false positive case) significantly reduce the gradient ascent rate and also introduce deviations from the target direction. The findings thus challenged the prevailing assumption that false negatives are the primary issue in the degeneracy of instruction-guided RL. See Appendix A.4 for more experimental details.
5 Solution to the Reward Noise Issue
5.1 Preferring Binary Signal
Our experiments have shown that the issue of false negatives may be less detrimental to learning than false positives. Therefore, we propose a reward function that provides one-time binary rewards only when the similarity between the agent’s current trajectory and the instruction exceeds a high confidence threshold. This method effectively reduces the likelihood of rewarding unacceptable trajectories. To achieve this, we introduce a thresholding mechanism using a calibration set of true positive trajectory-instruction pairs. This threshold, denoted as $\hat{q}$, is set to the empirical quantile of cosine similarity scores at the significance level $1-\alpha$. Pairs whose similarity scores fall below this threshold $\hat{q}$ receive no reward. Conversely, pairs exceeding $\hat{q}$ receive a one-time $+1$ reward, i.e., $r^{v}_{\textsc{Bi}}(\tau,l_{k})=\mathbf{1}_{\{p(l_{k}\mid\tau)\geq\hat{q}\}}$. This approach statistically guarantees a high probability (at least $1-\alpha$) that true positive pairs are recognized while minimizing frequency of false positives errors (Sadinle et al. 2019). See Appendix A.5 for detailed threshold calculation.
5.2 Mutual Information Maximization
Intuitively, when we observe rewards coming from a particular signal source too frequently, we tend to downplay the significance of that signal to avoid over-reliance. This intuition is effectively captured by incorporating a mutual information maximization term into the reward function. Specifically, the updated reward function $r^{v}_{\textsc{MI}}(\tau,l_{k})$ measures the mutual information between the agent’s trajectory and the instruction. Mathematically, it can be expressed as:
$r^{v}_{\textsc{MI}}(\tau,l_{k})$ $=I(l_{k};\tau)=D_{KL}(p(l_{k},\tau)\mid\mid p(l_{k})p(\tau))$
$=\mathbb{E}_{\tau\sim\pi_{\theta},l_{k}\sim L}[\log p(l_{k}\mid\tau)-\log p(l_{k})]$ (1)
where $p(l_{k}\mid\tau)$ comes from the similarity score provided by VLMs, and $p(l_{k})$ is the frequency of the instruction $l_{k}$ being fulfilled by the agent’s trajectory. Therefore, the second term in the equation serves as a regularization term that downplays the significance of the reward signal when it is too frequent. For instance, if a VLM frequently detects that the agent’s actions are fulfilling the “climbing the ladder” instruction, even when the agent is performing unrelated tasks, the reward signal for this instruction will be downplayed. $p(l_{k})$ is calculated as follows:
$p(l_{k})=\mathbb{E}_{\tau\sim\pi_{\theta_{-1}}}\left[\frac{\sum_{t=1}^{T_{\tau}}\mathbf{1}_{\{p(l_{k}\mid\tau_{t})\geq\hat{q}\}}}{T_{\tau}}\right]$ (2)
Here, $\tau_{t}$ is the agent’s trajectory up to time $t$, and if the VLM deems the trajectory as fulfilling the instruction (i.e., $p(l_{k}\mid\tau_{t})\geq\hat{q}$), we increment the count. Dividing the count by the total trajectory length $T_{\tau}$ gives the empirical frequency of the instruction being fulfilled. The subscript $\theta_{-1}$ in $\pi_{\theta_{-1}}$ indicates that the trajectories are sourced from rollouts in the previous policy iteration, acknowledging the impracticality of real-time computation of $p(l_{k})$ during an ongoing episode.
To enhance the stability of the training process, we adopt a linearized version of the mutual information maximization approach, as proposed by Li et al. (2023). Overall, BiMI, the proposed reward function that enhances the noise resilience of VLM-based reward models, can be expressed as follows:
$r^{v}_{\textsc{BiMI}}(\tau,l_{k})=\max(\mathbf{1}_{\{p(l_{k}\mid\tau)\geq\hat{q}\}}-p(l_{k}),0)$ (3)
It’s important to note that the BiMI approach primarily mitigates false positives (FP) rather than false negatives (FN). By implementing a high confidence threshold and a binary reward signal, we effectively reduce the likelihood of rewarding incorrect trajectories, thus addressing the FP issue. While this conservative approach may potentially increase FNs, as some correct trajectories might fall below the threshold, our experiments indicate that this trade-off is beneficial. We found that false positives are more detrimental to learning in instruction-guided RL tasks, justifying our preference for a method that prioritizes FP reduction.
In the following sections, we will demonstrate how BiMI significantly improves the effectiveness of VLM-based reward models in instruction-guided RL tasks.
6 Experiments
Figure 7: Besides the improvements of the score performance of agents across different environments with the BiMI reward function, it also collaborates well with intrinsic rewards. Combining both can lead to significant performance improvements
Figure 8: Ablation on the components of BiMI reward function. The binary reward (Bi) alone led to a 36.5% improvement compared to original models. Excluding Crafter, Mutual Information (MI) provided a 23% further improvement over Bi alone
We evaluate the BiMI reward function using baseline agents (Pixl2R and ELLM-) and their BiMI-enhanced counterparts, while also exploring potential synergies with intrinsic reward models. We use the score metric for evaluation.
Table 2: Model score across various environments. $\star$ is the baseline agents with a learned VLM-based reward model to compare with. BiMI significantly improves performance in Montezuma and Minigrid, while showing mixed results in Crafter due to task-specific characteristics
Methods Montezuma % vs. $\star$ Minigrid % vs. $\star$ Crafter % vs. $\star$
Pixl2R $\star$ $0.142\pm 0.003$ $12.422\pm 2.439$ $9.409\pm 0.022$
Pixl2R + Bi $0.137\pm 0.009$ $-3.5\%$ $31.236\pm 2.040$ $151\%$ $10.735\pm 0.784$ $14\%$
Pixl2R + BiMI $0.162\pm 0.022$ $14\%$ $37.507\pm 7.832$ $199\%$ $7.951\pm 0.351$ $-15\%$
Pixl2R + DEIR $0.176\pm 0.009$ $23\%$ $17.372\pm 0.514$ $39\%$ $10.440\pm 1.015$ $10\%$
Pixl2R + BiMI + DEIR $0.267\pm 0.016$ $88\%$ $57.759\pm 2.157$ $364\%$ $11.014\pm 0.190$ $17\%$
ELLM- $\star$ $0.150\pm 0.004$ $19.406\pm 10.067$ $10.826\pm 1.017$
ELLM- + Bi $0.151\pm 0.016$ $0.6\%$ $29.788\pm 1.290$ $53\%$ $11.175\pm 0.601$ $3.2\%$
ELLM- + BiMI $0.156\pm 0.014$ $4.0\%$ $33.683\pm 3.990$ $74\%$ $9.425\pm 0.267$ $-12\%$
ELLM + DEIR $0.178\pm 0.029$ $20\%$ $30.985\pm 3.507$ $59\%$ $11.857\pm 1.152$ $9.5\%$
ELLM- + BiMI + DEIR $0.279\pm 0.078$ $86\%$ $56.281\pm 6.193$ $190\%$ $13.170\pm 0.393$ $22\%$
Figure 9: BiMI reward showed faster and higher success rates on difficult tasks in Montezuma
6.1 Montezuma’s Revenge
In Montezuma, Pixl2R$+$BiMI demonstrated an 14% performance increase compared to the original models (see Table 2), which is slightly below our expectations. We attribute this result to BiMI’s intentional strategy of providing less frequent discrete rewards. While this strategy effectively reduces false positives, it does not substantially mitigate the inherent reward sparsity issue in Montezuma. However, we discovered a remarkable synergy between BiMI and intrinsic reward models. While previous models showed no significant improvements with DEIR (the intrinsic reward model) alone, combining BiMI and DEIR led to a 65% performance gain. The gap in collaboration effectiveness can be attributed to two factors. In the previous setup, the consistent presence of false positive rewards misled agents towards unacceptable behaviors and hindered further exploration. Now, BiMI’s less frequent but more meaningful rewards provide anchor points for the agent’s learning. Meanwhile, DEIR’s intrinsic rewards fill the gaps between these anchor points, encouraging the agent to explore efficiently in the interim.
As illustrated in Figure 6, BiMI rewards are now concentrated on key locations. A significant improvement is the minimal rewards given for falling off cliffs, which was a common source of false positives in the original model. Figure 9 demonstrates higher success rate grabbing the key in the first room, one of the most difficult task in Montezuma, highlighting the effectiveness of the proposed reward function and its collaboration with intrinsic reward models in guiding agents to solve difficult sparse-reward tasks.
6.2 Minigrid
ELLM-$+$BiMI achieved a remarkable 74% improvement in performance compared to the original models.This substantial gain is particularly noteworthy given the unique challenges presented by Minigrid. The abstract, shape-based visuals in Minigrid diverge drastically from the natural images used in VLMs’ pretraining, preventing the models from effectively utilizing their prior pretraining knowledge. Consequently, VLMs struggled to accurately assess similarities between Minigrid’s abstract visuals and textual instructions, resulting in highly noisy reward signals. The significant improvement demonstrated by BiMI underscores its effectiveness in handling noisy signals, directly addressing our primary research challenge. This capability is crucial for deploying instruction-guided agents in real-world, unfamiliar scenarios, where visual inputs often deviate from the VLMs’ training distribution, leading to noisy reward signals.
6.3 Crafter
We observed an intriguing pattern of results. The Bi component alone led to 14% and 3.2% improvement in performance over the original models. However, contrary to our observations in other environments, the addition of the MI component actually decreased this improvement. This unexpected outcome can be attributed to the unique nature of Crafter task, where agents must repeatedly achieve the same subtasks (e.g., drinking water) for survival. The MI component, designed to discourage over-reliance on frequently occurring signals, inadvertently penalized the necessary repetition of survival-critical actions. Furthermore, note that instruction-guided RL agents, regardless of the reward model employed, were unable to outperform pure RL agents in Crafter. This discrepancy is due to the open-world nature of Crafter, which requires dynamic strategies and real-time decision-making that our testing setups did not fully capture. Despite these challenges, it is noteworthy that Bi alone still managed to improve performance over vanilla VLM-based reward models, suggesting that reducing false positives is still beneficial across all testing environments. The combination of BiMI with DEIR (the intrinsic reward model) also showed promising results, indicating a productive balance between exploration (driven by DEIR) and exploitation (guided by BiMI instruction reward).
6.4 Ablation Study and Overall Performance
Our ablation study highlights the distinct contributions of the binary reward (Bi) and Mutual Information (MI) components within the BiMI framework. The binary reward mechanism alone accounted for a substantial 36.5% improvement in performance. When excluding the results from Crafter, MI component further contributes a 23% improvement over the binary reward alone. The overall improvements were substantial. BiMI led to a 67% improvement for Pixl2R and a 22% improvement for ELLM-. These results are illustrated in the line plots in Figure 8 and Figure 8 and summarized in Table 2. See Appendix A.6 for more details.
7 Conclusion
Our research reveals two key findings in instruction-guided reinforcement learning using VLM-based reward models: (1) false positive rewards, rather than false negatives, are more detrimental to policy learning; and (2) our proposed BiMI reward function significantly mitigates the impact of noise in reward models. Our work addresses a crucial gap in prior work, paving the way for more reliable and effective instruction-guided RL systems in practical settings.
Limitation and Future Work. Our study primarily focused on linear sequences of language instructions, excluding more complex cases. Future research should investigate conditional and ambiguous instructions, which likely introduce additional challenges for VLM-based reward models. We also did not explore finetuning the VLM during agent training, a useful strategy as discussed by Fu et al. (2024).
Appendix A Technical Appendix
Continued from the main text of Overcoming Reward Model Noise in Instruction-Guided Reinforcement Learning, the technical appendix consist of the following:
  • §A.1 Property of Sparse Reward Function, which discuss why sparse reward function results in the categorization of policies into a set of acceptable policies and a set of unacceptable policies but not others.
  • §A.2 Proof of the Reduction of Convergence Rate, which is referred in Section 3.
  • §A.3 Implementation Details of the Experiments, which specify the implementation details for both the first and second stage experiments in Section 4.3 and Section 6 respectively.
  • §A.4 Additional Details of the Experiments of False Positive Rewards, which is referred in Section 4.3.
  • §A.5 Pseudo-code for Empirical Quantile Calculation for Binary Signal Threshold, which is referred in Section 5.1.
  • §A.6 Additional Details of the Experiments of BiMI Reward Function, which is referred in Section 6.
  • §A.7 Cosine Similarity Score Evaluation for Different Environments, are details of the cosine similarity score evaluation for different environments, which is referred in Section 4.3.
  • §A.8 Line Plots over Average Reward for Different Environments and Ablation Study, which provides performance comparison of different models using average reward metric, the conventional metric used in RL, though they are less informative than the score metric used in the main text when dealing with sparse reward environments.
A.1 Sparse Reward Function and Range-SOAP
Formally, Abel et al. (2021) defined that:
Theorem A.1 (Abel et al. (2021)).
A reward function realizes a Range Set of Acceptable Policies (Range-SOAP) $\Pi_{G}$ when there exists an $\epsilon\geq 0$ such that every $\pi_{g}\in\Pi_{G}$ is $\epsilon$-optimal in start-state value, $V^{*}(s_{0})-V^{\pi_{g}}(s_{0})\leq\epsilon$, while all other policies are worse.
When the reward signal is sparse, the agent only receives a reward upon reaching some goal states, and the reward function does not provide any feedback during the intermediate steps. We argue that the sparse reward function realizes a Range-SOAP, leading to the categorization of policies into acceptable and unacceptable policies.
Proposition A.2.
Sparse reward function “realizes” Range-SOAP (i.e., Range Set of Acceptable Policies).
Justification:
A sparse reward function, where the agent only receives a reward upon completing the task, cannot prefer policies that lead to shorter task completion times. This is because either the agent completes the goal very quickly or slowly, they will receive nearly the same amount of cumulative rewards, and the reward function will not show a strong preference.
Since the sparse reward function does not induce a strict partial ordering on the policies, we say this reward function cannot realize a Partial Ordering (PO) task. Specifically, a PO on policies is a generalization of a Set of Acceptable Policies (SOAP) task. In a PO, the agent specifies a partial ordering on the policy space, where some policies are identified as “great”, some as “good”, and some as “bad” to strictly avoid, while remaining indifferent to the rest.
Therefore, the sparse reward function can realize a Set of Acceptable Policies (SOAP), where there is a set of policies that are all considered “good” or near-optimal, while all other policies are worse.
Furthermore, the sparse reward function will lead to a Range-SOAP, rather than an Equal-SOAP. Specifically, Equal-SOAP is a SOAP where all the acceptable policies are equally optimal in start-state value. This is because the good policies in the SOAP may differ slightly in their start-state values, as some may reach multiple goal states in the environment and thereby receiving different cumulative rewards. Therefore, the sparse reward function will realize a Range-SOAP, where there is a range of acceptable policies that are all near-optimal in start-state value.
A.2 Proof of the Reduction of Convergence Rate
Specifically, the update rule of Actor-Critic algorithm is:
  • Critic:
    $\phi\leftarrow\phi-\alpha_{\phi}\nabla_{\phi}(\delta)^{2}$ (4)
    where
    $\delta=\mathbb{E}_{\pi_{\theta}}[G^{\pi_{\theta}}-Q_{\phi}(s,a)]$ is the Monte Carlo (MC) estimation error, $Q_{\phi}(s,a)$ is the $Q$-value function which measures the expected discounted cumulative reward given the state $s$ and action $a$, and $\pi_{\theta}$ is the policy.
    $G^{\pi_{\theta}}$ is the rollout cumulative rewards from the trajectory $\tau^{\pi_{\theta}}$ generated from $\pi_{\theta}$.
  • Actor:
    $\theta\leftarrow\theta+\alpha_{\theta}\frac{Q_{\phi}(s,a)\nabla_{\theta}\pi_{\theta}(a|s)}{\pi_{\theta}(a|s)}$ (5)
We need to make the following assumptions to simplify the theoretical analysis:
  1. 1.
    We use $\Pi_{G}$ to represent the set of acceptable policies and use $\Pi_{B}$ as an arbitrary set of “not acceptable” policies that are learned to follow instruction but fail to reach subgoals, i.e., $\Pi_{B}\cap\Pi_{G}=\varnothing$.
  2. 2.
    Assume the policy class parameterized by $\theta$ should be expressive enough to capture optimal or near-optimal policies, and the policy is initialized randomly from uniform distribution, i.e., $\pi_{\theta_{0}}\sim\mathcal{U}(\Pi)$. Meanwhile, the Q-value is initialized as $Q_{\phi_{0}}(s,a)=0,\forall_{s\in S}\forall_{a\in A}$.
  3. 3.
    We assume that $|\Pi_{G}|$ and $|\Pi_{B}|$ is a fixed number predefined by the task environment while $\mathbb{E}[G^{\pi}|\pi\in\Pi_{B}]$ is non-zero as false positive rewards are unavoidable in real-world VLMs.
Since the update rule for $Q$-value is a gradient descent on $\|\mathbb{E}[G^{\pi_{\theta}}-Q_{\phi}(s,a)]\|^{2}$ and also we have that $\{\Pi_{B},\Pi_{G},\Pi\setminus(\Pi_{B}\cup\Pi_{G})\}$ is a countable partition of the policy universe $\Pi$, the updated $Q$-value will approach as follows:
$\quad Q_{\phi}(s,a)\rightarrow\mathbb{E}[G^{\pi_{\theta}}]$
$=P(\pi_{\theta}\in\Pi_{B})\cdot\mathbb{E}[G^{\pi_{\theta}}|\pi_{\theta}\in\Pi_{B}]$
$\qquad+P(\pi_{\theta}\in\Pi_{G})\cdot\mathbb{E}[G^{\pi_{\theta}}|\pi_{\theta}\in\Pi_{G}]$
$\qquad+P(\pi_{\theta}\in\Pi\setminus(\Pi_{B}\cup\Pi_{G}))$
$\qquad\cdot\mathbb{E}[G^{\pi_{\theta}}|\pi_{\theta}\in\Pi\setminus(\Pi_{B}\cup\Pi_{G})]$
$=P(\pi_{\theta}\in\Pi_{B})\cdot\mathbb{E}[G^{\pi_{\theta}}|\pi_{\theta}\in\Pi_{B}]$
$\qquad+P(\pi_{\theta}\in\Pi_{G})\cdot\mathbb{E}[G^{\pi_{\theta}}|\pi_{\theta}\in\Pi_{G}]$ (6)
Given that the update rule for policy $\pi$ is the gradient ascent on $Q_{\phi}(s,a)\pi_{\theta}(a|s)$, we have the following:
$\quad\nabla_{\theta}Q_{\phi}(s,a)\pi_{\theta}(a|s)$
$=(\nabla_{\theta}P(\pi_{\theta_{old}}\in\Pi_{B})\cdot\mathbb{E}[G^{\pi_{\theta_{old}}}|\pi_{\theta_{old}}\in\Pi_{B}]\pi_{\theta}(a|s))$
$\qquad+(\nabla_{\theta}P(\pi_{\theta_{old}}\in\Pi_{G})$
$\qquad\cdot\mathbb{E}[G^{\pi_{\theta_{old}}}|\pi_{\theta_{old}}\in\Pi_{G}]\pi_{\theta}(a|s))$
$=P(\Pi_{B})\cdot\mathbb{E}[G^{\Pi_{B}}]\nabla_{\theta}\pi_{\theta}(a|s)$
$\qquad+P(\pi_{G})\cdot\mathbb{E}[G^{\pi_{G}}]\nabla_{\theta}\pi_{\theta}(a|s)$
$=(1-P(\pi_{G})-P(\pi_{\theta}\in\Pi\setminus(\Pi_{B}\cup\Pi_{G})))$
$\qquad\cdot\mathbb{E}[G^{\Pi_{B}}]\nabla_{\theta}\pi_{\theta}(a|s)+P(\pi_{G})\cdot\mathbb{E}[G^{\pi_{G}}]\nabla_{\theta}\pi_{\theta}(a|s)$
$=\textit{const}\cdot\mathbb{E}[G^{\Pi_{B}}]\nabla_{\theta}\pi_{\theta}(a|s)$
$\qquad+(\mathbb{E}[G^{\pi_{G}}]-\mathbb{E}[G^{\Pi_{B}}])P(\pi_{G})\nabla_{\theta}\pi_{\theta}(a|s)$ (7)
Justification of the assumptions.
  • Regarding the non-zero probability of recovering the optimal policy at initialization, it is standard in theoretical analyses to assume a uniform distribution of a random variable at initialization (see Agarwal et al. (2021)). This assumption does not contradict the conclusion about the convergence rate deterioration in Theorem 3.1.
  • In reference to the realizability condition implied by Assumption 2, the expressiveness of the policy class parameterized by $\theta$ is an underlying assumption for deep learning models, supported by the The Universal Approximation Theorem (Hornik et al. 1989).
Observation. Since the goal of the learning agent is to maximize $P(\pi\in\Pi_{G})$ (i.e., to converge to an acceptable policy), we can see that the second term provides the target direction with rate $(\mathbb{E}[G^{\pi\in\Pi_{G}}]-\mathbb{E}[G^{\pi\in\Pi_{B}}])$. Therefore, the ascent rate will decrease when the cumulative rewards from trajectories that cannot reach the goal state (i.e., the false positive rewards) gets higher. Moreover, the first term $\textit{const}\cdot\mathbb{E}[G^{\Pi_{B}}]\nabla_{\theta}\pi_{\theta}(a|s)$ can be regarded as the deviation of target direction. It shows that the level of deviation is also positively proportional to the magnitude of rewards coming from failed trajectories. This theorem follows the intuition that the presence of false positive rewards can slow down the convergence rate of the learning agent.
A.3 Implementation Details of the Experiments
Environment Details
We describe each testing environment used in our experiments. More details introduction can be found in on the official project homepage of each benchmark (Chevalier-Boisvert et al. 2018; Bellemare et al. 2013; Hafner 2021).
  • Crafter features randomly generated 2D worlds where the player needs to forage for food and water, find shelter to sleep, defend against monsters, collect materials, and build tools. The original Crafter environment does not have a clear goal trajectory or instructions; agents are aimed at surviving as long as possible and exploring the environment to unlock new crafting recipes. We modified the environment to include a preset linear sequence of instructions to guide the agent to mine diamond. However, this instruction was found to hinder the agent’s performance. The nature of the task requires dynamic strategies and real-time decision-making, but the fixed instructions limited the agent. For example, the instruction did not account for what to do when the agent is attacked by zombies.
  • Montezuma’s Revenge is a classic adventure platform game where the player must navigate through a series of rooms to collect treasures and keys. The game is known for its sparse rewards and challenging exploration requirements. We manually annotate 97 instructions for the agent to follow, guiding it to conquer the game. The instructions were designed to guide the agent through the game’s key challenges, such as avoiding enemies, collecting keys, and unlocking doors.
  • Minigrid ‘Go to seq’ Task: We use the ‘Go to seq’ task in the Minigrid environment, where the agent must navigate through a sequence of rooms and touch target objects in the correct order. This is a sparse reward task where the agent receives a reward of 1 only upon completing the entire sequence correctly. During the training phase, we randomly generate 50 different tasks, each with a room size of 5, 3 rows, and 3 columns. Each task features a unique room layout and target object sequence. The instruction complexity is set to 3, meaning there are at least 3 target objects to interact with in a specific order.
Figure 10: Illustration of the Montezuma’s Revenge task. The agent must navigate through a series of rooms to collect treasures and keys.
Figure 11: Illustration of the Minigrid ‘Go to seq’ task. The agent must navigate through a sequence of rooms and touch target objects in the correct order.
Figure 12: Illustration of the Crafter task. The agent must survive as long as possible and explore for new crafting recipes.
Instruction-Guide Procedure Details
The VLM-based reward model will have a pointer to the sequence of the instruction sentence, starting at the first sentence. For original models Pixl2R and ELLM-, we follow the setting in there original work where for each instruction sentence, the reward model will have a maximum cap of rewards (2.0) it can assign to the agent in one episode. When the cap is reached, the reward model will move its pointer to the next instruction sentence. For the BiMI reward model, the reward model will move its pointer to the next instruction sentence when the binary signal is triggered.
Finetuning VLM-based Reward Models
In contrast to previous work on Instruction-guided RL where they rely on hand-crafted oracle multimodal reward models, we use actual pretrained VLMs to generate reward signals. 2 VLM backbone models are used in our experiments: 1) CLIP (Radford et al. 2021), pretrained by image-text pairs; and (2) X-CLIP (Ma et al. 2022), pretrained by video-text pairs. In particular, Pixl2R uses CLIP because it only uses the single latest frame as input. In contrast, ELLM- takes a slice of trajectory (i.e., multiple frames) as input, and thus uses either X-CLIP or CLIP with additional RNN encoder as the reward model.
Due to the cartoonish and abstract visuals of the testing environments, we need to further fine-tune the VLMs to adapt to this new visual domain. We use well-trained expert agents based on Moon et al. (2023) to generate expert trajectories for the Crafter environments and annotate them with instructions using internal information from the game engine. For Minigrid environments, we use classical search-based planning robots to generate expert trajectories and annotate them with the corresponding task instructions. For Montezuma’s Revenge, we manually annotate the expert trajectories.
For Minigrid and Crafter, we have 80,000 training pairs, while for Montezuma’s Revenge, we have around 300 training pairs. These training data are of high quality, as we have made every effort to avoid false positive rewards due to poor training data quality. To enhance our models’ robustness, we also employed contrastive learning techniques during VLM training, utilizing similar manipulated data as hard negatives. However, despite the fine-tuning process, false positive rewards remain unavoidable.
Figure 13: Example of training data for the Montezuma environment.
We used the threshold $\hat{q}$ introduced in Section 5.1 to make binary classification on the testing pairs to evaluate the performance of the fine-tuned VLM-based reward models. We found that VLM models had difficulty achieving high accuracy on Minigrid environment, which is likely due to the too abstract and cartoonish nature of the environment, causing the VLMs to struggle to learn the visual-textual correspondence. We also found that X-CLIP did not perform better than CLIP in our experiments. We hypothesize that the cartoonish nature of the testing environments may have caused the X-CLIP model to struggle to learn the visual-textual correspondence. Thus, we used CLIP as the backbone model throughout our following experiments. The performance of the fine-tuned VLM-based reward models is shown in Table 3. Even when the precision score reaches 0.98, indicating that only 2% of the rewards are false positives in the validation set, the agent can still significantly underperform in the testing environments. The core issue is that in out-of-distribution (O.O.D.) testing environments, false positive rewards are prevalent and inevitable. Therefore, it is crucial to design a reward function that is robust to reward noise.
Table 3: Performance of fine-tuned VLM reward model on the testing dataset using the 90th percentile empirical quantile as threshold
Environment Precision Accuracy F1 Score Recall Model
Crafter 0.9847 0.9466 0.8538 0.9702 CLIP ELLM-
Crafter 0.9799 0.9028 0.7618 0.9842 CLIP Pixl2R
Crafter 0.2095 0.2514 0.2868 0.9657 XCLIP ELLM-
Minigrid 0.7260 0.9200 0.7849 0.9763 CLIP ELLM-
Minigrid 0.6992 0.9086 0.7592 0.9616 CLIP Pixl2R
Minigrid 0.1716 0.2310 0.2642 0.9704 XCLIP ELLM-
Montezuma 0.8838 0.9638 0.8825 0.9478 CLIP ELLM-
Montezuma 0.8343 0.9108 0.7652 0.9842 CLIP Pixl2R
Montezuma 0.8044 0.9259 0.8045 0.9657 XCLIP ELLM-
Hyperparameters for Instruction-Guided RL Agents
In the experiments, all methods are implemented based on PPO with same model architecture. The Minigrid and Crafter environments use the same training hyperparameters as the Achievement Distillation paper (Moon et al. 2023). For Montezuma’s Revenge, we found that the performance of the agent was sensitive to the gamma and GAE lambda parameters. To improve the performance of agents in Montezuma’s Revenge, we took two additional steps: (1) normalizing the observation inputs when computing the rewards, and (2) not normalizing the advantage during the GAE calculation. The hyperparameters are shown in the following tables.
Table 4: Model Parameters
Parameter Value
model_cls “PPORNNModel”
hidsize 1024
gru_layers 1
impala_kwargs
- chans [64, 128, 128]
- outsize 256
- nblock 2
- post_pool_groups 1
- init_norm_kwargs
- batch_norm false
- group_norm_groups 1
dense_init_norm_kwargs
- layer_norm true
Table 5: Crafter and Minigrid RL Parameters
Parameter Value
gamma 0.95
gae_lambda 0.65
algorithm_cls “PPOAlgorithm”
algorithm_kwargs
- ppo_nepoch 3
- ppo_nbatch 8
- clip_param 0.2
- vf_loss_coef 0.5
- ent_coef 0.01
- lr 3.0e-4
- max_grad_norm 0.5
- aux_freq 8
- aux_nepoch 6
- pi_dist_coef 1.0
- vf_dist_coef 1.0
Table 6: Montezuma RL Training Parameters
Parameter Value
gamma 0.99
gae_lambda 0.95
int_rew_type “rnd”
pre_obs_norm_steps 50
algorithm_cls “PPOAlgorithm”
algorithm_kwargs
- update_proportion 0.25
- ppo_nepoch 3
- ppo_batch_size 256
- clip_param 0.1
- vf_loss_coef 0.5
- ent_coef 0.001
- lr 1.0e-4
A.4 Additional Details of the Experiments of False Positive Rewards
Evaluation Metric Details In our experiments, we used a score metric adapted from the Crafter paper to evaluate agent performance across different environments. This score metric aggregates success rates for individual subtasks using a geometric mean. This metric was chosen over the maximum total rewards metric for several reasons:
  1. 1.
    Consistency in Sparse Reward Settings: Sparse reward environments often pose significant challenges for reinforcement learning agents. An agent might occasionally achieve high rewards by chance in one rollout but fail to replicate this success consistently in subsequent rollouts. This variability can lead to misleading evaluations if only the maximum total rewards are considered. The Score metric, by measuring the success rate of achieving each subgoal, provides a more stable and consistent measure of an agent’s performance.
  2. 2.
    Capturing Learning Stability: The Score metric evaluates the agent’s ability to consistently reproduce successful behaviors across multiple episodes. This is crucial in sparse reward settings, where the agent’s performance can fluctuate significantly. By focusing on the success rates of individual subtasks, the Score metric offers a more granular and reliable assessment of the agent’s learning progress and stability.
  3. 3.
    Crafter Benchmark Standard: The Crafter benchmark, which introduces the Score metric, is a well-regarded standard.
Crafter codebase provides score metric calculation by default. For Minigrid and Montezuma environments, we use the internal information from the game engine to detect whether the subtasks are completed, thus facilitating the calculation of the score metric.
Details of Manipulated Trajectory-Instruction Pairs to Evaluate Robustness
We evaluated the models’ sensitivity by examining how cosine similarity scores change for manipulated trajectory-instruction pairs. These manipulations were designed to test the robustness of the models against various types of noise. Here’s a detailed breakdown of the manipulations:
  1. 1.
    Trajectory Reversal: We inverted the sequence of frames within each trajectory (i.e., frames = frame[::-1]) to assess the model’s ability to detect reversed state transitions. This manipulation tests whether the model can distinguish between forward and backward progression in the state transition.
  2. 2.
    Instruction Negation: We modified the original instructions by adding negation (e.g., changing “do $l_{k}$” to “do not do $l_{k}$”). This tests the model’s sensitivity to semantic changes in the instruction that fundamentally alter the goal.
  3. 3.
    Instruction Rephrasing: We rephrase the original instructions while maintaining their core meaning. This evaluates the model’s robustness to linguistic variations and its ability to capture the essential semantic content of instructions.
  4. 4.
    Concatenation and Order Swapping: Given two trajectory-instruction pairs $(\tau_{1},l_{1})$ and $(\tau_{2},l_{2})$, we created concatenated pairs and then swapped the order in one modality. For example:
    • Original concatenation: $(\tau_{1}+\tau_{2},l_{1}+l_{2})$
    • Swapped trajectory: $(\tau_{2}+\tau_{1},l_{1}+l_{2})$
    • Swapped instruction: $(\tau_{1}+\tau_{2},l_{2}+l_{1})$
    This tests the model’s sensitivity to the order of components in multi-step tasks.
  5. 5.
    Concatenation with Partial Content: We concatenated pairs but truncated one modality. For instance:
    • Truncated trajectory: $(\tau_{1},l_{1}+l_{2})$
    • Truncated instruction: $(\tau_{1}+\tau_{2},l_{1})$
    This assesses the model’s ability to detect partial mismatches in longer sequences.
A.5 Pseudo-code for Empirical Quantile Calculation for Binary Signal Threshold
Using empirical quantile as threshold guarantees a high probability (at least $1-\alpha$) that the true positive pairs are recognized while minimizing the average number of mistakes predicting false positives (Sadinle et al. 2019):
Algorithm 1 Calculate Empirical Quantile ($\hat{q}$)
Calibration set $\{\tau,l\}_{n}$, where $l$is the instruction sentence, $\tau$is the corresponding trajectory, and $n$is the number of samples; Significance level $\alpha$; VLM model reward model $v$Obtain the similarity-based score
[…]
2:$q_{\text{level}}\leftarrow\frac{\lceil(n-1)\times(1-\alpha)\rceil}{n}$LCommentCompute the empirical quantile
[…]
A.6 Details of the Experiments of BiMI Reward Function
We set confidence level for empirical quantile calculation to be $0.91-\alpha=$ 0.9. We adhered to the standard requirement of limiting the training budget to 1 million frames (Hafner 2021). This constraint poses a significant challenge, particularly in sparse reward settings, as it demands that agents both explore efficiently and exploit their knowledge effectively within this limited budget.
To achieve the 1 million frame budget, we used the following configuration:
  • nproc: 8 (Number of processes used for parallel environments)
  • nstep: 512 (Length of the rollout stored in the buffer)
  • nepoch: 250 (Number of epochs to train the RL policy)
This configuration results in approximately 1 million steps: 250 epochs $\times$ 512 steps $\times$ 8 processes $=$ 1,024,000 frames.
In the case of Montezuma’s Revenge, we found that the 1 million frame limit used in Crafter was insufficient due to the game’s complexity and sparse reward structure. To address this, we extended the training budget to 8 million frames. It’s important to note that even with this increased frame count, agents were still unable to fully solve the task. As Zhang et al. (2021) pointed out, about 1 billion frames are required to truly master Montezuma’s Revenge. This vast difference in required training time (8 million vs 1 billion frames) underscores the exceptional difficulty of Montezuma’s Revenge as a sparse reward task.
The implementation details for the BiMI reward function are consistent with those outlined in the first stage of experiments.
A.7 Cosine Similarity Score Evaluation for Different Environments
Figure 14: Cosine similarity scores over matched, not-matched and manipulated trajectory-instruction pairs in Montezuma’s Revenge.
Figure 15: Cosine similarity scores over matched, not-matched and manipulated trajectory-instruction pairs in Minigrid.
Figure 16: Cosine similarity scores over matched, not-matched and manipulated trajectory-instruction pairs in Crafter.
A.8 Line Plots over Average Reward for Different Environments and Ablation Study
Figure 17: The average reward of agents across different environments. The BiMI reward function significantly improves the average reward of agents in all environments
Figure 18: Ablation on the components of BiMI reward function and their impact on the average reward of agents.

### v4

@@
-Overcoming Reward Model Noise in Instruction-Guided Reinforcement Learning
+The Dark Side of Rich Rewards: Understanding and Mitigating Noise in VLM Rewards
 Abstract
-Vision-language models (VLMs) have gained traction as auxiliary reward models to provide more informative reward signals in sparse reward environments.
-However, our work reveals a critical vulnerability of this method: a small amount of noise in the reward signal can severely degrade agent performance.
-In challenging environments with sparse rewards, we show that reinforcement learning agents using VLM-based reward models without proper noise handling perform worse than agents relying solely on exploration-driven methods.
-We hypothesize that false positive rewards – where the reward model incorrectly assigns rewards to trajectories that do not fulfill the given instruction – are more detrimental to learning than false negatives.
-Our analysis confirms this hypothesis, revealing that the widely used cosine similarity metric, when applied to comparing agent trajectories and language instructions, is prone to generating false positive reward signals.
-To address this, we introduce BiMI (Binary Mutual Information), a novel noise-resilient reward function.
-Our experiments demonstrate that, BiMI significantly boosts the agent performance, with an average improvement ratio of 44.5% across diverse environments with learned, non-oracle VLMs, thereby making VLM-based reward models practical for real-world applications [^1] Code available at https://anonymous.4open.science/r/Official-Overcoming-Reward-Noise-in-Instruction-Guided-RL-E9B6.
+While Vision-Language Models (VLMs) are increasingly used to generate reward signals for training embodied agents to follow instructions, our research reveals that agents guided by VLM rewards often underperform compared to those employing only intrinsic (exploration-driven) rewards, contradicting expectations set by recent work.
+We hypothesize that false positive rewards – instances where unintended trajectories are incorrectly rewarded – are more detrimental than false negatives.
+We confirmed this hypothesis, revealing that the widely used cosine similarity metric is prone to false positive estimates.
+To address this, we introduce BiMI (Binary Mutual Information), a novel reward function designed to mitigate noise.
+BiMI significantly enhances learning efficiency across diverse and challenging embodied navigation environments.
+Our findings offer a nuanced understanding of how different types of reward noise impact agent learning and highlight the importance of addressing multimodal reward signal noise when training embodied agents [^1] Code available at https://shorturl.at/VsH70.
+The Dark Side of Rich Rewards:
+Understanding and Mitigating Noise in VLM Rewards
+Sukai Huang1, Shu-Wei Liu2, Nir Lipovetzky1 and Trevor Cohn1thanks: Now at Google DeepMind 1School of Computing and Information Systems, The University of Melbourne, Australia 2 Max Planck Institute for the Physics of Complex Systems, Nöthnitzer Strasse 38, 01187 Dresden, Germany sukaih@student.unimelb.edu.au, sliu@pks.mpg.de, {nir.lipovetzky, trevor.cohn}@unimelb.edu.au
 1 Introduction
-In real-world reinforcement learning (RL) applications, reward signals are often absent, sparse, or difficult to define explicitly (Clark 2016).
-To address this, natural language instructions have been adopted as an effective source of reward signals, providing a direct, goal-oriented guidance to agents.
-This approach, known as instruction-guided reinforcement learning, employs vision-language models (VLMs) to measure the semantic similarity between an agent’s behavior and the provided instructions and reward the agents based on the similarity scores (Kaplan et al.
-2017; Goyal et al.
-2019, 2020; Du et al.
-2023).
-Specifically, the VLM generates vector embeddings for both the agent’s trajectory history (a sequence of visual frames) and the instructions within a shared embedding space.
-Notably, prior approaches have consistently adopted cosine similarity to quantify the semantic similarity between vision and language, a metric first demonstrated and proven effective by Frome et al.
-(2013).
-Figure 1: Illustration of false positives in a VLM embedding space.
-The undesired agent’s trajectory (dashed line) may exhibit high cosine similarity to the instruction (solid line) in the embedding space, as indicated by the proximity of their endpoints in the embedding space.
-Despite this apparent similarity, these trajectories fundamentally fail to fulfill the intended instruction.
-Therefore, the rewards guides the agent towards incorrect behaviors.
-The figure shows three distinct cases of false positives.
-Please refer to Section 4 for more details.
-However, a crucial challenge in VLM-based reward models has been largely overlooked: these models are learned from data rather than being explicitly defined.
-Consequently, they inherently introduce reward noise due to imperfections in the learning process.
-This noise manifests in two forms: false positive rewards, where the model incorrectly assigns high rewards to trajectories that do not fulfill the given instructions, and false negative rewards, where the model assigns inadequate rewards to the matched pairs.
-Previous research has largely celebrated the potential of VLM reward models under controlled settings, neglecting the pervasive issue of noisy signals.
-Also, investigations into policy learning failures with learned reward models have typically attributed these failures to false negative rewards and poor training data quality (Du et al.
-2023).
-Our study, however, shifts the focus to the equally critical but less discussed issue of false positive rewards.
-We argue that (1) false positive rewards are more prevalent and detrimental to learning than false negatives; and (2) using cosine similarity metric is inadequate for capturing the true semantic similarity between trajectories and instructions in RL’s sequential decision-making context, resulting in false positive rewards.
-As illustrated in Figure 1, this inadequacy stems from two fundamental issues: state entanglement, where the metric fails to account for the impact of actions on state transitions, often assigning high similarity based solely on lexical similarities; and composition insensitivity, where the metric fails to adequately penalize incorrect execution sequences and incomplete task execution.
-Recognizing that false positive rewards in VLM-based reward models can significantly hinder performance, we introduce BiMI, a novel reward function designed to mitigate their impact.
-Specifically, BiMI leverages binary reward signals to directly reduce the occurrence of false positives and incorporates a mutual information term to prevent overfitting to potentially noisy signal sources.
-Our main contributions are as follows: (1) We provide a detailed analysis of how false positive rewards are generated and its impact on RL, challenging the prevailing assumption that the degeneracy of instruction-guided RL is due to false negative rewards.
-(2) We introduce BiMI, a noise-robust reward function that significantly improves instruction-guided RL, achieving an average 44.5% performance increase across diverse environments.
+Natural language instructions are increasingly recognized as a valuable source of reward signals for guiding embodied agents to learn complex tasks.
+In particular, a growing trend in embodied agent learning involves using vision-language models (VLMs) for reward modeling.
+This approach measures the semantic similarity – often quantified by cosine similarity – between the embedding representations of an agent’s behaviors (i.e., past trajectories) and the provided instructions, all within the same embedding space (Kaplan et al., 2017; Goyal et al., 2019, 2020; Du et al., 2023).
+Figure 1: Schematic diagram of false positives in a VLM embedding space.
+The unintended agent’s trajectory (dashed line) may exhibit high cosine similarity to the instruction (solid line) in the embedding space, as indicated by the proximity of their endpoints in the embedding space.
+Despite this apparent similarity, the unintended trajectory fails to fulfill the instruction.
+This illustrates how approximation errors can lead to false positive rewards.
+Refer to Section 5 for more details.
+However, we observed that embodied reinforcement learning (RL) agents trained with VLM rewards, while effective in simplified settings, often struggled with tasks involving complex dynamics and longer action horizons.
+This is evident in several recent works – for instance, Goyal et al.
+(2019) reported the effective use of VLM rewards in Montezuma’s Revenge, a notoriously challenging Atari game.
+However, we observed that this success was confined to individual sub-tasks and the agent struggled when attempting to scale up to the full game.
+Similarly, Du et al.
+(2023) demonstrated impressive performance of VLM rewards in guiding agents within a 2D survival game.
+However, their study was conducted in a modified environment with a reduced observation and action space using internal game state information and manually defined macro actions.
+Consequently, when tested in the original, unmodified environment, their agent’s performance did not exceed that of agents using only intrinsic (exploration-driven) rewards.
+The consistent underperformance of VLM rewards, particularly their unexpected failure to outperform intrinsic rewards, raised concerns about their reliability.
+This discrepancy, where VLM rewards underperformed contrary to their perceived potential, prompted us to investigate the underlying causes of this performance gap.
+Our findings indicate that noisy reward estimates in VLMs are a key factor contributing to poor learning efficacy.
+Specifically, our analysis centered around two classes of noise: false positives, which involve rewarding unintended trajectories, and false negatives, which occur when correct trajectories are not rewarded.
+We posit that false positive rewards (see Figure 1) are not only more prevalent but potentially more detrimental to the learning process than false negatives, a hypothesis supported by our empirical and theoretical findings.
+Among various sources of reward noise, our study particularly investigates the approximation errors within the commonly used cosine similarity metric.
+We examine how these errors generate false positive rewards, which in turn hinder learning.
+To this end, we propose a novel reward function, BiMI (Binary Mutual Information Reward).
+It uses binary reward signals to directly reduce the occurrence of false positives and incorporates a mutual information term to prevent overfitting to noisy signal sources.
+Our experiments demonstrate that BiMI significantly improves the learning efficacy of instruction following agents trained by VLM rewards across various challenging environments.
 2 Related Work
-Use of VLMs as Reward Models.
-VLMs have been pivotal in robotics studies by serving as reward models that guide agents to follow instructions (Wang et al.
-2018; Shridhar et al.
-2022; Mahmoudieh et al.
-2022).
-These studies primarily focus on leveraging VLMs t overcome the challenge of manual reward design for complex tasks, without explicitly considering the impact of reward noise or its implications for policy convergence rates.
-VLM-based reward models have also been commonly used to solve sparse reward environments such as Montezuma’s Revenge in Atari games (Kaplan et al.
-2017; Goyal et al.
-2019).
-However, their approaches, which are tested on simplified sub-tasks or require extensive prior knowledge of the environment, inadvertently evade the impact of noisy reward signal.
-Simplified sub-tasks prevent the accumulation of noise over longer horizons, while prior knowledge about the environment allows agents to reduce their dependence on the noisy reward signal.
-Some work sidesteps the noisy reward problem by constructing oracle VLMs, which use internal state information from the game engine rather than pixel information to calculate semantic similarity (Du et al.
-2023; Wang et al.
-2023).
-Despite these efforts, the issue of noisy rewards remains a significant challenge in VLM-based reward models.
-For example, Du et al.
-(2023) reported difficulties in policy learning with a vanilla learned VLM, attributing the failure to false negatives – instances where the VLM fails to recognize matched pairs.
-They claimed that an average false negative rate of 11% led to unsuccessful training.
-However, our research challenges this emphasis on false negatives as the primary obstacle in instruction-guided RL.
-Our findings indicate that false positive rewards are actually more prevalent and detrimental to the learning process.
-We provide a comprehensive analysis of noise in VLM-based reward models and examine the comparative impact of false positives and false negatives on the learning efficacy of agents.
-Reward Signal from Human Preference.
-Recent work on Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al.
-2022) also leverage human preference as a reward signal.
-However, our work differs in key aspects.
-Unlike RLHF’s focus on textual outputs, our approach involves evaluating cross-modal similarities between visual and textual data in environments that require long-horizon decision-making and frequent embodied interactions, a domain not typically covered by RLHF.
+Using VLMs as Reward Models.
+VLMs have been pivotal in robotics studies, serving as reward models that guide agents to follow instructions (Wang et al., 2018; Shridhar et al., 2022; Mahmoudieh et al., 2022).
+While most research primarily focuses on leveraging VLMs to overcome the challenge of manual reward design for complex tasks (Clark, 2016), the impact of reward noise and its implications for policy convergence rates are often overlooked.
+As mentioned in Section 1, some work sidesteps the noisy reward problem by accessing internal state information from the game engine as well as providing predefined action macros (Du et al., 2023; Wang et al., 2023), thereby preventing the accumulation of noise over longer horizon.
+However, understanding the impact of reward noise from VLMs is crucial for developing a reliable language interface for embodied learning agents in real-world applications.
 Mitigating Reward Model Failures.
@@
 (2022) and Fu et al.
-(2024) has introduced methods to counteract unreliable reward signals from learned VLMs.
+(2024) has introduced methods to counteract unreliable rewards from learned VLMs.
 These strategies involve employing a parallel exploration-based policy alongside the reward-maximizing policy, thereby reducing reliance on potentially misspecified VLM rewards.
-Our work extends these efforts by providing a comprehensive analysis of noise in VLM-based reward models, with a particular focus on the distinct impacts of false positive and false negative rewards.
-Furthermore, instead of relying on exploration policies to escape local optima induced by noisy VLM rewards, we directly address the noise within the reward function itself, preventing these pitfalls from arising in the first place.
-3 Background
-In the context of using instructions as auxiliary guide for RL agents, we frame our task as a MDP defined by a tuple $\langle\mathcal{S},\mathcal{A},\mathcal{T},s_{0},r^{\mathit{e}},\gamma\rangle$, where $\mathcal{S}$ represents a set of states $s\in\mathcal{S}$, $\mathcal{A}$ represents a set of actions $a\in\mathcal{A}$, and $\mathcal{T}(s^{\prime}\mid s,a)$ describes the dynamics of the environment.
-$s_{0}\in\mathcal{S}$ is the initial state, and $\gamma\in[0,1]$ is the reward discount factor.
-$r^{\mathit{e}}(s,a,s^{\prime})$ is the environmental reward function.
-In this work, we focus on a sparse reward setting, where the agent receives a non-zero reward only when reaching goal states $S_{G}\subset\mathcal{S}$, and 0 otherwise, with $|S_{G}|\ll|\mathcal{S}|$.
-An agent’s trajectory is a sequence of states and actions $\tau=\langle s_{t},a_{t},\ldots,s_{t+T}\rangle$, where $t$ is the initial time step and $T$ is the length of the trajectory.
-The objective is to learn a policy $\pi$ that maximizes the expected cumulative reward of the trajectories $\tau^{\pi}=\{\tau_{1},...,\tau_{n}\mid\tau_{i}\text{ induced by }\pi\}$.
-A sparse reward function only realizes a set of acceptable policies, rather than distinguishing between them more finely (as derived from Abel et al.
-(2021); see Appendix A.1 for details).
-Building on this, we define the following:
-(1) Policy Universe, $\Pi$: the set of all possible policies for the MDP.
-(2) Acceptable Policies, $\Pi_{G}$: the policies in $\Pi_{G}$ are those policies whose start-state values $V^{\pi}(s_{0})$ are within a small $\epsilon$ of the optimal value, while all other policies have strictly lower values.
-(3) Unacceptable Policies, $\Pi_{B}$: formally, $\Pi_{B}=\Pi\setminus\Pi_{G}$, which refers to the set of policies that do not meet the near-optimal criteria and are therefore considered not acceptable.
-A task instruction $L$ is a sequence of natural language sentences $L=\{l_{1},l_{2},\ldots,l_{n}\}$, where each $l_{k}$ induces a set of desired sub-trajectories $\tau^{l_{k}}$.
-A complete task execution is represented by a trajectory that concatenates instructions’ corresponding sub-trajectories in order, traversing from the initial state $s_{0}$ to a goal state in $S_{G}$.
-An auxiliary VLM-based reward model provides more frequent rewards by evaluating the semantic similarity between the agent’s trajectory and the specified instruction sentence.
-This can be formally represented as $r^{v}(\tau,l_{k})=p(l_{k}|\tau)$ (see Figure 2), where $r^{v}$ denotes the VLM-based reward function.
-The use of non-Markovian reward functions in MDP environments has been well-established in the literature, particularly through the work on reward machines (Icarte et al.
-2018; Corazza et al.
-2022).
-Such a VLM-based reward function consists of two components: an embedding encoder $\phi(\cdot)$ and a similarity function.
-The reward is calculated as the cosine similarity between the embeddings, which can be expressed as $\frac{\phi(\tau)\cdot\phi(l_{k})}{\|\phi(\tau)\|\|\phi(l_{k})\|}$.
-Figure 2: Illustration of reward signals from a vision-language reward model
-Reduction of Convergence Rate.
-To understand the effects of reward noise on policy learning, it’s crucial to analyze its impact on the convergence rate.
-In sparse reward settings, the convergence rate – defined as the number of training iterations needed to learn an $\epsilon$-optimal policy – can be significantly affected by false positive rewards that are incorrectly given for policies that do not lead to the goal states.
-To illustrate this, we analyze the convergence rate by measuring the probability of learning acceptable policies ($P(\pi\in\Pi_{G})$) over each update iteration.
-The following theorem demonstrates how false positive rewards impact the convergence rate:
-Theorem 3.1 (Convergence rate reduction).
-In Actor-Critic algorithm, gradient ascent on $Q(s,a)\pi_{\theta_{i}}(a|s)$ pushes the next updated policy $\pi_{\theta_{i+1}}$ in the direction provided by the $Q$ value function.
-In the direction of maximizing $P(\pi\in\Pi_{G})$, the gradient of $Q(s,a)\pi(a|s)$ can be expressed as follows:
-$\nabla_{\theta}Q_{\phi}(s,a)\pi_{\theta}(a|s)=\textit{const}\cdot\mathbb{E}[G^{\pi\in\Pi_{B}}]\nabla_{\theta}\pi_{\theta}(a|s)$
-$\qquad+(\mathbb{E}[G^{\pi\in\Pi_{G}}]-\mathbb{E}[G^{\pi\in\Pi_{B}}])P(\pi\in\Pi_{G})\nabla_{\theta}\pi_{\theta}(a|s)$
-where $\Pi_{G}$ is the set of acceptable policies; $\Pi_{B}$ is the set of unacceptable policies that follow instructions but fail to reach the goal; $\mathbb{E}[G^{\pi}]$ is expectation of the cumulative rewards by executing policy $\pi$ in one episode.
-The proof is in Appendix A.2.
-In this context, false positive rewards correspond to rewards assigned to policies within the set $\Pi_{B}$, which represents undesirable policies.
-Therefore, $\mathbb{E}[G^{\pi\in\Pi_{B}}]$ can be interpreted as the expected cumulative false positive rewards.
-This theorem highlights how the presence of $\mathbb{E}[G^{\pi\in\Pi_{B}}]$ reduce the convergence rate by: (1) introducing deviation in the policy update directions (the first term), and (2) reducing the gradient of the desired policy update (the second term).
-Both effects can significantly hinder the convergence process.
-The issue of noisy rewards is not merely a theoretical concern; it has practical implications for the effectiveness of instruction-guided RL in real-world applications.
-In the next section, we will investigate how cosine similarity metric used in VLM-based reward models inevitably introduce false positive rewards.
-Additionally, we will empirically evaluate the impact of both false positives and false negatives on the learning efficacy of agents in sparse reward environments.
-4 Approximation Error of Cosine Similarity
-Cosine similarity is insufficient for evaluating the alignment between trajectories and instructions.
-In this section, we will discuss two fundamental issues associated with cosine similarity scores in RL contexts: state entanglement and composition insensitivity.
-The former issue, state entanglement, refers to the metric’s inability to recognize trajectories $\{\tau_{1},...,\tau_{n}\mid\tau_{i}\text{ induced by }\pi\in\Pi_{B}\}$ that, while being cosine similar to the target instruction in the embedding space, fail to reach the goal states in $S_{G}$.
-The latter issue refers to the metric’s tendency to reward trajectories that execute the sub-tasks (corresponding to individual instruction sentences $l_{k}$) in an incorrect order.
-This is problematic in RL tasks where the specific sequence of executing sub-tasks, as defined by the order of instructions in $L=\{l_{1},l_{2},\ldots,l_{n}\}$, is crucial for reaching the goal states.
-4.1 The Issue of State Entanglement
-State entanglement refers to the issue where the cosine similarity metric erroneously pays more attention to lexical-level similarity while lacking comprehension of the underlying state transitions.
+Our work contributes to this growing body of research by proposing a novel reward function that directly mitigates the impact of false positive rewards from VLM-based models, complementing approaches that use exploration policies to escape local optima.
+Furthermore, we tested the synergy of combining our reward function with exploration strategies, demonstrating how these approaches can be integrated for further advancements.
+3 Formal Problem Statement
+We frame our task as an MDP defined by a tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{P},s_{0},r^{e},\gamma\rangle$, where $\mathcal{S}$ represents a set of states $s\in\mathcal{S}$, $\mathcal{A}$ represents a set of actions $a\in\mathcal{A}$, and $\mathcal{P}(s^{\prime}|s,a)$ describes the dynamics of the environment.
+$s_{0}\in\mathcal{S}$ is the initial state and $\gamma\in(0,1)$ is the reward discount factor.
+$r^{e}(s,a)$ is the environmental reward function.
+An agent’s trajectory is a sequence of states and actions $\tau_{t}=\langle s_{0},a_{0},\ldots,s_{t}\rangle$.
+In this work, we focus on a sparse reward setting, where the agent receives a $+1$ reward only when reaching goal states $S_{G}\subset\mathcal{S}$, and 0 otherwise, with $|S_{G}|\ll|\mathcal{S}|$.
+This sparse reward setting motivates the use of expert instructions and VLMs to provide auxiliary reward signals for more effective RL.
+Specifically, we have a walkthrough $L$ that breaks down a complex task into $n$ expert-defined sub-tasks, each represented by a natural language instruction that is not necessarily atomic and can encompass multiple finer sub-goals ($L=\{l_{1},l_{2},\ldots,l_{n}\}$).
+By following these sequential instructions, the agent can navigate from the initial state towards the goal states.
+A dedicated non-Markovian VLM-based reward model $r^{v}(\tau_{t},l_{m(t)})$ is used to assess how well the agent’s trajectory at current time $t$ fulfills an instruction sentence $l_{m(t)}$.
+Here, $m(t)$ is a pointer at time step $t$ that indicates the current instruction the agent is trying to complete [^2] The use of non-Markovian reward functions in MDP has been well-established, particularly through the work on reward machines (Icarte et al., 2018; Corazza et al., 2022).
+For a complete evaluation, we also tested Markovian version of VLM reward function (Pixl2R) in our experiments..
+The VLM provides auxiliary rewards by evaluating the semantic similarity between $\tau_{t}$ and $l_{m(t)}$, as illustrated in Figure 2 (see Appendix A.1 for details).
+Figure 2: Illustration of reward from a VLM
+4 Theoretical Analysis
+In this section, we theoretically show how auxiliary rewards influence policy convergence.
+Initially, we establish that providing auxiliary rewards that reflect progress towards the goal state typically accelerates the convergence process compared to scenarios where only sparse environmental reward $r^{e}$ is used.
+Then, we show that the presence of false positive rewards has a more detrimental effect on convergence than false negatives.
+4.1 Auxiliary Rewards and Convergence
+With sparse rewards, the gradient landscape is nearly flat, making gradient-based updates indistinguishable from a random walk in parameter space (Antognini and Sohl-Dickstein, 2018).
+In this problem, we are interested in $D:=\|\theta_{goal}-\theta_{0}\|$ which is the distance in parameter space from initial parameters $\theta_{0}$ to goal parameters $\theta_{goal}$.
+We make the following assumption
+Assumption 4.1.
+Expert knowledge can guide the parameter search along a path in parameter space, defined by a sequence of $n$ intermediate parameter vectors $\theta_{1},\dots,\theta_{n}$, where each $\theta_{i}$ represents the parameters after learning sub-task $l_{i}$.
+As a result, the overall distance $D$ can be decomposed into segments: $D\approx\sum_{k=1}^{n-1}d_{i}$, where $d_{i}=\|\theta_{i+1}-\theta_{i}\|$.
+We also prove the following proposition
+Proposition 4.2.
+The sum of expected time for a series of random walks, each covering the shorter distance of an individual sub-task, is less than the expected time to travel the entire distance $D$ in one long random walk:
+$\frac{1}{n-1}\mathbb{E}[T_{D}]\leq\mathbb{E}\left[\sum_{i=1}^{n-1}T_{d_{i}}\right]<\mathbb{E}[T_{D}].$
+and thus show that subgoal-based auxiliary rewards improve the convergence of random walk optimization in a sparse reward landscape up to a factor of $(n-1)$ with details in Appendix A.2.1.
+4.2 Connection to Heuristic-Guided RL
+The problem and the possible solutions can be framed within the context of Heuristic-Guided Reinforcement Learning (HuRL) by Cheng et al.
+(2021).
+HuRL mandates that auxiliary reward signals serve as heuristics, where $h:\mathcal{S}\rightarrow\mathbb{R}$ approximates the future total rewards an agent expects to get starting from state $s$ under the optimal policy $\pi^{*}$ (i.e., $h(s)\approx V^{*}(s)$).
+These heuristics typically come from domain knowledge, aligning with expert instructions.
+We make the following assumption
+Assumption 4.3.
+The fulfillment of $l_{m(t)}$ can be captured by intermediate goal states $\mathcal{S}^{m(t)}$.
+This assumption is justified because often instructions in VLM come with end goals that can be indicated by definite states.
+For further discussions, please refer to Section 9.
+This assumption is required so that we can construct a domain-knowledge Markovian heuristic reward based on non-Markovian VLM rewards (see Appendix A.1.3 for the exact formula).
+With this construction, we are estimating the possible reward in the future by evaluating how well the agent has followed expert instruction in the past.
+It is consistent with the intuition of a heuristic reward based on semantic similarity because the better the agent has been following expert instructions, the more likely it is on the right track to collect more future rewards.
+Using HuRL framework allows us to analyze how false positive rewards influence on policy performance gap, defined as $V^{*}(s_{0})-V^{\pi}(s_{0})$.
+We will demonstrate that false positive rewards increase the upper bound of this gap, whereas false negative rewards maintain this upper bound.
+4.3 False Positives vs. False Negatives
+To begin, we provide the formal definition of false positive auxiliary rewards from both instruction-following (IF) and heuristic perspectives:
+Definition 4.4 (False Positive Rewards).
+A false positive reward occurs when:
+IF Perspective: For a trajectory $\tau_{t}$ that does not satisfy instruction $l_{m(t)}$, the VLM-based reward $r^{v}(\tau_{t},l_{m(t)})$ ranges between 00 and $1$.
+Heuristic Perspective: The heuristic $h(s_{t})>V^{*}(s_{t})$, overestimating the optimal value of $s_{t}$.
+Proposition 4.5.
+False positive rewards in the IF perspective imply false positive rewards in the heuristic perspective.
+The heuristic is approximated as $h(s_{t})\approx V^{*}(s_{t})=1\cdot\gamma^{\widetilde{T}-t}$, calculated based on an assumed optimal path length $\widetilde{T}$.
+Here, $\widetilde{T}-t$ measures the remaining steps towards the goal.
+When a trajectory $\tau_{t}$ receives a high reward but fails to fulfill instruction $l_{m(t)}$, it corresponds to a high $h(s_{t})$, thus a low $\widetilde{T}-t$.
+However, since the agent must eventually redo $l_{m(t)}$, $s_{t}$ is actually further from the goal than estimated.
+Therefore, the actual distance $T-t$ to reach the goal will exceed $\widetilde{T}-t$, and $V^{*}(s_{t})$, calculated with the actual $T$, will be smaller than $h(s_{t})$, derived from $\widetilde{T}$.
+This thus explains how a false positive reward from an instruction-following (or VLM) perspective leads to an overestimation of the heuristic.
+Researchers have advocated the benefits of pessimistic value estimation to enhance the stability of RL algorithms (Kumar et al., 2020; Jin et al., 2021).
+In HuRL, Cheng et al.
+(2021) further identify a beneficial property: when a heuristic is pessimistic with respect to $\mathcal{M}$, it results in a smaller upper bound on the performance gap.
+Definition 4.6 (Pessimistic $h$).
+Let Bellman operator $(\mathcal{B}h)(s,a)=r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[h(s^{\prime})]$.
+A heuristic function $h$ is said to be pessimistic with respect to an MDP $\mathcal{M}$ if $\max_{a}(\mathcal{B}h)(s,a)\geq h(s)$.
+This condition essentially means that the heuristic $h$ never overestimates the true value of a state.
+Proposition 4.7.
+Even if the heuristic remains conservative for all successor states, a single overestimation ($h(s)>V^{*}(s)$) can violate the pessimistic condition by causing $max_{a}(\mathcal{B}h)(s,a)<h(s)$.
+Informally, our result on the impact of false positive/negative rewards on convergence is as follows:
+Theorem 4.8.
+Borrowing from Cheng et al.
+(2021), the performance gap in RL when using heuristics can be broken down into a regret term and a bias term, where false negatives maintain the upper bound of bias while false positives increase the bias by breaking the heuristic’s pessimism, thereby potentially leading to slower convergence.
+See Appendix A.2.2 for formal version and detailed proofs.
+In the next section, we present a case study on cosine similarity metrics, demonstrating how they contribute to false positive rewards in learned reward models.
+5 False Positives From Cosine Similarity
+This section identifies and discusses two fundamental issues with cosine similarity scores in sequential decision-making contexts: state entanglement and composition insensitivity.
+The former issue, state entanglement, refers to the metric’s inability to recognize trajectories that, while being cosine similar to the target instruction in the embedding space, fail to reach the goal states in $S_{G}$.
+The latter issue refers to the metric’s tendency to reward trajectories even when the temporal relationships between sub-tasks are not satisfied.
+The Issue of State Entanglement State entanglement refers to the issue where the cosine similarity metric erroneously pays more attention to lexical-level similarity while lacking comprehension of the underlying state transitions.
 Consequently, rewards are given to trajectory-instruction pairs that are cosine similar in embedding space but in fact result in distinct state transitions.
@@
 Despite their lexical similarity, they lead to vastly different states.
-However, the cosine similarity metric may represent them as similar due to the shared action “take” and shared object “pill”, disregarding the critical difference in state outcomes.
+However, the cosine similarity metric may represent them as similar due to the shared words, disregarding the critical difference in state outcomes.
 Understanding state transitions is crucial in sequential decision-making scenarios.
-Otherwise, rewards may be given to trajectories that lead to unintended states, resulting in the learning of unacceptable policies.
-4.2 The Issue of Composition Insensitivity
-Composition insensitivity in cosine similarity metrics gives rise to two issues: (1) rewarding incomplete task execution – cosine similarity may incorrectly reward partial task completion, as even incomplete trajectories can receive high similarity score in the embedding space.
+Otherwise, rewards may be given to trajectories that lead to unintended states, potentially prolonging the path to the goal state by necessitating corrective actions or re-attempts.
+The Issue of Composition Insensitivity Composition insensitivity in cosine similarity metrics gives rise to two issues: (1) rewarding incomplete task execution – cosine similarity may incorrectly reward partial task completion, as even incomplete trajectories can receive high similarity score in the embedding space.
 For instance, in a task to “pick up a tool, then fix a machine,” the model might prematurely reward the agent for merely picking up the tool, neglecting the crucial repair action.
-This can lead to suboptimal learning, with the agent potentially getting stuck in easier, incomplete subtasks rather than fully accomplishing the task.
+We also observed this phenomenon particularly in the Montezuma environment, where RL agents tend to hack the reward system by focusing on the easiest actions that yield rewards (e.g., moving towards a direction) rather than executing more complex, timely actions.
+Eventually, this leads to an overestimation of the agent’s progress towards the ultimate goal.
 (2) insensitivity to the ordering of execution – cosine similarity often fails to adequately penalize incorrect execution sequences.
 In a safety protocol requiring an agent to “turn off the machinery, then perform maintenance,” the metric might assign high rewards based merely on the presence of relevant actions, disregarding their order.
-While some large language models have become sensitive to the order of tokens (Su et al.
-2024), the compact visual and sentence embeddings from multimodal models remains largely insensitive to sequential information (Pham et al.
-2021).
-This limitation can lead to potentially dangerous outcomes in scenarios where execution sequence is critical, such as in safety protocols or complex manufacturing processes.
-Figure 1 illustrates various scenarios where high similarity scores are erroneously assigned to false positive cases.
-To empirically demonstrate the issue, Section 4.3 presents experiments on false positive rewards and their impact on agent learning in sparse reward environments.
-Figure 3: Learned VLM models differentiate between matched and not-matched pairs, but struggle with O.O.D.
-cases.
-They incorrectly assign high scores to manipulated pairs, which should be low as the trajectories in the manipulated pairs do not fulfill the instruction
+In contrast to some advancements in language models, compact visual and sentence embeddings from multimodal VLMs remain largely insensitive to sequential information (Pham et al., 2021).
+When the task is order-sensitive, executing actions in the wrong sequence prolongs the path towards the goal state, as agents need to re-attempt the correct order.
+Figure 1 illustrates various scenarios where false positive rewards are erroneously assigned.
+To empirically demonstrate the issue, Section 5.1 presents experiments on these issues and their impact on agent learning in sparse reward environments.
+Figure 3: Learned VLM models performed badly with O.O.D.
+examples.
+They incorrectly assign high scores to manipulated pairs, which should be low as the trajectories in the manipulated pairs fail the instruction.
 Figure 4: The false positive vs. false negative oracle model. The false positive model get a more severe drop in the final training score.
 5.1 Experiments on Reward Noise Impact
-Our experiments test the following hypothesis: (H1) The two issues of state entanglement and composition insensitivity exist; (H2) false positive rewards are prevalent during training; (H3) similarity-based reward models lacking noise handling mechanisms underperform against strong intrinsic reward models in sparse reward environments; (H4) false negatives may not be as harmful as false positives.
+Our experiments test the following hypothesis: (H1) The two issues of state entanglement and composition insensitivity exist; (H2) false positive rewards are prevalent during training; (H3) VLM reward models lacking noise handling mechanisms underperform against intrinsic reward models in sparse reward environments; (H4) false negatives may not be as harmful as false positives.
 Setup.
-We evaluate these hypotheses through various challenging sparse-reward environments: (1) Crafter, an open-ended 2D Minecraft environment (Hafner 2021); (2) Montezuma’s Revenge, a classic hard adventure game in Atari (Bellemare et al.
-2013); and (3) Minigrid ‘Go To Seq’, a hard task involving long-horizon navigation and object interactions (Chevalier-Boisvert et al.
-2018).
-Two cosine similarity-based reward models were tested: (1) Pixl2R by Goyal et al.
+We evaluate these hypotheses through various challenging sparse-reward environments: (1) Crafter, an open-ended 2D Minecraft (Hafner, 2021); (2) Montezuma, a classic hard adventure game in Atari (Bellemare et al., 2013); and (3) Minigrid ‘Go To Seq’, a hard task involving long-horizon navigation and object interactions (Chevalier-Boisvert et al., 2018).
+A Markovian and a Non-Markovian reward model were tested: (1) Pixl2R by Goyal et al.
 (2020), which uses only the current video frame to determine if the goal state specified in the instruction has been reached; and (2) ELLM-, a variant of ELLM by Du et al.
@@
 Unlike ELLM, which queries instructions from LLMs in real-time, ELLM- directly uses preset expert instructions and compares them with the transition differences of the agent’s trajectory.
-The VLM backbones used are: (1) CLIP (Radford et al.
-2021), pretrained by image-text pairs; and (2) X-CLIP (Ma et al.
-2022), pretrained by video-text pairs.
+The VLM backbones used are: (1) CLIP (Radford et al., 2021), pretrained by image-text pairs; and (2) X-CLIP (Ma et al., 2022), pretrained by video-text pairs.
 To ensure high-quality finetuning data, we used internal information from the game engine to annotate expert trajectories from expert agents.
-To demonstrate how noisy reward signals hinder learning, we selected a strong intrinsic reward model DEIR (Zhang et al.
-2021) for comparison.
+To demonstrate how noisy reward signals hinder learning, we selected a strong intrinsic reward model DEIR (Zhang et al., 2021) for comparison.
 It provides auxiliary rewards based on observation novelty to encourage exploration.
@@
 Evaluation Metric.
-We used the score metric adapted from the Crafter benchmark (Hafner 2021) for performance evaluation, as it better captures an agent’s consistent performance across multiple subtasks.
-This metric follows the intuitive ‘higher is better’ principle.
-Unlike maximum total rewards metric, which fail to capture consistent performance in sparse reward settings, the score metric provides a more reliable measure of learning progress.
+We adopted the score metric from the Crafter benchmark (Hafner, 2021) for performance evaluation, as it effectively measures consistent performance across multiple subtasks in sparse reward environments.
+Unlike the maximum total rewards metric, which does not adequately reflect consistent performance, the score metric offers a more reliable indicator of learning progress.
+See Appendix A.5 for its formal definition.
 Reward Noise Issue.
@@
 The composition insensitivity test examined concatenated pairs of trajectory-instruction data.
-For example, given $(\tau_{1},l_{1})$ and $(\tau_{2},l_{2})$, we create a concatenated pair $(\tau_{1}+\tau_{2},l_{1}+l_{2})$.
+Given $(\tau_{1},l_{1})$ and $(\tau_{2},l_{2})$, we create a concatenated pair $(\tau_{1}+\tau_{2},l_{1}+l_{2})$.
 We then test two types of manipulations – (1) swapping the order within one modality: e.g., $(\tau_{2}+\tau_{1},l_{1}+l_{2})$; and (2) truncating one modality: e.g., $(\tau_{1},l_{1}+l_{2})$.
-Overall, we categorized the trajectory-instruction pairs for evaluation into three types: (1) matched pairs – these are positive examples where the trajectory correctly corresponds to the given instruction.
-(2) not-matched pairs – these are negative examples where the trajectory does not correspond to the given instruction.
-(3) manipulated pairs: derived from matched pairs by altering either the trajectory or the instruction.
-Ideally, manipulated pairs should receive low similarity scores as essentially the trajectory are not fulfilling the instruction.
-However, our results reveal that the reward model assigns high scores to these manipulated pairs, particularly in the “do not do $l_{k}$” case.
-This finding highlights the noise issue in cosine similarity-based reward models (see Figure 4).
+Overall, three types are evaluated: (1) matched pairs; (2) not-matched pairs and (3) manipulated pairs which are derived from matched pairs by polluting either the trajectory or the instruction.
+Our results reveal a critical flaw in the reward model: despite manipulated pairs that fundamentally fail to fulfill the instruction, the model paradoxically assigns high similarity scores (see Figure 4 for overall results and also Appendix A.6 for individual environments).
 It’s worth noting that the poor performance in the negation case aligns with broader challenges in natural language processing.
-Recent studies (Hossain et al.
-2022; Truong et al.
-2023) have highlighted that negation is central to language understanding but is not properly captured by modern language models.
-This limitation extends to VLMs and significantly impacts their ability to provide accurate rewards in complex scenarios and instructions.
-Figure 5: The heatmap shows the cumulative rewards received at various locations, with larger circle sizes indicating higher rewards.
+Recent studies (Hossain et al., 2022; Truong et al., 2023) have highlighted that negation is central to language understanding but is not properly captured by modern language models.
+This limitation extends to VLMs and directly leads to false positive rewards.
+Figure 5: The heatmap shows rewards received at various locations, with larger circle sizes indicating higher rewards.
 The later figures shows the offsets between the state where rewards are given and the actual goal-reaching state.
-Agents are getting both issues of false positives and false negatives during training
+Agents are getting both issues of false positives and false negatives
 Figure 6: The ratio of false positive rewards is significantly reduced after applying BiMI
+Table 1: Score metric across environments (equivalent to total rewards, higher is better).
+$\star$ denotes baseline intrinsic reward model.
+VLM reward models without noise handling underperformed.
+All are based on PPO.
+Models Type Monte. Minigrid Crafter % vs. DEIR
+PPO Pure $0.151$ $24.9$ $16.8$ $-28\%$
+DEIR $\star$ Intrinsic $0.174$ $55.5$ $19.7$
+Pixl2R VLM $0.142$ $12.4$ $9.40$ $\mathbf{-49\%}$
+ELLM- VLM $0.150$ $19.4$ $10.8$ $-41\%$
+Pixl2R + DEIR VLM + intr. $0.176$ $17.3$ $10.4$ $-38\%$
+ELLM- + DEIR VLM + intr. $0.178$ $30.9$ $11.8$ $-27\%$
 Prevalence of False Positives.
@@
 The heatmap revealed a concerning trend: RL agents engage in reward hacking, receiving rewards across vast areas of the environment rather than just at goal states.
-For instance, in Montezuma environment where the goal is to grab the key and escape the room, we observed that agents received rewards even for falling off cliffs, which undoubtedly contribute to the detrimental $\mathbb{E}[G^{\pi\in\Pi_{B}}]$ specified in Theorem 3.1.
+For instance, in Montezuma where the goal is to grab the key and escape the room, we observed that agents received rewards even for falling off cliffs, which undoubtedly contribute to false positive rewards.
 For environments without fixed camera views, we calculated the step offset between the current rewarded state and the actual goal state.
-A positive offset indicates that a reward was given before reaching the goal state, signifying a false positive reward, while a negative offset suggests that the agent reached the goal state but the reward model failed to recognize it, indicating a false negative reward (see Figure 6).
-Interestingly, we observed a large amount of negative step offsets in Minigrid environments.
-We attribute this to Minigrid’s abstract shape-based visual representations, which fall outside the VLM’s pretraining distribution.
+A positive offset indicates a false positive reward, as the reward was given before reaching the goal.
+Conversely, a negative offset indicates a false negative, where the agent reached the goal but the reward model failed to acknowledge it (see Figure 6).
+Interestingly, besides positive offsets, we observed a large amount of negative offsets in Minigrid environments.
+We attribute this to Minigrid’s abstract shape-based visual representations, which fall outside the VLM’s training distribution.
 Impact on Learning.
 We trained agents using learned VLM reward models and compared their learning efficacy against intrinsic reward models.
-As shown in Table 1, our results confirmed H3: instruction-guided RL agents using learned VLM reward models without noise handling consistently underperform compared to DEIR, the intrinsic reward-based RL agent.
+As shown in Table 1, our results confirmed H3: instruction-following RL agents using learned VLM reward models without noise handling consistently underperform compared to DEIR, the intrinsic reward-based RL agent.
 To investigate the impact of false negatives versus false positives (H4), we designed an oracle Pixl2R model with two variants: a false negative model and a false positive model.
 The false negative model only rewards the agent for reaching subgoal states described in the instruction, with a probability of xthat some rewarding states in the map are removed.
-In contrast, the false positive model rewards the agent for reaching every subgoal, but also introduces a 0.1 one-off reward for certain locations, covering xof the map.
-The results indicate that false negatives may be less detrimental to agent performance than false positives (see Figure 4).
-This performance difference can be explained through our theoretical framework:
-Table 1: Score metric across environments (equivalent to total rewards, higher is better).
-$\star$ denotes baseline intrinsic reward model.
-VLM reward models without noise handling underperformed.
-All models are based on PPO.
-Models Type Monte. Minigrid Crafter % vs. DEIR
-PPO Pure $0.151$ $24.993$ $16.863$ $-28\%$
-DEIR $\star$ Intrinsic $0.174$ $55.556$ $19.758$
-Pixl2R VLM $0.142$ $12.422$ $9.409$ $\mathbf{-49\%}$
-ELLM- VLM $0.150$ $19.406$ $10.826$ $-41\%$
-Pixl2R + DEIR VLM + intr. $0.176$ $17.372$ $10.440$ $-38\%$
-ELLM- + DEIR VLM + intr. $0.178$ $30.985$ $11.857$ $-27\%$
-  1. 1.
-    False negative models maintain $\mathbb{E}[G^{\pi\in\Pi_{G}}]-\mathbb{E}[G^{\pi\in\Pi_{B}}]$, ensuring steady gradient ascent towards maximizing $P(\pi\in\Pi_{G})$.
-  2. 2.
-    By decreasing $\mathbb{E}[G^{\pi\in\Pi_{B}}]$, false negative models minimize deviations from the target direction, leading to more stable learning.
-In contrast, scenarios with high $\mathbb{E}[G^{\pi\in\Pi_{B}}]$ (i.e., the false positive case) significantly reduce the gradient ascent rate and also introduce deviations from the target direction.
-The findings thus challenged the prevailing assumption that false negatives are the primary issue in the degeneracy of instruction-guided RL.
-See Appendix A.4 for more experimental details.
-5 Solution to the Reward Noise Issue
-5.1 Preferring Binary Signal
-Our experiments have shown that the issue of false negatives may be less detrimental to learning than false positives.
-Therefore, we propose a reward function that provides one-time binary rewards only when the similarity between the agent’s current trajectory and the instruction exceeds a high confidence threshold.
-This method effectively reduces the likelihood of rewarding unacceptable trajectories.
+In contrast, the false positive model rewards the agent for reaching every subgoal, but also introduces a small (0.1) one-off reward for certain locations, covering xof the map.
+The results indicate that false negatives were less detrimental to agent performance than false positives (see Figure 4).
+This performance difference can be explained through our theoretical analysis in Appendix A.2.2.
+6 Addressing the False Positive Issue
+Our proposed solution is not specific to cosine similarity’s approximation error, which serves as a case study.
+Rather, it’s a broad strategy for mitigating false positive rewards across various reward models, including noisy VLM reward sources.
+6.1 Binary Signal and Conformal Prediction
+Our experiments have demonstrated that false positives are more detrimental to learning than false negatives.
+Based on these findings, we propose a reward function that issues a one-time binary reward only when the similarity between the agent’s current trajectory and the instruction exceeds a high confidence threshold.
+This approach contrasts with previous methods, which provide continuous rewards whenever the reward score exceeds a predefined threshold, and continue to do so until reaching a maximum cap.
+Our method, however, delivers this reward only once.
+This approach minimizes the likelihood of accumulating false positive rewards while maintaining adherence to Assumption 4.3.
 To achieve this, we introduce a thresholding mechanism using a calibration set of true positive trajectory-instruction pairs.
@@
 Conversely, pairs exceeding $\hat{q}$ receive a one-time $+1$ reward, i.e., $r^{v}_{\textsc{Bi}}(\tau,l_{k})=\mathbf{1}_{\{p(l_{k}\mid\tau)\geq\hat{q}\}}$.
-This approach statistically guarantees a high probability (at least $1-\alpha$) that true positive pairs are recognized while minimizing frequency of false positives errors (Sadinle et al.
-2019).
-See Appendix A.5 for detailed threshold calculation.
+This thresholding approach leverages statistical properties studied by Sadinle et al.
+(2019), which ensures a high probability (at least $1-\alpha$) that true positive pairs are recognized while minimizing frequency of false positives errors.
+See Appendix A.4.2 for the pseudo-code of the instruction-following RL algorithm and Appendix A.8 for detailed threshold calculation.
 6.2 Mutual Information Maximization
@@
 $r^{v}_{\textsc{MI}}(\tau,l_{k})$ $=I(l_{k};\tau)=D_{KL}(p(l_{k},\tau)\mid\mid p(l_{k})p(\tau))$
-$=\mathbb{E}_{\tau\sim\pi_{\theta},l_{k}\sim L}[\log p(l_{k}\mid\tau)-\log p(l_{k})]$ (1)
-where $p(l_{k}\mid\tau)$ comes from the similarity score provided by VLMs, and $p(l_{k})$ is the frequency of the instruction $l_{k}$ being fulfilled by the agent’s trajectory.
+$=\mathbb{E}_{\tau\sim\pi_{\theta},l_{k}\sim L}[\log p(l_{k}\mid\tau)-\log p(l_{k})]$
+where $\tau=\langle s_{t-W},a_{t-W},\ldots,s_{t}\rangle$ is the agent’s trajectory up to current time step $t$, and $W$ is the memory size of the agent for its past trajectory.
+$p(l_{k}\mid\tau)$ comes from the similarity score provided by VLMs, referring to the likelihood of the instruction $l_{k}$ being fulfilled by a trajectory $\tau$.
+$p(l_{k})$ is overall likelihood of encountering the instruction $l_{k}$ in the learning environment.
 Therefore, the second term in the equation serves as a regularization term that downplays the significance of the reward signal when it is too frequent.
-For instance, if a VLM frequently detects that the agent’s actions are fulfilling the “climbing the ladder” instruction, even when the agent is performing unrelated tasks, the reward signal for this instruction will be downplayed.
+For instance, if a VLM frequently detects that the agent’s actions are fulfilling the “climbing the ladder” instruction, even when the agent is performing unrelated tasks, any reward signal from this instruction will be downplayed.
 $p(l_{k})$ is calculated as follows:
-$p(l_{k})=\mathbb{E}_{\tau\sim\pi_{\theta_{-1}}}\left[\frac{\sum_{t=1}^{T_{\tau}}\mathbf{1}_{\{p(l_{k}\mid\tau_{t})\geq\hat{q}\}}}{T_{\tau}}\right]$ (2)
+$p(l_{k})=\mathbb{E}_{\tau\sim\pi_{\theta_{-1}}}\left[\left(\sum\nolimits_{t=1}^{T_{\tau}}\mathbf{1}_{\{p(l_{k}\mid\tau_{t})\geq\hat{q}\}}\right)/T_{\tau}\right]$
 Here, $\tau_{t}$ is the agent’s trajectory up to time $t$, and if the VLM deems the trajectory as fulfilling the instruction (i.e., $p(l_{k}\mid\tau_{t})\geq\hat{q}$), we increment the count.
@@
 The subscript $\theta_{-1}$ in $\pi_{\theta_{-1}}$ indicates that the trajectories are sourced from rollouts in the previous policy iteration, acknowledging the impracticality of real-time computation of $p(l_{k})$ during an ongoing episode.
+Table 2: Model score across various environments.
+$\star$ is the baseline agents with a learned VLM-based reward model to compare with.
+BiMI significantly improves performance in Montezuma and Minigrid, while showing mixed results in Crafter due to task-specific characteristics
+Methods Montezuma % vs. $\star$ Minigrid % vs. $\star$ Crafter % vs. $\star$
+Pixl2R $\star$ $0.142\pm 0.003$ $12.4\pm 2.43$ $9.40\pm 0.022$
+Pixl2R + Bi $0.137\pm 0.009$ $-3.5\%$ $31.2\pm 2.04$ $+151\%$ $10.7\pm 0.784$ $+14\%$
+Pixl2R + BiMI $0.162\pm 0.022$ $+14\%$ $37.5\pm 7.83$ $+199\%$ $7.95\pm 0.351$ $-15\%$
+Pixl2R + DEIR $0.176\pm 0.009$ $+23\%$ $17.3\pm 0.51$ $+39\%$ $10.4\pm 1.015$ $+10\%$
+Pixl2R + BiMI + DEIR $0.267\pm 0.016$ $+88\%$ $57.7\pm 2.15$ $+364\%$ $11.0\pm 0.190$ $+17\%$
+ELLM- $\star$ $0.150\pm 0.004$ $19.4\pm 10.06$ $10.8\pm 1.017$
+ELLM- + Bi $0.151\pm 0.016$ $+0.6\%$ $29.7\pm 1.29$ $+53\%$ $11.1\pm 0.601$ $+3.2\%$
+ELLM- + BiMI $0.156\pm 0.014$ $+4.0\%$ $33.6\pm 3.99$ $+74\%$ $9.42\pm 0.267$ $-12\%$
+ELLM + DEIR $0.178\pm 0.029$ $+20\%$ $30.9\pm 3.50$ $+59\%$ $11.8\pm 1.152$ $+9.5\%$
+ELLM- + BiMI + DEIR $0.279\pm 0.078$ $+86\%$ $56.2\pm 6.19$ $+190\%$ $13.1\pm 0.393$ $+22\%$
+Figure 7: BiMI reward showed faster and higher success rates on difficult tasks in Montezuma
 To enhance the stability of the training process, we adopt a linearized version of the mutual information maximization approach, as proposed by Li et al.
@@
 It’s important to note that the BiMI approach primarily mitigates false positives (FP) rather than false negatives (FN).
-By implementing a high confidence threshold and a binary reward signal, we effectively reduce the likelihood of rewarding incorrect trajectories, thus addressing the FP issue.
-While this conservative approach may potentially increase FNs, as some correct trajectories might fall below the threshold, our experiments indicate that this trade-off is beneficial.
-We found that false positives are more detrimental to learning in instruction-guided RL tasks, justifying our preference for a method that prioritizes FP reduction.
-In the following sections, we will demonstrate how BiMI significantly improves the effectiveness of VLM-based reward models in instruction-guided RL tasks.
+Both Bi and MI aim to reduce the likelihood of rewarding unintended trajectories, thus addressing the FP issue.
+While this conservative approach may increase false negatives by pruning some correct trajectories, this trade-off is beneficial, as demonstrated by both our empirical and theoretical results.
 7 Experiments
-Figure 7: Besides the improvements of the score performance of agents across different environments with the BiMI reward function, it also collaborates well with intrinsic rewards.
-Combining both can lead to significant performance improvements
-Figure 8: Ablation on the components of BiMI reward function.
-The binary reward (Bi) alone led to a 36.5% improvement compared to original models.
-Excluding Crafter, Mutual Information (MI) provided a 23% further improvement over Bi alone
-We evaluate the BiMI reward function using baseline agents (Pixl2R and ELLM-) and their BiMI-enhanced counterparts, while also exploring potential synergies with intrinsic reward models.
-We use the score metric for evaluation.
-Table 2: Model score across various environments.
-$\star$ is the baseline agents with a learned VLM-based reward model to compare with.
-BiMI significantly improves performance in Montezuma and Minigrid, while showing mixed results in Crafter due to task-specific characteristics
-Methods Montezuma % vs. $\star$ Minigrid % vs. $\star$ Crafter % vs. $\star$
-Pixl2R $\star$ $0.142\pm 0.003$ $12.422\pm 2.439$ $9.409\pm 0.022$
-Pixl2R + Bi $0.137\pm 0.009$ $-3.5\%$ $31.236\pm 2.040$ $151\%$ $10.735\pm 0.784$ $14\%$
-Pixl2R + BiMI $0.162\pm 0.022$ $14\%$ $37.507\pm 7.832$ $199\%$ $7.951\pm 0.351$ $-15\%$
-Pixl2R + DEIR $0.176\pm 0.009$ $23\%$ $17.372\pm 0.514$ $39\%$ $10.440\pm 1.015$ $10\%$
-Pixl2R + BiMI + DEIR $0.267\pm 0.016$ $88\%$ $57.759\pm 2.157$ $364\%$ $11.014\pm 0.190$ $17\%$
-ELLM- $\star$ $0.150\pm 0.004$ $19.406\pm 10.067$ $10.826\pm 1.017$
-ELLM- + Bi $0.151\pm 0.016$ $0.6\%$ $29.788\pm 1.290$ $53\%$ $11.175\pm 0.601$ $3.2\%$
-ELLM- + BiMI $0.156\pm 0.014$ $4.0\%$ $33.683\pm 3.990$ $74\%$ $9.425\pm 0.267$ $-12\%$
-ELLM + DEIR $0.178\pm 0.029$ $20\%$ $30.985\pm 3.507$ $59\%$ $11.857\pm 1.152$ $9.5\%$
-ELLM- + BiMI + DEIR $0.279\pm 0.078$ $86\%$ $56.281\pm 6.193$ $190\%$ $13.170\pm 0.393$ $22\%$
-Figure 9: BiMI reward showed faster and higher success rates on difficult tasks in Montezuma
-6.1 Montezuma’s Revenge
-In Montezuma, Pixl2R$+$BiMI demonstrated an 14% performance increase compared to the original models (see Table 2), which is slightly below our expectations.
+We continue to evaluate using Markovian Pixl2R and Non-Markovian ELLM- and their BiMI-enhanced counterparts, while also exploring potential synergies with the intrinsic reward model DEIR.
+We follow the same experimental setup as in Section 5.1, with additional details in Appendix A.9.
+7.1 Main Results
+In Montezuma, Pixl2R$+$BiMI demonstrated 14% performance increase compared to the original models (see Table 2), which is slightly below our expectations.
 We attribute this result to BiMI’s intentional strategy of providing less frequent discrete rewards.
@@
 Meanwhile, DEIR’s intrinsic rewards fill the gaps between these anchor points, encouraging the agent to explore efficiently in the interim.
-As illustrated in Figure 6, BiMI rewards are now concentrated on key locations.
+See Figure 6 for a quantitative analysis: BiMI rewards are now concentrated on key locations.
 A significant improvement is the minimal rewards given for falling off cliffs, which was a common source of false positives in the original model.
-Figure 9 demonstrates higher success rate grabbing the key in the first room, one of the most difficult task in Montezuma, highlighting the effectiveness of the proposed reward function and its collaboration with intrinsic reward models in guiding agents to solve difficult sparse-reward tasks.
-6.2 Minigrid
-ELLM-$+$BiMI achieved a remarkable 74% improvement in performance compared to the original models.This substantial gain is particularly noteworthy given the unique challenges presented by Minigrid.
-The abstract, shape-based visuals in Minigrid diverge drastically from the natural images used in VLMs’ pretraining, preventing the models from effectively utilizing their prior pretraining knowledge.
-Consequently, VLMs struggled to accurately assess similarities between Minigrid’s abstract visuals and textual instructions, resulting in highly noisy reward signals.
-The significant improvement demonstrated by BiMI underscores its effectiveness in handling noisy signals, directly addressing our primary research challenge.
-This capability is crucial for deploying instruction-guided agents in real-world, unfamiliar scenarios, where visual inputs often deviate from the VLMs’ training distribution, leading to noisy reward signals.
-6.3 Crafter
-We observed an intriguing pattern of results.
-The Bi component alone led to 14% and 3.2% improvement in performance over the original models.
-However, contrary to our observations in other environments, the addition of the MI component actually decreased this improvement.
-This unexpected outcome can be attributed to the unique nature of Crafter task, where agents must repeatedly achieve the same subtasks (e.g., drinking water) for survival.
-The MI component, designed to discourage over-reliance on frequently occurring signals, inadvertently penalized the necessary repetition of survival-critical actions.
-Furthermore, note that instruction-guided RL agents, regardless of the reward model employed, were unable to outperform pure RL agents in Crafter.
-This discrepancy is due to the open-world nature of Crafter, which requires dynamic strategies and real-time decision-making that our testing setups did not fully capture.
-Despite these challenges, it is noteworthy that Bi alone still managed to improve performance over vanilla VLM-based reward models, suggesting that reducing false positives is still beneficial across all testing environments.
-The combination of BiMI with DEIR (the intrinsic reward model) also showed promising results, indicating a productive balance between exploration (driven by DEIR) and exploitation (guided by BiMI instruction reward).
-6.4 Ablation Study and Overall Performance
+Figure 7 demonstrates a higher success rate in grabbing the key in the first room, one of the most difficult tasks in Montezuma, highlighting the effectiveness of the proposed reward function and its synergy with intrinsic reward models in guiding agents to solve difficult sparse-reward tasks.
+See results for other environments in Appendix A.10.
+7.2 Overall Performance and Ablation Study
+The overall improvements were substantial.
+As shown in Table 2, BiMI led to a 67% improvement for Pixl2R (Markovian) and a 22% improvement for ELLM- (Non-Markovian).
+These results are also illustrated in Figure 17 in Appendix A.10.
 Our ablation study highlights the distinct contributions of the binary reward (Bi) and Mutual Information (MI) components within the BiMI framework.
@@
 When excluding the results from Crafter, MI component further contributes a 23% improvement over the binary reward alone.
-The overall improvements were substantial.
-BiMI led to a 67% improvement for Pixl2R and a 22% improvement for ELLM-.
-These results are illustrated in the line plots in Figure 8 and Figure 8 and summarized in Table 2.
-See Appendix A.6 for more details.
 8 Conclusion
-Our research reveals two key findings in instruction-guided reinforcement learning using VLM-based reward models: (1) false positive rewards, rather than false negatives, are more detrimental to policy learning; and (2) our proposed BiMI reward function significantly mitigates the impact of noise in reward models.
-Our work addresses a crucial gap in prior work, paving the way for more reliable and effective instruction-guided RL systems in practical settings.
-Limitation and Future Work.
+We reveal two findings in VLM-based reward functions for RL agents: (1) false positive rewards, rather than false negatives, are more detrimental to policy learning; and (2) our proposed BiMI reward function advocates for pessimistic rewarding, significantly mitigating the slowdown in learning caused by false positives.
+Our results are supported by both theoretical analysis and empirical validation across three challenging embodied tasks.
+9 Limitations
 Our study primarily focused on linear sequences of language instructions, excluding more complex cases.
@@
 (2024).
+However, we believe that finetuning large VLMs simultaneously during reinforcement learning (RL) is computationally expensive and may not be practical in real-world scenarios.
+We acknowledge that mutual information (MI) has been previously explored as a reward objective in RL.
+However, our work uncovers a novel insight: MI is particularly well-suited for preventing over-reliance on excessively frequent reward sources, which in turn helps to mitigate the issue of too many false positive rewards.
+This research is innovative in explicitly addressing the practical challenges posed by false positive rewards in instruction-following learning agents that utilize VLM rewards.
+We believe that our proposed BiMI reward function represents a novel application of MI specifically tailored for VLM-based rewards.
+There is a gap in providing a rigorous theoretical foundation for why our theoretical findings extend to non-Markovian reward models like ELLM-.
+Our Assumption 4.3 facilitates the integration of non-Markovian reward models into the HuRL framework, though some might argue this assumption is strong.
+However, with advancements in deep RL, the distinction between non-Markovian and Markovian models has become increasingly blurred.
+For instance, the pioneering DQN (Mnih, 2013) utilized sequences of past observations as input, ostensibly making it non-Markovian.
+Yet, with deep learning’s capacity to encapsulate complex state histories into a representation that can be treated as Markovian at a higher abstraction level, some have argued that non-Markovian elements can be effectively reconsidered as Markovian within this new context (Hausknecht and Stone, 2015).
+Our experiments were conducted in extremely sparse reward settings, with tasks intentionally designed to be challenging by requiring long-horizon planning and sequential execution.
+This setup may overemphasize the impact of false positive rewards.
+We acknowledge that in certain cases, rewarding partially correct trajectories may be beneficial for exploration.
+However, to prevent the agent from overvaluing these partial solutions, it is essential that the reward signal remains conservative, providing a pessimistic estimate of the value of partially correct trajectories.
 Appendix A Technical Appendix
-Continued from the main text of Overcoming Reward Model Noise in Instruction-Guided Reinforcement Learning, the technical appendix consist of the following:
-  • §A.1 Property of Sparse Reward Function, which discuss why sparse reward function results in the categorization of policies into a set of acceptable policies and a set of unacceptable policies but not others.
-  • §A.2 Proof of the Reduction of Convergence Rate, which is referred in Section 3.
+Continued from the main text of Understanding and Mitigating Noise in VLM Rewards, the technical appendix consists of the following:
+  • §A.1 Detailed Problem Setting, which provides a detailed problem setting of the instruction-following reinforcement learning. This is referred to in Section 3 and Section 4.
+  • §A.2 Theoretical Analysis and Proofs, which is referred to in Section 4
+  • §A.3 Additional Notes on Theoretical Setting
   • §A.4 Implementation Details of the Experiments, which specify the implementation details for both the first and second stage experiments in Section 5.1 and Section 7 respectively.
   • §A.5 Additional Details of the Experiments of False Positive Rewards, which is referred in Section 5.1.
+  • §A.6 Details of Showing the Prevalence of False Positives in VLM Cosine Similarity Scores, which provides extra figures that cannot be in the main body due to page limit.
+  • §A.7 Impact of Reward Model Quality on Policy Learning Efficiency: Supplementary Experiments and Quantitative Analysis in Montezuma, which provides extra evaluation results to further illustrate false positive rewards hinders agent learning.
   • §A.8 Pseudo-code for Empirical Quantile Calculation for Binary Signal Threshold, which is referred in Section 6.1.
-  • §A.6 Additional Details of the Experiments of BiMI Reward Function, which is referred in Section 6.
-  • §A.7 Cosine Similarity Score Evaluation for Different Environments, are details of the cosine similarity score evaluation for different environments, which is referred in Section 4.3.
-  • §A.8 Line Plots over Average Reward for Different Environments and Ablation Study, which provides performance comparison of different models using average reward metric, the conventional metric used in RL, though they are less informative than the score metric used in the main text when dealing with sparse reward environments.
-A.1 Sparse Reward Function and Range-SOAP
-Formally, Abel et al. (2021) defined that:
-Theorem A.1 (Abel et al. (2021)).
-A reward function realizes a Range Set of Acceptable Policies (Range-SOAP) $\Pi_{G}$ when there exists an $\epsilon\geq 0$ such that every $\pi_{g}\in\Pi_{G}$ is $\epsilon$-optimal in start-state value, $V^{*}(s_{0})-V^{\pi_{g}}(s_{0})\leq\epsilon$, while all other policies are worse.
-When the reward signal is sparse, the agent only receives a reward upon reaching some goal states, and the reward function does not provide any feedback during the intermediate steps.
-We argue that the sparse reward function realizes a Range-SOAP, leading to the categorization of policies into acceptable and unacceptable policies.
-Proposition A.2.
-Sparse reward function “realizes” Range-SOAP (i.e., Range Set of Acceptable Policies).
-Justification:
-A sparse reward function, where the agent only receives a reward upon completing the task, cannot prefer policies that lead to shorter task completion times.
-This is because either the agent completes the goal very quickly or slowly, they will receive nearly the same amount of cumulative rewards, and the reward function will not show a strong preference.
-Since the sparse reward function does not induce a strict partial ordering on the policies, we say this reward function cannot realize a Partial Ordering (PO) task.
-Specifically, a PO on policies is a generalization of a Set of Acceptable Policies (SOAP) task.
-In a PO, the agent specifies a partial ordering on the policy space, where some policies are identified as “great”, some as “good”, and some as “bad” to strictly avoid, while remaining indifferent to the rest.
-Therefore, the sparse reward function can realize a Set of Acceptable Policies (SOAP), where there is a set of policies that are all considered “good” or near-optimal, while all other policies are worse.
-Furthermore, the sparse reward function will lead to a Range-SOAP, rather than an Equal-SOAP.
-Specifically, Equal-SOAP is a SOAP where all the acceptable policies are equally optimal in start-state value.
-This is because the good policies in the SOAP may differ slightly in their start-state values, as some may reach multiple goal states in the environment and thereby receiving different cumulative rewards.
-Therefore, the sparse reward function will realize a Range-SOAP, where there is a range of acceptable policies that are all near-optimal in start-state value.
-A.2 Proof of the Reduction of Convergence Rate
-Specifically, the update rule of Actor-Critic algorithm is:
-  • Critic:
-    $\phi\leftarrow\phi-\alpha_{\phi}\nabla_{\phi}(\delta)^{2}$ (4)
+  • §A.9 Additional Implementation Details of the Experiments of BiMI Reward Function, which is referred in Section 7.
+  • §A.10 Detailed Experiment Results of BiMI Reward Function, which provides detailed experiment results of BiMI reward function, which is referred in Section 7.
+  • §A.11 Deprecated Convergence Analysis, It is retained here solely for reference and backtracking purposes.
+As the more rigorous theoretical analysis is constructed, the content of this section is now deprecated.
+A.1 Detailed Problem Setting
+A.1.1 The Pointer Mechanism in Reward Machine
+In our problem setting, $r^{v}(\tau_{t},l_{m(t)})$ is the VLM based reward at time step $t$ evaluated using the sub-trajectory $\tau_{t}$ is the sub-trajectory up to time $t$.
+Existing VLM-reward implementations applied a pointer mechanism that decide which instruction should the agent consider at the current step.
+We follow this implementation and denote the pointer as $m(t)$.
+It indicates the current instruction of the agent is trying to complete at time step $t$.
+$m(t)$ is updated according to the following rule:
+$m(t+1)=\begin{cases}1&\text{if $t=0$}\\ m(t)+1&\text{if instr. $l_{m(t)}$ completed at $t$}\\ m(t)&\text{otherwise}\end{cases}$ (2)
+In both the ELLM and the Pixl2R framework that uses VLM-based rewards, the assessment of whether instruction $l_{m(t)}$ is completed at time $t$ is based on the cumulative reward for that instruction.
+The pointer remains on the current instruction until the accumulated reward reaches a predetermined threshold.
+For a detailed pseudo-code of the existing VLM reward + RL algorithm, please refer to Appendix A.4.2.
+Note that our proposed BiMI reward function uses a different mechanism to determine if instruction $l_{m(t)}$ is completed at time $t$.
+In logic, an atomic sentence is a type of statement which cannot be broken down into other simpler sentences.
+While the pointer mechanism effectively enforces the sequential ordering of instructions at a higher level, each individual instruction within the expert walkthrough is not necessarily atomic.
+This means that a single instruction can encapsulate multiple finer-grained requirements for the agent’s behavior.
+Therefore, even though the pointer mechanism is implemented, the internal composition of each instruction is not strictly enforced.
+Consequently, the issue of “composition insensitivity” can arise when the VLM tries to align agents’ trajectories with non-atomic instruction sentences.
+Just as potential-based reward shaping (PBRS) modifies the original MDP into a reshaped MDP through the use of auxiliary rewards, the HuRL framework similarly transforms the original MDP $\mathcal{M}$ into a new MDP ~$\widetilde{\mathcal{M}}=\langle\mathcal{S},\mathcal{A},\mathcal{P},s_{0},\widetilde{r},\widetilde{\gamma}\rangle$, employing a heuristic-based auxiliary reward $h$, along with a coefficient $\beta\in[0,1]$ that scales the auxiliary rewards.
+Thus, the updated reward function can be written as:
+$\widetilde{r}(s,a)=r^{e}(s,a,s^{\prime})+(1-\beta)\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a}[h(s^{\prime})]\quad\text{and $\widetilde{\gamma}=\beta\gamma$}$
+We also provide the formal definition of false negative reward as below.
+Definition A.1 (False Negative Rewards).
+A false negative reward occurs when:
+Instruction-Following Perspective: The VLM-based reward $r^{v}(\tau_{t},l_{m(t)})\approx 0$ for a trajectory $\tau_{t}$ that does satisfy instruction $l_{m(t)}$.
+Heuristic Perspective: The heuristic $h(s_{t})\approx 0<V^{*}(s_{t})$, underestimating the optimal value of $s_{t}$.
+A.1.2 Detailed Terminology
+This section contains detailed terminology from both the standard RL and the HuRL framework.
+We start by setting clear the terminology.
+The original MDP is defined by a tuple $\langle\mathcal{S},\mathcal{A},\mathcal{P},s_{0},r^{e},\gamma\rangle$ as set out in the main text.
+Let $G^{e}(\tau)$ denote the cumulative environment reward of a single trajectory $\tau=\langle s_{0},a_{0},\ldots,s_{T}\rangle$ where $T$ is the total time of the trajectory, we have
+$G^{e}(\tau)=\sum_{t=0}^{T}\gamma^{t}r^{e}(s_{t},a_{t}).$ (3)
+A policy $\pi_{\theta}$ is a conditional distribution $\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})$ parameterized by $\theta$, where $\Delta(\cdot)$ denotes the space of probability distribution.
+The probability of a trajectory $\tau$ being generated, given that actions come from $\pi_{\theta}$ and the starting state is $s_{0}$, is expressed as:
+$P(\tau\mid\pi_{\theta})=\prod_{t=0}^{T}\mathcal{P}(s_{t+1}\mid s_{t},a_{t})\pi_{\theta}(a_{t}\mid s_{t})$ (4)
+Because a policy $\pi$ is stochastic, different trajectories could be generated from the same policy for a given initial state.
+The objective function to be optimized is the expected cumulative environment reward of trajectories generated by $\pi$ starting from state $s_{0}$.
+It is also called the state value function of $s_{0}$ and denoted as:
+$V^{\pi}(s_{0}):=\mathbb{E}_{\rho_{s}^{\pi}}\left[G^{e}(\tau)\right]=\sum_{\textnormal{all $\tau_{i}$ generated by $\pi_{\theta}$}}G^{e}(\tau_{i})\cdot P(\tau_{i}|\pi)$ (5)
+where $\rho_{s}^{\pi}$ denotes the trajectory distribution of $s_{0},a_{0},s_{1},...$ induced by running $\pi$ starting from $s_{0}=s$. We defined an acceptable policy $\pi_{A}$ as
+$\pi_{A}:V^{\pi_{A}}\geq V^{\pi^{*}}-\epsilon$ (6)
+where $\pi^{*}$ is the optimal policy and $\epsilon$ is a control parameter that quantifies the acceptance of other policies.
+Acceptable policies defined in this way are also referred to as $\epsilon$-optimal policies.
+We set $V^{*}=V^{\pi^{*}}$ for easy reference.
+Any policy that is not an acceptable policy is an unacceptable policy $\pi_{U}$.
+We define a good trajectory $\tau_{G}$ as
+$\tau_{G}:G^{e}(\tau_{G})\geq V^{*}-\epsilon.$ (7)
+Any trajectory that is not a good trajectory is a bad trajectory $\tau_{B}$.
+Note that this way of distinguishing good/bad trajectories and acceptable/unacceptable policies is explicitly based on the environment reward $r^{e}$, and introducing additional or alternative rewards into the system does not change this definition.
+For easy reference, we call any trajectory that reaches a goal state a goal trajectory, and we call any policy that can generate at least one goal trajectory a goal policy.
+We denote the state distribution of a policy $\pi$ at time $t$ as $d^{\pi}_{t}$.
+Thus, the discounted average state distribution of a policy $\pi$ can be expressed as $d^{\pi}=(1-\gamma)\sum^{\infty}_{t=0}\gamma^{t}d^{\pi}_{t}$, where $(1-\gamma)$ is the normalization factor to ensure the result is a proper probability distribution.
+We also define a shorthand notation – for a state distribution $d\in\Delta(\mathcal{S})$, we define $V(d)=\mathbb{E}_{s\sim d}\left[V(s)\right]$.
+When the Heuristic-Guided Reinforcement Learning (HuRL) framework is applied, the reshaped MDP $\widetilde{\mathcal{M}}$ has separated reshaped reward $\widetilde{r}$ and separated discount factor $\widetilde{\gamma}=\beta\gamma$, where $\beta\in[0,1]$ is the hyperparameter that scales the auxiliary heuristic rewards.
+We denote the value function of the reshaped MDP as $\widetilde{V}$, and the optimal policy and its value function under $\widetilde{\mathcal{M}}$ as $\widetilde{\pi}^{*}$ and $\widetilde{V}^{*}(\widetilde{V}^{\widetilde{\pi}^{*}})$.
+The discussion regarding the convergence guarantee of learning $\pi$ in $\widetilde{\mathcal{M}}$ resulting in the optimal policy in the original MDP $\mathcal{M}$ can be found in Appendix A.3.1.
+For the purposes of this paper, it is sufficient to note that the authors of HuRL have provided a trick to ensure convergence.
+Moreover, empirical experiments conducted by the authors demonstrate that HuRL converges even without this trick.
+It is also important to distinguish the Bellman backup equation under the two different MDPs: the original $\mathcal{M}$ and the reshaped $\widetilde{\mathcal{M}}$.
+By definition, $(\mathcal{B}h)(s,a)=r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[h(s^{\prime})]$.
+In contrast, (~$(\widetilde{\mathcal{B}}h)(s,a)=\widetilde{r}(s,a)+\widetilde{\gamma}\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[h(s^{\prime})]$.
+Nevertheless, the two Bellman backup equations possess a remarkable property – they are essentially equivalent to each other:
+(~$(\widetilde{\mathcal{B}}h)(s,a)$ $=\widetilde{r}(s,a)+\widetilde{\gamma}\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[h(s^{\prime})]$
+$=\left(r(s,a)+(1-\beta)\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[h(s^{\prime})]\right)+\beta\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[h(s^{\prime})]$
+$=r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[h(s^{\prime})]$
+$=(\mathcal{B}h)(s,a)$ (8)
+A.1.3 Detailed Construction of Markovian Heuristic using Non-Markovian Rewards from VLM
+Specifically,
+$h(s_{t}):=\begin{cases}A(s_{t})V^{*}(s_{t})&\text{if $s_{t}\in S^{m(t)}$}\\ 0&\text{otherwise}\end{cases}$ (9)
 where
-    $\delta=\mathbb{E}_{\pi_{\theta}}[G^{\pi_{\theta}}-Q_{\phi}(s,a)]$ is the Monte Carlo (MC) estimation error, $Q_{\phi}(s,a)$ is the $Q$-value function which measures the expected discounted cumulative reward given the state $s$ and action $a$, and $\pi_{\theta}$ is the policy.
-    $G^{\pi_{\theta}}$ is the rollout cumulative rewards from the trajectory $\tau^{\pi_{\theta}}$ generated from $\pi_{\theta}$.
-  • Actor:
-    $\theta\leftarrow\theta+\alpha_{\theta}\frac{Q_{\phi}(s,a)\nabla_{\theta}\pi_{\theta}(a|s)}{\pi_{\theta}(a|s)}$ (5)
-We need to make the following assumptions to simplify the theoretical analysis:
+$A(s_{t}):=\frac{\Pi_{t^{\prime}=0}^{t}r^{v}(\tau_{t^{\prime}},l_{m(t^{\prime})})}{\Pi_{t^{\prime}=0}^{t}r^{v}(\tau^{*}_{t^{\prime}},l_{m(t^{\prime})})}\in[0,1]$ (10)
+is a scaling factor that quantifies how closely the agent is following previous instructions.
+A.2 Theoretical Analysis and Proofs
+A.2.1 Convergence time on a sparse-reward landscape
+We can pin down a characteristic convergence time on a sparse-reward landscape. The sparse-reward setting enforces that
+$r^{e}(s_{t})=\begin{cases}1&\text{if $s_{t}\in\mathcal{S}_{G}$}\\ 0&\text{otherwise}\end{cases}.$ (11)
+It is clear from the definition that a good trajectory must always be a goal trajectory, and the cumulative reward is simply the reward at the final goal state so $G^{e}(\tau)=\gamma^{T}$ in this setting.
+For a randomly initialized policy, it is highly unlikely that the initial distribution of trajectories contains any goal trajectory due to the sparsity of goal states.
+The optimization of $V^{e}_{\pi_{\theta}}$ thus consists of two parts.
+The first part is to search for a goal trajectory.
+The gradient landscape is almost 0 everywhere, except for cases where a trajectory is $\delta$-close to a goal trajectory.
+Here $\delta$ is the differential unit in the numerical differentiation used in the gradient calculation
+$\theta=\theta+\alpha\nabla_{\theta}V_{\pi_{\theta}}$ (12)
+such that $\delta$-close means being numerically accessible within a distance of $|\delta|$ in parameter space.
+And the second part is to reduce $T$ so that goal trajectories become good trajectories and consequently achieving acceptable policies.
+For the first part, searching for a trajectory for the target is effectively a random walk in the $d$-dimensional parameter space due to the flat gradient landscape.
+Lemma A.2.
+For a random walk in $n$-dimensional space, the expected number of steps $T_{D}$ needed to travel a distance of $D$ scales with $D^{2}$.
+Proof: let $\vec{X_{1}}$, $\vec{X_{2}}$, … $\vec{X_{T}}$ be IID random unit vectors uniformly distributed on a $(d-1)$-dimensional sphere $\mathcal{S}^{d-1}\subset\mathbb{R}^{d}$, where $\vec{X_{i}}=(X_{i1},...,X_{id})$ and $|\vec{X_{i}}|^{2}=\sum_{j=1}^{d}X^{2}_{ij}=1$.
+Let $\vec{S_{T}}:=\sum_{i=1}^{T}\vec{X_{i}}$.
+By a $n$-dimensional cosine rule we have
+$|\vec{S}_{T}|^{2}$ $=|\vec{S}_{T-1}|^{2}+2\vec{S}_{T-1}\cdot\vec{X}_{T}+|\vec{X}_{T}|^{2},$ (13)
+and because $\mathbb{E}[\vec{X}_{T}]=\vec{0}$
+$\mathbb{E}[|\vec{S}_{T}|^{2}]$ $=\mathbb{E}[|\vec{S}_{T-1}|^{2}]+\mathbb{E}[2\vec{S}_{T-1}\cdot\vec{X}_{T}]+1$ (14)
+$=\mathbb{E}[|\vec{S}_{T-1}|^{2})]+2\vec{S}_{T-1}\mathbb{E}[\vec{X}_{T}]+1$ (15)
+$=\mathbb{E}[|\vec{S}_{T-1}|^{2}]+1$ (16)
+$=\mathbb{E}[|\vec{S}_{T-2}|^{2}]+1+1$ (17)
+$\vdots$ (18)
+$=T$ (19)
+i.e.
+$\mathbb{E}[|\vec{S}_{T}|]\sim\sqrt{T}$.
+When a policy is randomly initialized with $\theta_{0}$, the distance $D$ to a goal policy, $D:=||\theta_{goal}-\theta_{0}||$ is fixed and is the distance the random walk needs to travel ($\mathbb{E}[|\vec{S}_{T}|]=D\sim\sqrt{T_{D}}$), so the characteristic time needed to travel this distance $T_{D}\sim D^{2}$ as we have shown above.
+See 4.1 Auxiliary rewards essentially open the path for a divide-and-conquer approach by introducing intermediate rewards in the learning process. We introduce BiMI rewards as
+$r^{v}_{\textsc{BiMI}}(\tau,l_{k})=\max(\mathbf{1}_{\{p(l_{k}\mid\tau)\geq\hat{q}\}}-p(l_{k}),0)$ (20)
+and the cumulative reward of a single trajectory in the presence of BiMI rewards becomes
+$G^{v}_{\textsc{BiMI}}(\tau)=\sum_{t=0}^{T}\gamma^{t}r^{v}_{\textsc{BiMI}}(\tau,l_{m(t)}).$ (21)
+[…]
+See 4.2 Proof: the expected time taken for each of the sub-tasks then scales with ${}d_{i}^{2}$ respectively (Lemma A.2), and we have
+$d_{1}^{2}+d_{2}^{2}+...+d_{n-1}^{2}$ $<(d_{1}+d_{2}+...+d_{n-1})^{2}$ (22)
+$\mathbb{E}\left[\sum_{i=1}^{n-1}T_{d_{i}}\right]$ $<\mathbb{E}[T_{D}]$ (23)
+because $d_{i}>0\forall i$.
+We can also work out that the upper bound for this improvement is a factor of $n-1$ by invoking the Cauchy–Schwarz inequality $\left(\sum_{i=1}^{n}u_{i}^{2}\right)\left(\sum_{i=1}^{n}v_{i}^{2}\right)\geq\left(\sum_{i=1}^{n}u_{i}v_{i}\right)^{2}$:
+$(d_{1}^{2}+d_{2}^{2}+...+d_{n-1}^{2})(1^{2}+1^{2}+...+1^{2})$ $\geq(d_{1}+d_{2}+...+d_{n-1})^{2}$ (24)
+$(d_{1}^{2}+d_{2}^{2}+...+d_{n-1}^{2})$ $\geq\frac{1}{n-1}D^{2}$ (25)
+$\mathbb{E}\left[\sum_{i=1}^{n-1}T_{d_{i}}\right]\geq\frac{1}{n-1}\mathbb{E}[T_{D}]$ (26)
+and the equality sign holds (indicating maximal improvement) when $d_{1}=d_{2}=...=d_{n-1}$.
+The intuitive interpretation is that the divide-and-conquer approach is the most effective when the task is divided evenly into subtasks.
+Additionally, the presence of $p(l_{k})$ introduces some lexical-level tolerance in the learning process and results in a considerably larger $\delta$-close radius which could result in further improvement in the convergence rate.
+However, as we will discuss in the next section, introducing too much tolerance could harm the convergence rate rather than improve it.
+A.2.2 False Positive and the Violation of Pessimistic Property of Heuristics
+See 4.7
+Proof.
+We want to show that given:
   1. 1.
-    We use $\Pi_{G}$ to represent the set of acceptable policies and use $\Pi_{B}$ as an arbitrary set of “not acceptable” policies that are learned to follow instruction but fail to reach subgoals, i.e., $\Pi_{B}\cap\Pi_{G}=\varnothing$.
+    for all successor states $s^{\prime}$: $h(s^{\prime})\leq V^{*}(s^{\prime})$ for an arbitrary $s$.
   2. 2.
-    Assume the policy class parameterized by $\theta$ should be expressive enough to capture optimal or near-optimal policies, and the policy is initialized randomly from uniform distribution, i.e., $\pi_{\theta_{0}}\sim\mathcal{U}(\Pi)$.
-Meanwhile, the Q-value is initialized as $Q_{\phi_{0}}(s,a)=0,\forall_{s\in S}\forall_{a\in A}$.
+    false positive at current arbitrary state $s$: $V^{*}(s)<h(s)$.
+The objective is to show that under the above conditions, we obtain $\max_{a}(\mathcal{B}h)(s,a)<h(s)$.
+  1. 1.
+    Express the Bellman backup for $h$: $(\mathcal{B}h)(s,a)=R(s,a)+\gamma\sum_{s^{\prime}}P(s^{\prime}|s,a)h(s^{\prime})$
+  2. 2.
+    Express the Bellman version of $V^{*}$: $V^{*}(s)=\max_{a\in\mathcal{A}}\left[R(s,a)+\gamma\sum_{s^{\prime}}P(s^{\prime}|s,a)V^{*}(s^{\prime})\right]$.
+    Given that $h(s^{\prime})\leq V^{*}(s^{\prime})$ for all $s^{\prime}$, we can expand it to $\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)h(s^{\prime})\leq\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)V^{*}(s^{\prime})$.
+Therefore, we have:
+    $R(s,a)+\gamma\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)h(s^{\prime})\leq R(s,a)+\gamma\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)V^{*}(s^{\prime})=Q^{*}(s,a)$ (27)
   3. 3.
-    We assume that $|\Pi_{G}|$ and $|\Pi_{B}|$ is a fixed number predefined by the task environment while $\mathbb{E}[G^{\pi}|\pi\in\Pi_{B}]$ is non-zero as false positive rewards are unavoidable in real-world VLMs.
-Since the update rule for $Q$-value is a gradient descent on $\|\mathbb{E}[G^{\pi_{\theta}}-Q_{\phi}(s,a)]\|^{2}$ and also we have that $\{\Pi_{B},\Pi_{G},\Pi\setminus(\Pi_{B}\cup\Pi_{G})\}$ is a countable partition of the policy universe $\Pi$, the updated $Q$-value will approach as follows:
-$\quad Q_{\phi}(s,a)\rightarrow\mathbb{E}[G^{\pi_{\theta}}]$
-$=P(\pi_{\theta}\in\Pi_{B})\cdot\mathbb{E}[G^{\pi_{\theta}}|\pi_{\theta}\in\Pi_{B}]$
-$\qquad+P(\pi_{\theta}\in\Pi_{G})\cdot\mathbb{E}[G^{\pi_{\theta}}|\pi_{\theta}\in\Pi_{G}]$
-$\qquad+P(\pi_{\theta}\in\Pi\setminus(\Pi_{B}\cup\Pi_{G}))$
-$\qquad\cdot\mathbb{E}[G^{\pi_{\theta}}|\pi_{\theta}\in\Pi\setminus(\Pi_{B}\cup\Pi_{G})]$
-$=P(\pi_{\theta}\in\Pi_{B})\cdot\mathbb{E}[G^{\pi_{\theta}}|\pi_{\theta}\in\Pi_{B}]$
-$\qquad+P(\pi_{\theta}\in\Pi_{G})\cdot\mathbb{E}[G^{\pi_{\theta}}|\pi_{\theta}\in\Pi_{G}]$ (6)
-Given that the update rule for policy $\pi$ is the gradient ascent on $Q_{\phi}(s,a)\pi_{\theta}(a|s)$, we have the following:
-$\quad\nabla_{\theta}Q_{\phi}(s,a)\pi_{\theta}(a|s)$
-$=(\nabla_{\theta}P(\pi_{\theta_{old}}\in\Pi_{B})\cdot\mathbb{E}[G^{\pi_{\theta_{old}}}|\pi_{\theta_{old}}\in\Pi_{B}]\pi_{\theta}(a|s))$
-$\qquad+(\nabla_{\theta}P(\pi_{\theta_{old}}\in\Pi_{G})$
-$\qquad\cdot\mathbb{E}[G^{\pi_{\theta_{old}}}|\pi_{\theta_{old}}\in\Pi_{G}]\pi_{\theta}(a|s))$
-$=P(\Pi_{B})\cdot\mathbb{E}[G^{\Pi_{B}}]\nabla_{\theta}\pi_{\theta}(a|s)$
-$\qquad+P(\pi_{G})\cdot\mathbb{E}[G^{\pi_{G}}]\nabla_{\theta}\pi_{\theta}(a|s)$
-$=(1-P(\pi_{G})-P(\pi_{\theta}\in\Pi\setminus(\Pi_{B}\cup\Pi_{G})))$
-$\qquad\cdot\mathbb{E}[G^{\Pi_{B}}]\nabla_{\theta}\pi_{\theta}(a|s)+P(\pi_{G})\cdot\mathbb{E}[G^{\pi_{G}}]\nabla_{\theta}\pi_{\theta}(a|s)$
-$=\textit{const}\cdot\mathbb{E}[G^{\Pi_{B}}]\nabla_{\theta}\pi_{\theta}(a|s)$
-$\qquad+(\mathbb{E}[G^{\pi_{G}}]-\mathbb{E}[G^{\Pi_{B}}])P(\pi_{G})\nabla_{\theta}\pi_{\theta}(a|s)$ (7)
-Justification of the assumptions.
-  • Regarding the non-zero probability of recovering the optimal policy at initialization, it is standard in theoretical analyses to assume a uniform distribution of a random variable at initialization (see Agarwal et al.
-(2021)).
-This assumption does not contradict the conclusion about the convergence rate deterioration in Theorem 3.1.
-  • In reference to the realizability condition implied by Assumption 2, the expressiveness of the policy class parameterized by $\theta$ is an underlying assumption for deep learning models, supported by the The Universal Approximation Theorem (Hornik et al.
-1989).
-Observation.
-Since the goal of the learning agent is to maximize $P(\pi\in\Pi_{G})$ (i.e., to converge to an acceptable policy), we can see that the second term provides the target direction with rate $(\mathbb{E}[G^{\pi\in\Pi_{G}}]-\mathbb{E}[G^{\pi\in\Pi_{B}}])$.
-Therefore, the ascent rate will decrease when the cumulative rewards from trajectories that cannot reach the goal state (i.e., the false positive rewards) gets higher.
-Moreover, the first term $\textit{const}\cdot\mathbb{E}[G^{\Pi_{B}}]\nabla_{\theta}\pi_{\theta}(a|s)$ can be regarded as the deviation of target direction.
-It shows that the level of deviation is also positively proportional to the magnitude of rewards coming from failed trajectories.
-This theorem follows the intuition that the presence of false positive rewards can slow down the convergence rate of the learning agent.
-A.3 Implementation Details of the Experiments
-Environment Details
+    Taking the maximum over actions to both sides, we have:
+    $\max_{a}\left(R(s,a)+\gamma\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)h(s^{\prime})\right)$ $\leq max_{a}Q^{*}(s,a)$ (28)
+    $\max_{a}(\mathcal{B}h)(s,a)$ $\leq V^{*}(s)$ (29)
+  4. 4.
+    Apply false positive condition: Given $V^{*}(s)<h(s)$, substituting into the above inequality, we have $\max_{a}(\mathcal{B}h)(s,a)<h(s)$
+Implication: Maintaining a pessimistic heuristic is inherently fragile because the introduction of a false positive in any state disrupts the pessimistic condition. ∎
+See 4.8 We analyze the convergence speed of policy learning by examining the performance gap, which is defined as the difference between the optimal value of the initial state $s_{0}$, $V^{*}(s_{0})$, and the value of the initial state under an arbitrary policy $\pi$, $V^{\pi}(s_{0})$.
+Specifically, we focus on deriving an upper bound for this performance gap.
+The key intuition is that a smaller upper bound implies faster convergence to the optimal policy, as fewer iterations of policy updates will be required to reach the optimum.
+We begin by stating the theorem made by HuRL authors:
+Theorem A.3 (Performance Gap Decomposition (Cheng et al., 2021)).
+For any policy $\pi$, heuristic $h:\mathcal{S}\rightarrow\mathbb{R}$, and mixing coefficient $\beta\in[0,1]$,
+$V^{*}(s_{0})-V^{\pi}(s_{0})=\mathrm{Regret}(h,\beta,\pi)+\mathrm{Bias}(h,\beta,\pi)$ (30)
+where the regret and the bias term are expressed as follows:
+$\mathrm{Regret}(h,\beta,\pi)$ $\coloneqq\beta\left(\widetilde{V}^{*}(s_{0})-\widetilde{V}^{\pi}(s_{0})\right)+\frac{1-\beta}{1-\gamma}\left(\widetilde{V}^{*}(d^{\pi})-\widetilde{V}^{\pi}(d^{\pi})\right)$ (31)
+$\mathrm{Bias}(h,\beta,\pi)$ $\coloneqq\left(V^{*}(s_{0})-\widetilde{V}^{*}(s_{0})\right)+\frac{\gamma(1-\beta)}{1-\gamma}\mathbb{E}_{s,a\sim d^{\pi}}\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}\left[h(s^{\prime})-\widetilde{V}^{*}(s^{\prime})\right]$ (32)
+We will not provide a proof for Theorem A.3 in this paper.
+Please refer to (Cheng et al., 2021) for details.
+The theorem demonstrates that the performance gap can be elegantly decomposed into two components: a regret term and a bias term such that:
+  1. 1.
+    The regret term quantifies the difference between $\widetilde{V^{*}}$ and $\widetilde{V^{\pi}}$, representing the error caused by $\pi$ being suboptimal in the reshaped MDP $\widetilde{\mathcal{M}}$.
+Since $\pi$ is trained by our selected RL algorithm directly on the reshaped MDP $\widetilde{\mathcal{M}}$, the primary responsibility for minimizing this regret term falls to the RL algorithm itself, not to the design of the auxiliary reward signal.
+Thus, when evaluating the effects of false positive or false negative rewards, we choose not to focus on bounding the regret term.
+  2. 2.
+    The bias term captures two key discrepancies: first, between the true optimal value function $V^{*}$ of the original MDP $\mathcal{M}$ and the optimal value function $\widetilde{V}^{*}$ of the reshaped MDP $\widetilde{\mathcal{M}}$; second, between $\widetilde{V}^{*}$ and the heuristic $h(s)$.
+This term, therefore, reflects the error introduced by addressing the reshaped MDP instead of the original one, alongside how well the heuristic $h$ approximates the optimal value function in the reshaped MDP.
+Consequently, this bias term directly relates to the quality of the heuristic reward signal $h$.
+Hence, we focus on analyzing its upper bound to assess the impact of false positive or false negative rewards on this heuristic.
+Cheng et al. (2021) have further derived the upper bound for the bias term, which we present here as a lemma. We omit the proof for brevity; for detailed proof, please refer to (Cheng et al., 2021).
+Lemma A.4 (Upper Bound of the Bias Term).
+$\mathrm{Bias}(h,\beta,\pi)\leq(1-\beta)\gamma\left(\mathbb{E}_{\rho^{\pi^{*}}}\left[\sum_{t=1}^{\infty}(\beta\gamma)^{t-1}(V^{*}(s_{t})-h(s_{t}))\right]+\mathbb{E}_{\rho^{\pi}}\left[\sum_{t=1}^{\infty}\gamma^{t-1}(h(s_{t})-\widetilde{V}^{*}(s_{t}))\right]\right)$ (33)
+where $\rho^{\pi}$ denotes the trajectory distribution of $s_{0},a_{0},s_{1},...$ induced by running $\pi$ starting from $s_{0}$.
+This upper bound elegantly illustrates a trade-off: if the heuristic $h$ is set too high (overestimation error), it reduces the first term but increases the second term.
+Conversely, if $h$ is set too low (underestimation error), it increases the first term but reduces the second term.
+Just by inspection, it is not immediately clear whether overestimation or underestimation results in a better upper bound, as both impact the bias term in opposing ways.
+We now prove that underestimation of $h$ (pessimistic $h$) is better than overestimation (i.e., $\exists s_{t}\in\mathcal{S},h(s_{t})>V^{*}(s_{t})$) for minimizing bias.
+Proof.
+  1. 1.
+    Define Terms:
+    • Let $B_{1}=\mathbb{E}_{\rho^{\pi^{*}}}\left[\sum_{t=1}^{\infty}(\beta\gamma)^{t-1}(V^{*}(s_{t})-h(s_{t}))\right]$
+    • Let $B_{2}=\mathbb{E}_{\rho^{\pi}}\left[\sum_{t=1}^{\infty}\gamma^{t-1}(h(s_{t})-\widetilde{V}^{*}(s_{t}))\right]$
+  2. 2.
+    Underestimation Case (Pessimistic $h$):
+    Lemma A.5.
+    If $h$ is pessimistic with respect to $\mathcal{M}$, $\forall\beta\in[0,1],s\in\mathcal{S},\widetilde{V}^{*}(s)\geq h(s)$
+    Proof.
+    To begin with, we need to use another lemma from Cheng et al. (2021), shown as follows:
+    Lemma A.6 (Bellman backup of reshaped MDP (Cheng et al., 2021)).
+    For any policy $\pi$, we have
+    $\widetilde{V}^{\pi}(s_{0})-h(s_{0})=\frac{1}{1-\beta\gamma}\mathbb{E}_{\widetilde{d}_{s_{0}}^{\pi}}[(\widetilde{\mathcal{B}}h)(s,a)-h(s)]$ (34)
+    where $\widetilde{d}_{s_{0}}^{\pi}$ refers to the discounted average state distribution of policy $\pi$ in reshaped MDP $\widetilde{\mathcal{M}}$.
+It can be expressed as $\widetilde{d}_{s_{0}}^{\pi}=(1-\widetilde{\gamma})\sum_{t=0}^{\infty}\widetilde{\gamma}^{t}d_{t}^{\pi}$, and $d_{t}^{\pi}$ is the state distribution of policy $\pi$ at time $t$ with $d_{0}^{\pi}=\mathbf{1}\{s=s_{0}\}$.
+    For brevity, we do not include the proof for Lemma A.6.
+    First of all, due to the definition of optimal value $V^{*}$, we have
+    $\widetilde{V}^{*}(s_{0})\geq\widetilde{V}^{\pi}(s_{0})$ (35)
+    Then, according to Lemma A.6, we can get
+    $\widetilde{V}^{*}(s_{0})\geq\widetilde{V}^{\pi}(s_{0})=h(s_{0})+\frac{1}{1-\beta\gamma}\mathbb{E}_{\widetilde{d}_{s_{0}}^{\pi}}[(\widetilde{\mathcal{B}}h)(s,a)-h(s)]$ (36)
+    Let $\pi$ denote the greedy policy of $\arg\max_{a}(\mathcal{B}h)(s,a)$ and then trace back to Equation 8, we have (~$(\widetilde{\mathcal{B}}h)(s,a)=(\mathcal{B}h)(s,a)$, that means
+    $\widetilde{V}^{*}(s_{0})\geq\widetilde{V}^{\pi}(s_{0})$ $=h(s_{0})+\frac{1}{1-\beta\gamma}\mathbb{E}_{\widetilde{d}_{s_{0}}^{\pi}}[(\widetilde{\mathcal{B}}h)(s,a)-h(s)]$ (37)
+    $=h(s_{0})+\frac{1}{1-\beta\gamma}\mathbb{E}_{\widetilde{d}_{s_{0}}^{\pi}}[(\mathcal{B}h)(s,a)-h(s)]$ (38)
+    $\geq h(s_{0})\quad\text{[direct result of $h$ being pessimistic]}$] (39)
+    Generalization to Any State: The lemma from Cheng et al.
+(2021) (Lemma A.6) is stated in terms of any starting state $s_{0}$.
+Therefore, we can replace $s_{0}$ with any state $s\in\mathcal{S}$ in our analysis.
+Thus we get $\forall s\in\mathcal{S},\widetilde{V}^{*}(s)\geq h(s)$
+    Lemma A.5 implies that when $h$ is pessimistic, $B_{2}=\mathbb{E}_{\rho^{\pi}}\left[\sum_{t=1}^{\infty}\gamma^{t-1}(h(s_{t})-\widetilde{V}^{*}(s_{t}))\right]\leq 0$ because $\widetilde{V}^{*}(s)\geq h(s)$ for all $s$.
+Therefore, the bias error is only bounded by $B_{1}$ when $h$ is pessimistic, i.e., $\mathrm{Bias}(h,\beta,\pi)\leq B_{1}$.
+  3. 3.
+    Overestimation Case ($\exists s_{t}\in\mathcal{S},h(s_{t})>V^{*}(s_{t})$):
+    We show briefly that, unlike the underestimation error, the overestimation error does not have a closed-form upper bound expression.
+The difficulty, as pointed out by (Cheng et al., 2021), originates from the trajectory-dependence on $\mathbb{E}_{\rho^{\pi}}[\cdot]$ as the $l_{\infty}$ approximation error here can be difficult to control in large state spaces.
+It is possible that, upon picking up falsely high $h$ states, $\rho^{\pi}$ gets further distorted away from $\rho^{\pi^{*}}$ and accumulates more falsely high $h$ states.
+In other words, this trajectory dependence makes $B_{2}$ prone to a feedback loop of accumulating overestimation errors and resulting in a much larger upper bound for the bias.
+A.3 Additional Notes on Theoretical Setting
+A.3.1 Convergence Guarantee of HuRL Compared With Potential-Based Reward Shaping
+Previous work on potential-based reward shaping (PBRS) has established that learning a policy on a reshaped MDP can converge to the optimum of the original MDP, as proved by Ng et al.
+(1999).
+They proved that the optimal $Q$ value of the reshaped MDP is equivalent to the optimal $Q$ value of the original MDP minus a state-dependent function.
+Consequently, for an optimal policy defined as $\pi^{*}=\arg\max_{a\in\mathcal{A}}Q^{*}(s,a)$, this additional state-dependent value does not alter the policy.
+Thus, the optimal policy of the original MDP $\mathcal{M}$ is also the optimal policy for the reshaped $\widetilde{\mathcal{M}}$.
+However, the theoretical convergence for the reshaped MDP within the framework of Heuristic-Guided Reinforcement Learning (HuRL) has not been proven with the same rigor.
+The authors of HuRL made a trick which involves manipulating the coefficient $\beta$, which scales the heuristic-based auxiliary rewards.
+If $\beta$ increases gradually from 0 to 1 over the training process, the agent effectively transitions from interacting with the modified MDP back to the original MDP.
+This method ensures that, at the end of training, the agent is exactly solving the original problem.
+Moreover, according to the Blackwell optimal property (Chen et al., 2018), convergence can occur before $\beta$ reaches 1, thereby allowing the trained policy to maintain optimality in the original MDP under HuRL.
+However, in the original HuRL paper, among the five test environments, the $\beta$ hyperparameter was actually fixed for four of them.
+Surprisingly, HuRL still converged to the optimal policy faster than the original MDP without updating $\beta$.
+This observation poses an enigma within the HuRL framework, as it is unclear why this convergence to the optimum of the original MDP occurs without dynamically adjusting $\beta$.
+Thus, while the practical success of HuRL in these environments is evident, the underlying reasons for such convergence remain to be fully understood.
+This discrepancy between theoretical expectation and empirical results leaves room for future research to explore why HuRL converges under fixed $\beta$.
+Nevertheless, proving the convergence properties goes beyond the scope of our paper, which aims to highlight the prevalence of false positive rewards and their impact.
+An alternative perspective on this issue is to consider the learning process of maximizing auxiliary rewards as akin to performing behavior cloning from experts.
+In this sense, we do not need to focus on convergence within the current MDP but can view it as a form of pretraining.
+A.4 Implementation Details of the Experiments
+A.4.1 Environment Details
 We describe each testing environment used in our experiments.
-More details introduction can be found in on the official project homepage of each benchmark (Chevalier-Boisvert et al.
-2018; Bellemare et al.
-2013; Hafner 2021).
+More details introduction can be found in on the official project homepage of each benchmark (Chevalier-Boisvert et al., 2018; Bellemare et al., 2013; Hafner, 2021).
   • Crafter features randomly generated 2D worlds where the player needs to forage for food and water, find shelter to sleep, defend against monsters, collect materials, and build tools.
@@
 Figure 10: Illustration of the Crafter task. The agent must survive as long as possible and explore for new crafting recipes.
-Instruction-Guide Procedure Details
+A.4.2 Instruction-Guide Procedure Details
 The VLM-based reward model will have a pointer to the sequence of the instruction sentence, starting at the first sentence.
-For original models Pixl2R and ELLM-, we follow the setting in there original work where for each instruction sentence, the reward model will have a maximum cap of rewards (2.0) it can assign to the agent in one episode.
+For original models Pixl2R and ELLM-, we follow the setting in there original work where for each instruction sentence (yes, the full instruction essay will be split into multiple sentences and treat each sentence as atomic instruction $l_{k}$), the reward model will have a maximum cap of rewards (2.0) it can assign to the agent in one episode.
 When the cap is reached, the reward model will move its pointer to the next instruction sentence.
 For the BiMI reward model, the reward model will move its pointer to the next instruction sentence when the binary signal is triggered.
-Finetuning VLM-based Reward Models
-In contrast to previous work on Instruction-guided RL where they rely on hand-crafted oracle multimodal reward models, we use actual pretrained VLMs to generate reward signals.
-2 VLM backbone models are used in our experiments: 1) CLIP (Radford et al.
-2021), pretrained by image-text pairs; and (2) X-CLIP (Ma et al.
-2022), pretrained by video-text pairs.
+Below is the pseudo-code for the instruction-following RL training procedure for both Pixl2R and ELLM- models.
+Algorithm 1 ELLM and Pixl2R pseudo-code for instruction-following RL
+1:Initialize policy network $\pi_{\theta}$
+2:Initialize value network $V_{\phi}$
+3:Setup VLM-based reward model $E(\cdot)$
+4:Split instruction essay into sentences $\{l_{1},l_{2},...,l_{K}\}$
+5:Initialize instruction pointer $p=1$
+6:Initialize cumulative VLM reward $r_{\text{cum}}=0$
+7:Initialize cumulative VLM reward threshold $q$
+8:Initialize replay buffer $\mathcal{D}$
+9:Initialize agent trajectory memory queue $\tau$ with length $W$Foreach episode
+[…]
+11:Select action $a_{t}\sim\pi_{\theta}(a_{t}|s_{t})$
+12:Execute $a_{t}$, observe next state $s_{t+1}$ and extrinsic reward $r^{e}_{t}$
+13:Enqueue $(s_{t},a_{t},r_{t},s_{t+1})$ in $\tau$
+14:Compute VLM reward:
+15:$r^{v}_{t}=\frac{E(\tau)\cdot E(l_{p})}{\|E(\tau)\|\|E(l_{p})\|}$triangleright Apply BiMI here if needed
+16:Combine rewards: $r_{t}=r^{e}_{t}+(1-\beta)\gamma r^{v}_{t}$triangleright $\beta$ is a scaling factor
+17:Store $(s_{t},a_{t},r_{t},s_{t+1})$ in $\mathcal{D}$
+[…]
+19:$p\leftarrow\min(p+1,K)$
+[…]
+21:Sample mini-batch $\{(s_{j},a_{j},r_{j},s_{j+1})\}$ from $\mathcal{D}$
+22:Compute TD errors:
+23:$\delta_{j}=r_{j}+\gamma V_{\phi}(s_{j+1})-V_{\phi}(s_{j})$
+24:Update value network:
+25:$\phi\leftarrow\phi+\alpha_{v}\sum_{j}\delta_{j}\nabla_{\phi}V_{\phi}(s_{j})$
+26:Update policy network:
+[…]
+A.4.3 Finetuning VLM-based Reward Models
+In contrast to previous work on instructing following RL where they rely on hand-crafted oracle multimodal reward models, we use actual pretrained VLMs to generate reward signals.
+2 VLM backbone models are used in our experiments: 1) CLIP (Radford et al., 2021), pretrained by image-text pairs; and (2) X-CLIP (Ma et al., 2022), pretrained by video-text pairs.
 In particular, Pixl2R uses CLIP because it only uses the single latest frame as input.
@@
 Montezuma 0.8044 0.9259 0.8045 0.9657 XCLIP ELLM-
-Hyperparameters for Instruction-Guided RL Agents
+A.4.4 Hyperparameters for Instruction-Following RL Agents
 In the experiments, all methods are implemented based on PPO with same model architecture.
-The Minigrid and Crafter environments use the same training hyperparameters as the Achievement Distillation paper (Moon et al.
-2023).
+The Minigrid and Crafter environments use the same training hyperparameters as the Achievement Distillation paper (Moon et al., 2023).
 For Montezuma’s Revenge, we found that the performance of the agent was sensitive to the gamma and GAE lambda parameters.
@@
 - lr 1.0e-4
-A.4 Additional Details of the Experiments of False Positive Rewards
+A.5 Additional Details of the Experiments on the Impact of Noisy Rewards
 Evaluation Metric Details In our experiments, we used a score metric adapted from the Crafter paper to evaluate agent performance across different environments.
 This score metric aggregates success rates for individual subtasks using a geometric mean.
+Formally, the score metric is defined as follows:
+$\text{Score}=\exp\left(\frac{1}{N}\sum_{k=1}^{N}\ln(1+s_{k})\right)-1$ (40)
+where $s_{k}$ is the agent’s success rate of achieving instruction $l_{k}$, and $N$ is the total number of instructions.
 This metric was chosen over the maximum total rewards metric for several reasons:
@@
 For Minigrid and Montezuma environments, we use the internal information from the game engine to detect whether the subtasks are completed, thus facilitating the calculation of the score metric.
-Details of Manipulated Trajectory-Instruction Pairs to Evaluate Robustness
+A.5.1 Details of Manipulated Trajectory-Instruction Pairs to Evaluate Robustness
 We evaluated the models’ sensitivity by examining how cosine similarity scores change for manipulated trajectory-instruction pairs.
@@
   2. 2.
-    Instruction Negation: We modified the original instructions by adding negation (e.g., changing “do $l_{k}$” to “do not do $l_{k}$”).
+    Instruction Negation: We modified the original instructions by adding negation (e.g., changing “do $l_{k}$” to “do not do $l_{k}$” or “avoid $l_{k}$”).
 This tests the model’s sensitivity to semantic changes in the instruction that fundamentally alter the goal.
@@
     This assesses the model’s ability to detect partial mismatches in longer sequences.
-A.5 Pseudo-code for Empirical Quantile Calculation for Binary Signal Threshold
-Using empirical quantile as threshold guarantees a high probability (at least $1-\alpha$) that the true positive pairs are recognized while minimizing the average number of mistakes predicting false positives (Sadinle et al.
-2019):
+A.6 Details of Showing the Prevalence of False Positives in VLM Cosine Similarity Scores
+We list the figures of reward signals from learned VLMs for different types of trajectory-instruction pairs for each individual environment.
+All figures show that the VLM-based reward models assign high rewards to manipulated trajectory-instruction pairs, indicating the prevalence of false positive rewards.
+Figure 12: Cosine similarity scores for match, mismatch and manipulated trajectory-instruction pairs in Montezuma.
+Figure 13: Cosine similarity scores for match, mismatch and manipulated trajectory-instruction pairs in Minigrid.
+Figure 14: Cosine similarity scores for match, mismatch and manipulated trajectory-instruction pairs in Crafter.
+A.7 Impact of Reward Model Quality on Policy Learning Efficiency: Supplementary Experiments and Quantitative Analysis in Montezuma
+In our efforts to assess the impact of false positive rewards from auxiliary reward model without the interference of other factors such as domain shift, poor data quality, and errors from other issues such as the choice of multimodal architectures, we devised a simulated auxiliary reward model (also known as oracle auxiliary reward model) that access to internal state information from the game engine.
+The model compares the sequence of past actions and states of the agent with predefined target intermediate state sets that map to each instruction sentence.
+This is feasible in Montezuma’s Revenge environment as we are able to access coordinate system information directly from the game engine.
+This access allows us to locate the current positional information of the agent and also label specific intermediate states as targets and assign rewards to the agent accordingly.
+We therefore designed three types of simulated reward model:
+  • Sim RM 1 (Perfect) generates rewards accurately whenever the agent reaches the designated intermediate states.
+Furthermore, it strictly adheres to the chronological sequence of instructions; rewards for subsequent instructions are only awarded if all preceding instructions have been fulfilled.
+  • Sim RM 2 (False Positive) introduces a tolerance for false positive rewards but with reduced reward magnitudes, all while maintaining the temporal sequence.
+This is implemented by defining a radius $\sigma$, where if the agent enters a circle centered at the target state with a radius of $\sigma$, it receives a small amount of rewards.
+  • Sim RM 3 (Temporal Insensitive False Positive) disregards the chronological order of instructions, allowing rewards for later tasks even if earlier ones remain unfulfilled.
+However, note that fulfilling every sub-task will still result in the agent receiving the maximum total rewards.
+Therefore, in theory, the policy will eventually converge.
+Results are reported in Table 7. Several important observations are as follows:
+Table 7: Agent performance in Montezuma’s Revenge, evaluated across three rooms with the goal state being the exit through a designated door.
+Metrics measured are the Area Under the Curve (AUC) for total reward (where higher is better) and success rate (SR) of reaching the goal state.
+The baseline is denoted as B, with $\star$ marking statistical significance ($p<0.05$).
+{tblr}
+width = 0.95colspec = Q[473]Q[346]Q[106],row1 = p,hline1-2,8 = -,hline4,6 = -dashed Model & AUC SR
+(1) PPO (Schulman et al., 2017) all failed 0%
+(2) PPO+RND (B) (Burda et al., 2018) 0.550$\pm$0.066 100%
+(3) PPO + Sim RM 1 (Perfect) 0.287$\pm$0.048 100%
+(4) PPO+RND + Sim RM 1 (Perfect) 0.608$\pm$0.073 100%
+(5) PPO+RND + Sim RM 2 (False Positive) 0.183$\pm$0.187 $\star$ 73.3%
+(6) PPO+RND + Sim RM 3 (Temp. Insen.) 0.051$\pm$0.116 $\star$ 16.7%
+  • Perfect auxiliary reward model have shown enhanced performance compared to weak RL baselines like PPO.
+While agents trained solely with PPO struggled to play Montezuma, incorporating Perfect auxiliary reward model into PPO (i.e., (3) in Table 7) did find the goal state, with the success rate increasing from 0% to 100%.
+However, we observed that the auxiliary reward model learns more slowly than the intrinsic reward model (see (2) in Table 7) when attempting to reach the goal state.
+This finding was initially surprising, but upon examining the agent movement heatmaps in Figure 15(a) and Figure 15(b), an explanation becomes clear: the PPO+RND model discovers shorter paths to the goal state.
+The heatmap reveals that the PPO+RND agent learns to directly jump to a rope, bypassing the use of a ladder and conveyor belt as suggested by the expert instructions.
+This observation raises a concern about instruction-following based reward signals; they might limit the agent’s exploration, confining it to states favored by the expert.
+If the quality of the expert’s knowledge about the task is not optimal, this can result in the agent failing to compete with an exploration-based reward model.
+  • Again, we observed a remarkable synergy between the instruction-following-based reward model and the intrinsic reward model, as discussed in §7.1.
+As shown in (4) in Table 7, the PPO+RND + Sim RM 1 (Perfect) agent achieves the highest AUC score, demonstrating that it learns to reach the goal state faster than any other model.
+  • The involvement of false positive rewards significantly slows down the learning process of the model, as evidenced by entries (5) and (6) in Table 7.
+Additionally, both (5) and (6) show lower AUC scores compared to (2), which is the vanilla PPO+RND model.
+If we consider the vanilla PPO+RND model to represent a scenario with full false negative instruction-following rewards (since no instruction-following reward model is implemented here), then comparing the scores of (2) with (5) and (6) further indicates that false positive rewards are more detrimental than false negative ones.
+  • The RL + VLM reward approach completely failed to converge when the reward model disregarded temporal ordering, as evidenced by entry (6) in Table 7.
+The success rate had a significant drop to 16.7%, and even the RND intrinsic reward was unable to recover the success rate within the time budget.
+This illustrates the severe impact of false positive rewards due to temporal ordering insensitivity, an example of composition insensitivity discussed in §5; they trap the agent in a cycle of ineffective actions.
+However, note that if every sub-task is fulfilled, the agent will still receive the maximum total rewards, suggesting that, in theory, the policy could eventually converge.
+Yet, the practical implications of such false positives emphasize how they can drastically extend the learning period or lead to non-convergence in real-world scenarios where time or computational resources are limited.
+(a) PPO+RND agent (same as + Sim RM (False Negative))
+(b) PPO+RND + Sim RM 1 (Perfect)
+(c) PPO+RND + Sim RM 2 (False Positive)
+(d) PPO+RND + Sim RM 3 (Temporal Insensitivity)
+Figure 15: Movement heatmap for PPO+RND agents when different simulated auxiliary instruction-following-based reward models are involved.
+We also provided movement heatmap of different models for further qualitative analysis.
+As shown in Figure 15(a), intrinsic reward model explored a shortcut towards the goal state by directly jump from the initial state to the rope, thereby learning to reach the goal state faster than other models.
+Figures 15(b) and 15(c) do not exhibit a significant contrast, but they do reveal that the presence of false positive rewards leads the agent to more frequently visit dead ends, such as falling off cliffs.
+This behavior is consistent with the mechanics of the simulated false positive reward model, which defines a radius $\sigma$ around target states.
+Within this radius, even if an agent falls off a cliff, it might receive rewards because its position could be close enough to the intended intermediate rewarding state.
+Consequently, agents had a chance of mistakenly perceiving falling off as advantageous.
+Figure 15(d) illustrates the consequences when temporal ordering restrictions are removed: the learning agent persistently pursues rewards associated with the final instruction, “walk left to the door.” However, simply executing this last instruction does not suffice for escaping the room because access to the door requires a key.
+This suggests that false positive rewards cause the agent to get trapped in local minima closer to $s_{0}$, indicating that deep RL can easily hack the reward system and effectively exploit shortcuts (Clark, 2016).
+A critical issue arises here: even with intrinsic rewards, the agent fails to escape these local optima, suggesting it has reached an equilibrium in the MDP where policy updates cease.
+This scenario is particularly relevant when the initial state distribution does not cover the entire state space.
+If the initial distribution does not have a positive probability for all states (i.e., $foralls\in\mathcal{S},d_{0}(s)>0$), it is not guaranteed that the RL algorithm will eventually reach the global optimum with maximum total rewards, as highlighted by (Agarwal et al., 2021).
+A.8 Pseudo-code for Empirical Quantile Calculation for Binary Signal Threshold
+Using empirical quantile as threshold guarantees a high probability (at least $1-\alpha$) that the true positive pairs are recognized while minimizing the average number of mistakes predicting false positives (Sadinle et al., 2019):
 Algorithm 2 Calculate Empirical Quantile ($\hat{q}$)
@@
 […]
-A.6 Details of the Experiments of BiMI Reward Function
+A.9 Implementation Details of the Experiments of BiMI Reward Function
 We set confidence level for empirical quantile calculation to be $0.91-\alpha=$ 0.9.
-We adhered to the standard requirement of limiting the training budget to 1 million frames (Hafner 2021).
+We adhered to the standard requirement of limiting the training budget to 1 million frames (Hafner, 2021).
 This constraint poses a significant challenge, particularly in sparse reward settings, as it demands that agents both explore efficiently and exploit their knowledge effectively within this limited budget.
@@
 The implementation details for the BiMI reward function are consistent with those outlined in the first stage of experiments.
-A.7 Cosine Similarity Score Evaluation for Different Environments
-Figure 14: Cosine similarity scores over matched, not-matched and manipulated trajectory-instruction pairs in Montezuma’s Revenge.
-Figure 15: Cosine similarity scores over matched, not-matched and manipulated trajectory-instruction pairs in Minigrid.
-Figure 16: Cosine similarity scores over matched, not-matched and manipulated trajectory-instruction pairs in Crafter.
-A.8 Line Plots over Average Reward for Different Environments and Ablation Study
-Figure 17: The average reward of agents across different environments. The BiMI reward function significantly improves the average reward of agents in all environments
-Figure 18: Ablation on the components of BiMI reward function and their impact on the average reward of agents.
+A.10 Detailed Experiment Results of BiMI Reward Function
+Figure 16: Besides the improvements of the score performance of agents across different environments with the BiMI reward function, it also collaborates well with intrinsic rewards.
+Combining both can lead to significant performance improvements
+Figure 17: Ablation on the components of BiMI reward function.
+The binary reward (Bi) alone led to a 36.5% improvement compared to original models.
+Excluding Crafter, Mutual Information (MI) provided a 23% further improvement over Bi alone
+A.10.1 Minigrid
+ELLM-$+$BiMI achieved a remarkable 74% improvement in performance compared to the original models.This substantial gain is particularly noteworthy given the unique challenges presented by Minigrid.
+The abstract, shape-based visuals in Minigrid diverge drastically from the natural images used in VLMs’ pretraining, preventing the models from effectively utilizing their prior pretraining knowledge.
+Consequently, VLMs struggled to accurately assess similarities between Minigrid’s abstract visuals and textual instructions, resulting in highly noisy reward signals.
+The significant improvement demonstrated by BiMI underscores its effectiveness in handling noisy signals, directly addressing our primary research challenge.
+This capability is crucial for deploying instruction-following agents in real-world, unfamiliar scenarios, where visual inputs often deviate from the VLMs’ training distribution, leading to noisy reward signals.
+A.10.2 Crafter
+We observed an intriguing pattern of results.
+The Bi component alone led to 14% and 3.2% improvement in performance over the original models.
+However, contrary to our observations in other environments, the addition of the MI component actually decreased this improvement.
+This unexpected outcome can be attributed to the unique nature of Crafter task, where agents must repeatedly achieve the same subtasks (e.g., drinking water) for survival.
+The MI component, designed to discourage over-reliance on frequently occurring signals, inadvertently penalized the necessary repetition of survival-critical actions.
+Furthermore, note that instruction-following RL agents, regardless of the reward model employed, were unable to outperform pure RL agents in Crafter.
+This discrepancy is due to the open-world nature of Crafter, which requires dynamic strategies and real-time decision-making that our testing setups did not fully capture.
+Despite these challenges, it is noteworthy that Bi alone still managed to improve performance over vanilla VLM-based reward models, suggesting that reducing false positives is still beneficial across all testing environments.
+The combination of BiMI with DEIR (the intrinsic reward model) also showed promising results, indicating a productive balance between exploration (driven by DEIR) and exploitation (guided by BiMI instruction reward).
+A.11 DEPRECATED Proof of the Reduction of Convergence Rate
+WARNING: The following content is deprecated.
+This section contains the original convergence analysis for the ARR October Version, which has been identified as problematic by Reviewer nSDu.
+It is retained here solely for reference and backtracking purposes.
+Formally, Abel et al. (2021) defined that:
+Theorem A.7 (Abel et al. (2021)).
+A reward function realizes a Range Set of Acceptable Policies (Range-SOAP) $\Pi_{G}$ when there exists an $\epsilon\geq 0$ such that every $\pi_{g}\in\Pi_{G}$ is $\epsilon$-optimal in start-state value, $V^{*}(s_{0})-V^{\pi_{g}}(s_{0})\leq\epsilon$, while all other policies are worse.
+When the reward signal is sparse, the agent only receives a reward upon reaching some goal states, and the reward function does not provide any feedback during the intermediate steps.
+We argue that the sparse reward function realizes a Range-SOAP, leading to the categorization of policies into acceptable and unacceptable policies.
+Proposition A.8.
+Sparse reward function “realizes” Range-SOAP (i.e., Range Set of Acceptable Policies).
+Justification:
+A sparse reward function, where the agent only receives a reward upon completing the task, cannot prefer policies that lead to shorter task completion times.
+This is because either the agent completes the goal very quickly or slowly, they will receive nearly the same amount of cumulative rewards, and the reward function will not show a strong preference.
+Since the sparse reward function does not induce a strict partial ordering on the policies, we say this reward function cannot realize a Partial Ordering (PO) task.
+Specifically, a PO on policies is a generalization of a Set of Acceptable Policies (SOAP) task.
+In a PO, the agent specifies a partial ordering on the policy space, where some policies are identified as “great”, some as “good”, and some as “bad” to strictly avoid, while remaining indifferent to the rest.
+Therefore, the sparse reward function can realize a Set of Acceptable Policies (SOAP), where there is a set of policies that are all considered “good” or near-optimal, while all other policies are worse.
+Furthermore, the sparse reward function will lead to a Range-SOAP, rather than an Equal-SOAP.
+Specifically, Equal-SOAP is a SOAP where all the acceptable policies are equally optimal in start-state value.
+This is because the good policies in the SOAP may differ slightly in their start-state values, as some may reach multiple goal states in the environment and thereby receiving different cumulative rewards.
+Therefore, the sparse reward function will realize a Range-SOAP, where there is a range of acceptable policies that are all near-optimal in start-state value.
+This proposition forms the basis of our theoretical analysis in Section A.11.
+Below is the convergence analysis when the total return can be classified by returns of trajectories from acceptable policies $P(\tau\in\mathcal{T}_{G}\mid\theta)\cdot\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{G}]$ versus returns of trajectories from unacceptable policies $P(\tau\in\mathcal{T}_{B}\mid\theta)\cdot\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{B}]$, where $\mathcal{T}_{G}\sim\Pi_{G}$, $\mathcal{T}_{B}\sim\Pi_{B}$ and $\Pi_{B}=\Pi\setminus\Pi_{G}$.
+Note on Derivation: The proof presented here can be viewed as a reformulation of the Q-value function’s gradient in the context of our specific problem setup.
+While this derivation may appear straightforward to readers well-versed in reinforcement learning, we include it to provide a clear mathematical foundation for our analysis of false positive rewards in the VLM-RL context.
+This formulation helps bridge the gap between standard RL theory and our specific problem domain on instruction-following RL with VLM-based rewards.
+Specifically, the update rule of Actor-Critic algorithm is:
+  • Critic:
+    $\phi\leftarrow\phi-\alpha_{\phi}\nabla_{\phi}(\delta)^{2}$ (41)
+    where
+    $\delta=\mathbb{E}_{\pi_{\theta}}[G^{\pi_{\theta}}-Q_{\phi}(s,a)]$ is the Monte Carlo (MC) estimation error, $Q_{\phi}(s,a)$ is the $Q$-value function which measures the expected discounted cumulative reward given the state $s$ and action $a$, and $\pi_{\theta}$ is the policy.
+    $G^{\pi_{\theta}}$ is the rollout cumulative rewards from the trajectory $\tau^{\pi_{\theta}}$ generated from $\pi_{\theta}$.
+  • Actor:
+    $\theta\leftarrow\theta+\alpha_{\theta}\frac{Q_{\phi}(s,a)\nabla_{\theta}\pi_{\theta}(a|s)}{\pi_{\theta}(a|s)}$ (42)
+We need to make the following assumptions to simplify the theoretical analysis:
+  1. 1.
+    We use $\mathcal{T}_{G}$ to represent the set of acceptable trajectories and use $\mathcal{T}_{B}$ as a non-overlapping set of “not acceptable” trajectories, we can say:
+    $\mathcal{T}_{G}\cap\mathcal{T}_{B}=\emptyset$ (43)
+    where $\mathcal{T}$ is the trajectory distribution space of the agent policy $\pi_{\theta}$.
+  2. 2.
+    When we define $\mathcal{T}_{B}$ and $\mathcal{T}_{G}$ within the universe $\mathcal{T}$, we must account for the possibility of a remaining set to ensure a complete partition of $\mathcal{T}$.
+The remaining set $\mathcal{T}\setminus(\mathcal{T}_{B}\cup\mathcal{T}_{G})\}$ can be interpreted as a trajectories that partially fulfill the instruction.
+Given our focus on the influence of false positive rewards within challenging sequential decision-making tasks, we assume that partially correct trajectories are also unacceptable.
+This stance is justified by the fact that partially correct trajectories can be viewed as false positives: they appear to be plausible behaviors but ultimately fail to fulfill the instructions.
+Rewarding such trajectories could reinforce incomplete or potentially dangerous behaviors.
+The errors discussed in Section 5 can be seen as instances of partially correct trajectories that are ultimately unacceptable.
+For example, incompletely following a safety-critical task such as “turn off the machinery before performing maintenance” could pose significant risks, and rewarding such behavior would be undesirable.
+Consequently, for the purposes of our analysis, we can assume:
+    $\mathcal{T}_{G}\cup\mathcal{T}_{B}=\mathcal{T}$ (44)
+  3. 3.
+    Assume the policy class parameterized by $\theta$ should be expressive enough to capture optimal or near-optimal policies, and the policy is initialized randomly from uniform distribution, i.e., $\pi_{\theta_{0}}\sim\mathcal{U}(\Pi)$.
+Meanwhile, the Q-value is initialized as $Q_{\phi_{0}}(s,a)=0,\forall_{s\in S}\forall_{a\in A}$.
+  4. 4.
+    We assume that $|\mathcal{T}_{G}|$ and $|\mathcal{T}_{B}|$ is a fixed number predefined by the task environment while $\mathbb{E}_{\tau\in\pi_{\theta}}[G(\tau)\mid\tau\in\mathcal{T}_{B}]$ is non-zero as false positive rewards are unavoidable in real-world VLMs.
+Since the update rule for $Q$-value is a gradient descent on $\|\mathbb{E}_{\tau\in\pi_{\theta}}[G(\tau)-Q_{\phi}(s,a)]\|^{2}$, the updated $Q$-value will approach as follows:
+$Q_{\phi}(s,a)\rightarrow\mathbb{E}[G(\tau)]$ (45)
+$=$ $P(\tau\in\mathcal{T}_{B}\mid\theta)\cdot\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{B}]+P(\tau\in\mathcal{T}_{G}\mid\theta)\cdot\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{G}]$
+$+P(\tau\in\mathcal{T}\setminus(\mathcal{T}_{B}\cup\mathcal{T}_{G})\mid\theta)\cdot\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}\setminus(\mathcal{T}_{B}\cup\mathcal{T}_{G})]$ (46)
+[…]
+Given that the update rule for policy $\pi$ is the gradient ascent on $Q_{\phi}(s,a)\pi_{\theta}(a\mid s)$, we have the following:
+$\nabla_{\theta}Q_{\phi}(s,a)\pi_{\theta}(a\mid s)$ (49)
+$=$ $\big{(}\nabla_{\theta}P(\tau\in\mathcal{T}_{B}\mid\theta_{\text{old}})\cdot\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{B}]\pi_{\theta}(a\mid s)\big{)}$
+$+\big{(}\nabla_{\theta}P(\tau\in\mathcal{T}_{G}\mid\theta_{\text{old}})\cdot\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{G}]\pi_{\theta}(a\mid s)\big{)}$ (50)
+$=$ $P(\tau\in\mathcal{T}_{B}\mid\theta_{\text{old}})\cdot\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{B}]\nabla_{\theta}\pi_{\theta}(a\mid s)$
+$+P(\tau\in\mathcal{T}_{G}\mid\theta_{\text{old}})\cdot\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{G}]\nabla_{\theta}\pi_{\theta}(a\mid s)\quad\text{$\theta_{\text{old}}$ terms are constants w.r.t.
+$\theta$}$ (51)
+$=$ $\big{(}1-P(\tau\in\mathcal{T}_{G}\mid\theta_{\text{old}})\big{)}\cdot\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{B}]\nabla_{\theta}\pi_{\theta}(a\mid s)$
+$+P(\tau\in\mathcal{T}_{G}\mid\theta_{\text{old}})\cdot\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{G}]\nabla_{\theta}\pi_{\theta}(a\mid s)$ (52)
+$=$ $\textit{const}\cdot\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{B}]\nabla_{\theta}\pi_{\theta}(a\mid s)$
+[…]
+Note the highlighted term in green ($P(\tau\in\mathcal{T}_{G}\mid\theta_{\text{old}})$) is the direction of the gradient ascent on parameter $\theta$.
+Justification of the assumptions.
+  • Regarding the non-zero probability of recovering the optimal policy at initialization, it is standard in theoretical analyses to assume a uniform distribution of a random variable at initialization (see Agarwal et al.
+(2021)).
+  • In reference to the realizability condition implied by Assumption 3, the expressiveness of the policy class parameterized by $\theta$ is an underlying assumption for deep learning models, supported by the The Universal Approximation Theorem (Hornik et al., 1989).
+Observation.
+Since the goal of the learning agent is to maximize $P(\tau\in\mathcal{T}_{G}\mid\theta)$ (i.e., to converge the agent policy to the distribution of acceptable trajectories), we can see that the second term provides the target direction with rate $(\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{G}]-\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{B}])$.
+Therefore, the ascent rate will decrease when the cumulative rewards from unacceptable trajectories (i.e., the false positive rewards) gets higher.
+In addition, the first term $\textit{const}\cdot\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{B}]\nabla_{\theta}\pi_{\theta}(a\mid s)$ can be regarded as the deviation term of target direction.
+This formalization follows the intuition that the presence of false positive rewards can slow down the convergence rate of the learning agent.

### v5

@@
 BiMI significantly enhances learning efficiency across diverse and challenging embodied navigation environments.
-Our findings offer a nuanced understanding of how different types of reward noise impact agent learning and highlight the importance of addressing multimodal reward signal noise when training embodied agents [^1] Code available at https://shorturl.at/VsH70.
-The Dark Side of Rich Rewards:
-Understanding and Mitigating Noise in VLM Rewards
-Sukai Huang1, Shu-Wei Liu2, Nir Lipovetzky1 and Trevor Cohn1thanks: Now at Google DeepMind 1School of Computing and Information Systems, The University of Melbourne, Australia 2 Max Planck Institute for the Physics of Complex Systems, Nöthnitzer Strasse 38, 01187 Dresden, Germany sukaih@student.unimelb.edu.au, sliu@pks.mpg.de, {nir.lipovetzky, trevor.cohn}@unimelb.edu.au
+Our findings offer a nuanced insight of how different types of reward noise impact agent learning and highlight the importance of addressing multimodal reward signal noise when training embodied agents.
 1 Introduction
@@
 In particular, a growing trend in embodied agent learning involves using vision-language models (VLMs) for reward modeling.
-This approach measures the semantic similarity – often quantified by cosine similarity – between the embedding representations of an agent’s behaviors (i.e., past trajectories) and the provided instructions, all within the same embedding space (Kaplan et al., 2017; Goyal et al., 2019, 2020; Du et al., 2023).
-Figure 1: Schematic diagram of false positives in a VLM embedding space.
-The unintended agent’s trajectory (dashed line) may exhibit high cosine similarity to the instruction (solid line) in the embedding space, as indicated by the proximity of their endpoints in the embedding space.
-Despite this apparent similarity, the unintended trajectory fails to fulfill the instruction.
-This illustrates how approximation errors can lead to false positive rewards.
-Refer to Section 5 for more details.
-However, we observed that embodied reinforcement learning (RL) agents trained with VLM rewards, while effective in simplified settings, often struggled with tasks involving complex dynamics and longer action horizons.
-This is evident in several recent works – for instance, Goyal et al.
-(2019) reported the effective use of VLM rewards in Montezuma’s Revenge, a notoriously challenging Atari game.
+This approach measures the semantic similarity – often quantified by cosine similarity – between the embedding representations of an agent’s behaviors (i.e., past trajectories) and the provided instructions, all within the same embedding space (Kaplan et al.
+2017; Goyal, Niekum, and Mooney 2019, 2020; Du et al.
+2023; Rocamonde et al.
+2024; Wang et al.
+2024).
+Figure 1: Illustration of embodied RL agents using VLM reward model
+However, we observed that embodied RL agents trained with VLM rewards, while effective in simplified settings, often struggled with tasks involving complex dynamics and longer action horizons [^1] Code available at https://shorturl.at/3uVGP.
+This is evident in several recent works – for instance, Goyal, Niekum, and Mooney (2019) reported the effective use of VLM rewards in Montezuma’s Revenge, a notoriously challenging Atari game.
 However, we observed that this success was confined to individual sub-tasks and the agent struggled when attempting to scale up to the full game.
@@
 (2023) demonstrated impressive performance of VLM rewards in guiding agents within a 2D survival game.
-However, their study was conducted in a modified environment with a reduced observation and action space using internal game state information and manually defined macro actions.
-Consequently, when tested in the original, unmodified environment, their agent’s performance did not exceed that of agents using only intrinsic (exploration-driven) rewards.
+However, their study was conducted in a modified environment with a reduced observation and action space using internal game state information and manually defined macro actions (see Appendix B).
+Consequently, when we tested their methods in the original, unmodified environment, we found that their agent’s performance did not exceed that of agents using only intrinsic (exploration-driven) rewards.
 The consistent underperformance of VLM rewards, particularly their unexpected failure to outperform intrinsic rewards, raised concerns about their reliability.
@@
 Our findings indicate that noisy reward estimates in VLMs are a key factor contributing to poor learning efficacy.
-Specifically, our analysis centered around two classes of noise: false positives, which involve rewarding unintended trajectories, and false negatives, which occur when correct trajectories are not rewarded.
-We posit that false positive rewards (see Figure 1) are not only more prevalent but potentially more detrimental to the learning process than false negatives, a hypothesis supported by our empirical and theoretical findings.
+Specifically, our analysis centered around two classes of noise: false positives, which involve rewarding unintended trajectories and can mislead the agent into reinforcing suboptimal behaviors; and false negatives, which occur when correct trajectories are not rewarded, thus providing little feedback and reintroduce the reward sparsity issue.
+We posit that false positive rewards are not only more prevalent but potentially more detrimental to the learning process than false negatives, a hypothesis supported by our empirical and theoretical findings.
 Among various sources of reward noise, our study particularly investigates the approximation errors within the commonly used cosine similarity metric.
@@
 To this end, we propose a novel reward function, BiMI (Binary Mutual Information Reward).
-It uses binary reward signals to directly reduce the occurrence of false positives and incorporates a mutual information term to prevent overfitting to noisy signal sources.
-Our experiments demonstrate that BiMI significantly improves the learning efficacy of instruction following agents trained by VLM rewards across various challenging environments.
+It uses binary signals to directly reduce the occurrence of false positives and incorporates a mutual information term to prevent overfitting to noisy signal sources.
+Our experiments demonstrate that BiMI significantly improves the learning efficacy of instruction-following agents trained with VLM rewards across various complex, long-horizon tasks in embodied environments.
+We validate the effectiveness of BiMI in both Markovian and non-Markovian reward settings, showcasing its robustness and generalizability across different types of reward models.
 2 Related Work
-Using VLMs as Reward Models.
-VLMs have been pivotal in robotics studies, serving as reward models that guide agents to follow instructions (Wang et al., 2018; Shridhar et al., 2022; Mahmoudieh et al., 2022).
-While most research primarily focuses on leveraging VLMs to overcome the challenge of manual reward design for complex tasks (Clark, 2016), the impact of reward noise and its implications for policy convergence rates are often overlooked.
-As mentioned in Section 1, some work sidesteps the noisy reward problem by accessing internal state information from the game engine as well as providing predefined action macros (Du et al., 2023; Wang et al., 2023), thereby preventing the accumulation of noise over longer horizon.
-However, understanding the impact of reward noise from VLMs is crucial for developing a reliable language interface for embodied learning agents in real-world applications.
-Mitigating Reward Model Failures.
-Research by Ghosal et al.
-(2022) and Fu et al.
-(2024) has introduced methods to counteract unreliable rewards from learned VLMs.
-These strategies involve employing a parallel exploration-based policy alongside the reward-maximizing policy, thereby reducing reliance on potentially misspecified VLM rewards.
-Our work contributes to this growing body of research by proposing a novel reward function that directly mitigates the impact of false positive rewards from VLM-based models, complementing approaches that use exploration policies to escape local optima.
-Furthermore, we tested the synergy of combining our reward function with exploration strategies, demonstrating how these approaches can be integrated for further advancements.
+We provide a brief overview of the related work for this study. For detailed ones, please refer to Appendix B.
+VLMs as Reward Models.
+Despite the rising use of VLM-based reward models in embodied RL (Kaplan et al.
+2017; Goyal, Niekum, and Mooney 2019, 2020; Du et al.
+2023; Rocamonde et al.
+2024; Wang et al.
+2024, 2019; Shridhar et al.
+2022; Mahmoudieh, Pathak, and Darrell 2022), many works focus on proof-of-concept results in simplified environments, sidestepping the challenges posed by noisy reward signals.
+Critical issues such as reward noise and its effect on policy convergence remain largely understudied.
+In particular, some approaches sidestep noise by leveraging internal state information or predefined action macros (Du et al.
+2023; Wang et al.
+2023).
+However, these methods struggle to scale to real-world applications due to their reliance on environment-specific priors, which lack generalizability in dynamic, unstructured settings.
+Pessimistic Value Estimation.
+The RL community has demonstrated that pessimism under uncertainty improves robustness in settings like offline RL optimization (Uehara and Sun 2021; Bai et al.
+2024).
+These methods conservatively estimate state values to avoid overoptimism in poorly explored or uncertain regions.
+However, such techniques are rarely applied to settings with learned VLM-based rewards, where noise arises not from environmental uncertainty but from misaligned or overconfident reward signals.
+Moreover, these methods remain largely alien to recent studies of VLM reward + RL embodied agent research.
+Our work bridges this gap by applying pessimism to mitigate learned reward noise, addressing a critical yet overlooked challenge in this emerging field.
+Optimistic Exploration in Online RL.
+Pessimistic estimates in RL are mostly an offline phenomenon; online RL instead typically embraces optimism under uncertainty to drive efficient exploration.
+Classic model-based schemes such as RMax initialize all unknown state-action pairs with the maximum possible reward, thereby provably encouraging the agent to visit under-explored regions of the MDP (Brafman and Tennenholtz 2002).
+In the model-free regime, intrinsic reward methods measure the surprise of an agent’s actions, such as the prediction error of a forward model (Burda et al.
+2018; Wan et al.
+2023), or the novelty of state visits (Zhang et al.
+2021).
+Our work investigates the integration of pessimistic, VLM-based reward signals with optimistic intrinsic rewards in an online embodied environment.
+Reward Signal from Human Preference.
+Recent work on RL from Human Feedback (RLHF) (Ouyang et al.
+2022) also leverage expert preference as a reward signal.
+However, our work differs in key aspects.
+Unlike RLHF’s focus on textual outputs, our approach involves evaluating cross-modal similarities between visual and textual data in environments requiring long-horizon decision-making and frequent embodied interactions, a domain not typically covered by RLHF.
+Mitigating Misspecified Rewards.
+Prior works proposed mitigating false positive rewards by training a parallel exploration policy to escape local optima caused by misspecified rewards (Ghosal et al.
+2022; Fu et al.
+2024).
+In contrast, we propose a novel reward function that directly penalizes likely false positive reward signals during training.
+We further show that our method complements exploration-based methods and achieves superior performance when combined.
 3 Formal Problem Statement
@@
 An agent’s trajectory is a sequence of states and actions $\tau_{t}=\langle s_{0},a_{0},\ldots,s_{t}\rangle$.
-In this work, we focus on a sparse reward setting, where the agent receives a $+1$ reward only when reaching goal states $S_{G}\subset\mathcal{S}$, and 0 otherwise, with $|S_{G}|\ll|\mathcal{S}|$.
+In this work, we focus on a sparse reward setting, where the agent receives a +1+1 reward only when reaching goal states $S_{G}\subset\mathcal{S}$, and 0 otherwise, with $|S_{G}|\ll|\mathcal{S}|$.
 This sparse reward setting motivates the use of expert instructions and VLMs to provide auxiliary reward signals for more effective RL.
+Here, it’s crucial to note that by VLM, we specifically mean encoder-based VLMs, as opposed to generative VLMs (see Appendix M).
+For conciseness, we will simply use ‘VLM’ to refer to encoder-based VLMs.
 Specifically, we have a walkthrough $L$ that breaks down a complex task into $n$ expert-defined sub-tasks, each represented by a natural language instruction that is not necessarily atomic and can encompass multiple finer sub-goals ($L=\{l_{1},l_{2},\ldots,l_{n}\}$).
 By following these sequential instructions, the agent can navigate from the initial state towards the goal states.
-A dedicated non-Markovian VLM-based reward model $r^{v}(\tau_{t},l_{m(t)})$ is used to assess how well the agent’s trajectory at current time $t$ fulfills an instruction sentence $l_{m(t)}$.
-Here, $m(t)$ is a pointer at time step $t$ that indicates the current instruction the agent is trying to complete [^2] The use of non-Markovian reward functions in MDP has been well-established, particularly through the work on reward machines (Icarte et al., 2018; Corazza et al., 2022).
-For a complete evaluation, we also tested Markovian version of VLM reward function (Pixl2R) in our experiments..
-The VLM provides auxiliary rewards by evaluating the semantic similarity between $\tau_{t}$ and $l_{m(t)}$, as illustrated in Figure 2 (see Appendix A.1 for details).
-Figure 2: Illustration of reward from a VLM
+A dedicated non-Markovian VLM-based reward model [^2] The use of non-Markovian reward functions in MDP has been well-established, particularly through the work on reward machines (Icarte et al.
+2018; Corazza, Gavran, and Neider 2022).
+For a complete evaluation, we also tested Markovian version of VLM reward function (Pixl2R) in our experiments.
+$r^{v}(\tau_{t},l_{m(t)})$ is used to assess how well the agent’s trajectory at current time $t$ fulfills an instruction sentence $l_{m(t)}$.
+For the detailed mechanism and the pseudo-code of the existing VLM reward + RL algorithm, please refer to Appendix C.
+Note that our proposed BiMI reward function uses a different mechanism to determine if instruction $l_{m(t)}$ is completed at time $t$ (explained shortly in Section 7).
+The VLM provides auxiliary rewards by evaluating the semantic similarity between $\tau_{t}$ and $l_{m(t)}$, as illustrated in Figure 1.
+The use of VLM-based reward model transform the original MDP into a shaped MDP $\widetilde{\mathcal{M}}=\langle\mathcal{S},\mathcal{A},\mathcal{P},s_{0},\tilde{r},\gamma\rangle$, where the reward function becomes $\tilde{r}=r^{e}(s,a)+r^{v}(\tau_{t},l_{m(t)})$.
 4 Theoretical Analysis
-In this section, we theoretically show how auxiliary rewards influence policy convergence.
-Initially, we establish that providing auxiliary rewards that reflect progress towards the goal state typically accelerates the convergence process compared to scenarios where only sparse environmental reward $r^{e}$ is used.
-Then, we show that the presence of false positive rewards has a more detrimental effect on convergence than false negatives.
+In this work, we define “reward noise” as errors made by the VLM-based reward model when evaluating an agent’s trajectory: either spuriously rewarding a trajectory that doesn’t satisfy the instruction (false positive) or failing to reward one that does (false negative).
+In this section, we first show that, in the absence of noise, auxiliary rewards derived from language instructions, reflecting progress toward the goal, accelerate convergence compared to relying solely on sparse environmental rewards, $r^{e}$.
+We then prove that false positive rewards impede convergence more severely than false negatives under the Heuristic guided RL (HuRL) framework (Cheng et al.
+2021).
 4.1 Auxiliary Rewards and Convergence
-With sparse rewards, the gradient landscape is nearly flat, making gradient-based updates indistinguishable from a random walk in parameter space (Antognini and Sohl-Dickstein, 2018).
+With sparse rewards, the gradient landscape is nearly flat, making gradient-based updates indistinguishable from a random walk in parameter space (see Proposition N.1 in Appendix N).
 In this problem, we are interested in $D:=\|\theta_{goal}-\theta_{0}\|$ which is the distance in parameter space from initial parameters $\theta_{0}$ to goal parameters $\theta_{goal}$.
-We make the following assumption
+We make the following assumption:
 Assumption 4.1.
@@
 As a result, the overall distance $D$ can be decomposed into segments: $D\approx\sum_{k=1}^{n-1}d_{i}$, where $d_{i}=\|\theta_{i+1}-\theta_{i}\|$.
-We also prove the following proposition
+We therefore prove the following proposition:
 Proposition 4.2.
-The sum of expected time for a series of random walks, each covering the shorter distance of an individual sub-task, is less than the expected time to travel the entire distance $D$ in one long random walk:
-$\frac{1}{n-1}\mathbb{E}[T_{D}]\leq\mathbb{E}\left[\sum_{i=1}^{n-1}T_{d_{i}}\right]<\mathbb{E}[T_{D}].$
-and thus show that subgoal-based auxiliary rewards improve the convergence of random walk optimization in a sparse reward landscape up to a factor of $(n-1)$ with details in Appendix A.2.1.
+The sum of expected time for a series of random walks, each covering the shorter distance of an individual sub-task, is less than the expected time to travel the entire distance $D$ in one long random walk: $\frac{1}{n-1}\mathbb{E}[T_{D}]\leq\mathbb{E}\left[\sum_{i=1}^{n-1}T_{d_{i}}\right]<\mathbb{E}[T_{D}]$.
+and thus show that subgoal-based auxiliary rewards improve the convergence of random walk optimization in a sparse reward landscape up to a factor of (n1)(n-1).
+See Appendix D for the proof.
+The theoretical takeaway is that auxiliary rewards aligned with intermediate subgoals can accelerate convergence of RL algorithms compared to learning from a single delayed reward at the goal state.
+In practice, however, this improvement depends critically on the quality of the auxiliary reward signal — which we examine in the next section.
 4.2 Connection to Heuristic-Guided RL
-The problem and the possible solutions can be framed within the context of Heuristic-Guided Reinforcement Learning (HuRL) by Cheng et al.
+The reward noise issue can be framed within the context of HuRL by Cheng et al.
 (2021).
 HuRL mandates that auxiliary reward signals serve as heuristics, where $h:\mathcal{S}\rightarrow\mathbb{R}$ approximates the future total rewards an agent expects to get starting from state $s$ under the optimal policy $\pi^{*}$ (i.e., $h(s)\approx V^{*}(s)$).
-These heuristics typically come from domain knowledge, aligning with expert instructions.
-We make the following assumption
+First of all, we establish the following assumption:
 Assumption 4.3.
-The fulfillment of $l_{m(t)}$ can be captured by intermediate goal states $\mathcal{S}^{m(t)}$.
-This assumption is justified because often instructions in VLM come with end goals that can be indicated by definite states.
-For further discussions, please refer to Section 9.
-This assumption is required so that we can construct a domain-knowledge Markovian heuristic reward based on non-Markovian VLM rewards (see Appendix A.1.3 for the exact formula).
-With this construction, we are estimating the possible reward in the future by evaluating how well the agent has followed expert instruction in the past.
-It is consistent with the intuition of a heuristic reward based on semantic similarity because the better the agent has been following expert instructions, the more likely it is on the right track to collect more future rewards.
-Using HuRL framework allows us to analyze how false positive rewards influence on policy performance gap, defined as $V^{*}(s_{0})-V^{\pi}(s_{0})$.
-We will demonstrate that false positive rewards increase the upper bound of this gap, whereas false negative rewards maintain this upper bound.
-4.3 False Positives vs. False Negatives
-To begin, we provide the formal definition of false positive auxiliary rewards from both instruction-following (IF) and heuristic perspectives:
+The expert instruction sequence $L={l_{1},l_{2},\dots,l_{n}}$ specifies a sequence of state and action landmarks guiding the agent toward goal states $\mathcal{S}_{G}$.
+Landmarks are key states or key actions that must hold or occur in any trajectories that reach $\mathcal{S}_{G}$.
+Assumption 4.3 is reasonable and reflects the quality of expert instructions, which are crucial for the agent to reach the goal efficiently.
+We now present the important proposition connecting VLM-based reward models to HuRL:
+Proposition 4.4.
+In sparse reward settings, VLM-based reward models, as implemented in Algorithm 1, can be viewed as a heuristic function $h(s_{t})$ in HuRL that estimates $V^{*}(s_{t})$.
+The proof is in Appendix E.
+HuRL framework allows us to analyze how false positive rewards influence on optimality gap, defined as $V^{*}(s_{0})-V^{\pi}(s_{0})$, a metric that is used for theoretical analysis of the convergence speed of RL algorithms — the smaller the gap, the faster the convergence.
+We demonstrate that false positive rewards increase the upper bound of this gap, whereas false negative rewards maintain this upper bound.
+4.3 False Positives and Heuristic Overestimation
+We provide the formal definition of false positive and false negative rewards from both instruction-following (IF) and heuristic perspectives:
 Definition 4.5 (False Positive Rewards).
 A false positive reward occurs when:
-IF Perspective: For a trajectory $\tau_{t}$ that does not satisfy instruction $l_{m(t)}$, the VLM-based reward $r^{v}(\tau_{t},l_{m(t)})$ ranges between 00 and $1$.
+IF Perspective: For a trajectory $\tau_{t}$ that does not satisfy instruction $l_{m(t)}$, the VLM-based reward $r^{v}(\tau_{t},l_{m(t)})>0$.
 Heuristic Perspective: The heuristic $h(s_{t})>V^{*}(s_{t})$, overestimating the optimal value of $s_{t}$.
+Definition 4.6 (False Negative Rewards).
+A false negative reward occurs when:
+IF Perspective: The VLM-based reward $r^{v}(\tau_{t},l_{m(t)})\approx 0$ for a trajectory $\tau_{t}$ that does satisfy instruction $l_{m(t)}$.
+Heuristic Perspective: The heuristic $h(s_{t})\approx 0<V^{*}(s_{t})$, underestimating the optimal value of $s_{t}$.
 Proposition 4.7.
 False positive rewards in the IF perspective imply false positive rewards in the heuristic perspective.
-The heuristic is approximated as $h(s_{t})\approx V^{*}(s_{t})=1\cdot\gamma^{\widetilde{T}-t}$, calculated based on an assumed optimal path length $\widetilde{T}$.
+Proof.
+The heuristic is approximated as $h(s_{t})\approx V^{*}(s_{t})=1\cdot\gamma^{\widetilde{T}-t}$ in sparse reward setting, calculated based on an assumed optimal path length $\widetilde{T}$.
 Here, $\widetilde{T}-t$ measures the remaining steps towards the goal.
 When a trajectory $\tau_{t}$ receives a high reward but fails to fulfill instruction $l_{m(t)}$, it corresponds to a high $h(s_{t})$, thus a low $\widetilde{T}-t$.
-However, since the agent must eventually redo $l_{m(t)}$, $s_{t}$ is actually further from the goal than estimated.
-Therefore, the actual distance $T-t$ to reach the goal will exceed $\widetilde{T}-t$, and $V^{*}(s_{t})$, calculated with the actual $T$, will be smaller than $h(s_{t})$, derived from $\widetilde{T}$.
+Yet the agent must eventually reattempt $l_{m(t)}$ due to Assumption 4.3, $s_{t}$ is actually further from the goal than estimated.
+Therefore, the actual distance TtT-t to reach the goal will exceed $\widetilde{T}-t$, and $V^{*}(s_{t})$, calculated with the actual $T$, will be smaller than $h(s_{t})$, calculated with $\widetilde{T}$.
 This thus explains how a false positive reward from an instruction-following (or VLM) perspective leads to an overestimation of the heuristic.
-Researchers have advocated the benefits of pessimistic value estimation to enhance the stability of RL algorithms (Kumar et al., 2020; Jin et al., 2021).
+∎
+Researchers have advocated the benefits of pessimistic value estimation to enhance the stability of RL algorithms (Kumar et al.
+2020; Jin, Yang, and Wang 2021).
 In HuRL, Cheng et al.
-(2021) further identify a beneficial property: when a heuristic is pessimistic with respect to $\mathcal{M}$, it results in a smaller upper bound on the performance gap.
-Definition 4.6 (Pessimistic $h$).
-Let Bellman operator $(\mathcal{B}h)(s,a)=r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[h(s^{\prime})]$.
-A heuristic function $h$ is said to be pessimistic with respect to an MDP $\mathcal{M}$ if $\max_{a}(\mathcal{B}h)(s,a)\geq h(s)$.
+(2021) further identify a beneficial property: when a heuristic is Bellman-consistent pessimism (i.e., $\max_{a}(\mathcal{B}h)(s,a)\geq h(s)$), it results in a capped upper bound on the optimality gap.
+Definition 4.8 (Bellman-consistent Pessimistic $h$).
+Recall Bellman equation of $h$: $(\mathcal{B}h)(s,a)=r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[h(s^{\prime})]$.
+A heuristic function $h$ is said to be Bellman-consistent pessimism with respect to an MDP $\mathcal{M}$ if $\max_{a}(\mathcal{B}h)(s,a)\geq h(s)$.
 This condition essentially means that the heuristic $h$ never overestimates the true value of a state.
 Proposition 4.9.
-Even if the heuristic remains conservative for all successor states, a single overestimation ($h(s)>V^{*}(s)$) can violate the pessimistic condition by causing $max_{a}(\mathcal{B}h)(s,a)<h(s)$.
-Informally, our result on the impact of false positive/negative rewards on convergence is as follows:
+Even if the heuristic remains conservative for all successor states, a single false positive overestimation, i.e., $h(s)>V^{*}(s)$, can violate the pessimistic condition by causing $max_{a}(\mathcal{B}h)(s,a)<h(s)$.
+The proof is in Appendix F. This implies that maintaining a pessimistic heuristic is inherently fragile, as the presence of a false positive in any state disrupts the pessimistic condition.
+Our result on the impact of false positive and false negative rewards on convergence is captured as follows:
 Theorem 4.10.
-Borrowing from Cheng et al.
-(2021), the performance gap in RL when using heuristics can be broken down into a regret term and a bias term, where false negatives maintain the upper bound of bias while false positives increase the bias by breaking the heuristic’s pessimism, thereby potentially leading to slower convergence.
-See Appendix A.2.2 for formal version and detailed proofs.
+Adapting the optimality gap analysis from Cheng et al.
+(2021), the gap $V^{*}(s_{0})-V^{\pi}(s_{0})$ in heuristic-guided RL decomposes into regret and bias terms.
+False negatives preserve the upper bound on bias, as they maintain the heuristic’s pessimism, i.e., $h(s)\leq V^{*}(s)$.
+In contrast, false positives breaks the pessimistic property, strictly increase the bias without an upper bound, thereby leading to a larger optimality gap and slower convergence.
+See Appendix G for the formal version and detailed proof.
 In the next section, we present a case study on cosine similarity metrics, demonstrating how they contribute to false positive rewards in learned reward models.
 5 False Positives From Cosine Similarity
+Figure 2: Schematic diagram of false positives in embedding space.
 This section identifies and discusses two fundamental issues with cosine similarity scores in sequential decision-making contexts: state entanglement and composition insensitivity.
@@
 For instance, consider the significant contrast between “take the red pill” and “take the blue pill”.
-Despite their lexical similarity, they lead to vastly different states.
-However, the cosine similarity metric may represent them as similar due to the shared words, disregarding the critical difference in state outcomes.
+Despite their lexical similarity, they lead to different states.
+However, the cosine similarity metric may represent them as similar due to the shared words, disregarding the critical difference in state outcomes (see Figure 2 top left).
 Understanding state transitions is crucial in sequential decision-making scenarios.
 Otherwise, rewards may be given to trajectories that lead to unintended states, potentially prolonging the path to the goal state by necessitating corrective actions or re-attempts.
-The Issue of Composition Insensitivity Composition insensitivity in cosine similarity metrics gives rise to two issues: (1) rewarding incomplete task execution – cosine similarity may incorrectly reward partial task completion, as even incomplete trajectories can receive high similarity score in the embedding space.
-For instance, in a task to “pick up a tool, then fix a machine,” the model might prematurely reward the agent for merely picking up the tool, neglecting the crucial repair action.
-We also observed this phenomenon particularly in the Montezuma environment, where RL agents tend to hack the reward system by focusing on the easiest actions that yield rewards (e.g., moving towards a direction) rather than executing more complex, timely actions.
-Eventually, this leads to an overestimation of the agent’s progress towards the ultimate goal.
-(2) insensitivity to the ordering of execution – cosine similarity often fails to adequately penalize incorrect execution sequences.
-In a safety protocol requiring an agent to “turn off the machinery, then perform maintenance,” the metric might assign high rewards based merely on the presence of relevant actions, disregarding their order.
-In contrast to some advancements in language models, compact visual and sentence embeddings from multimodal VLMs remain largely insensitive to sequential information (Pham et al., 2021).
-When the task is order-sensitive, executing actions in the wrong sequence prolongs the path towards the goal state, as agents need to re-attempt the correct order.
-Figure 1 illustrates various scenarios where false positive rewards are erroneously assigned.
-To empirically demonstrate the issue, Section 5.1 presents experiments on these issues and their impact on agent learning in sparse reward environments.
+The Issue of Composition Insensitivity One might wonder why composition insensitivity persists despite using a pointer mechanism (Appendix C) that enforces sequential instruction order.
+The issue arises because sentences $l$ in the instruction sequence $L$ are not always atomic: a single instruction like “tidy the room before leaving” may implicitly encode multiple sub-tasks.
+While the pointer ensures progress through high-level steps, it does not help the VLM parse the internal structure of complex, non-atomic instructions.
+As a result, “composition insensitivity” emerges, misaligning the agent’s trajectory and undermining reward accuracy.
+Unfortunately, even state-of-the-art natural language understanding models struggle with decomposing atomic sentences without losing critical sequential or contextual information (Dziri et al.
+2024).
+Under this circumstance, composition insensitivity in cosine similarity metrics gives rise to two issues: (1) rewarding incomplete task execution – cosine similarity may incorrectly reward incomplete task execution by giving high scores even when critical steps are missing.
+We observed this phenomenon particularly in the Montezuma environment, where RL agents tend to hack the reward system by focusing on the easiest actions that yield rewards (e.g., moving towards a direction) rather than executing more complex, timely actions.
+Eventually, this leads to an overestimation of the agent’s progress towards the ultimate goal (see Figure 2 bottom left).
+(2) insensitivity to the ordering of execution – VLM models often fails to adequately penalize incorrect execution sequences, rather, it assigns high rewards based merely on the presence of relevant actions, disregarding their order (see Figure 2 top right).
+In contrast to some advancements in language models, compact visual and sentence embeddings from multimodal VLMs remain largely insensitive to sequential information (Pham et al.
+2020).
+When the task is order-sensitive, executing actions in the wrong sequence prolongs the path towards the goal state, as agents need to re-attempt the correct order — yet the VLM’s reward may still incentivize the agent to continue with the incorrect sequence.
+We empirically demonstrate these issues in Section 6, showing their substantial impact on agent learning in sparse reward environments.
+6 Experiments on Reward Noise Impact
+Our experiments test the following hypothesis: (H1) The two issues of state entanglement and composition insensitivity exist, (H2) false positive rewards are prevalent during training, (H3) VLM reward models lacking noise handling mechanisms underperform against intrinsic reward models in sparse reward environments, (H4) false negatives may not be as harmful as false positives.
 Figure 3: Learned VLM models performed badly with O.O.D.
@@
 Figure 4: The false positive vs. false negative oracle model. The false positive model get a more severe drop in the final training score.
-5.1 Experiments on Reward Noise Impact
-Our experiments test the following hypothesis: (H1) The two issues of state entanglement and composition insensitivity exist; (H2) false positive rewards are prevalent during training; (H3) VLM reward models lacking noise handling mechanisms underperform against intrinsic reward models in sparse reward environments; (H4) false negatives may not be as harmful as false positives.
-Setup.
-We evaluate these hypotheses through various challenging sparse-reward environments: (1) Crafter, an open-ended 2D Minecraft (Hafner, 2021); (2) Montezuma, a classic hard adventure game in Atari (Bellemare et al., 2013); and (3) Minigrid ‘Go To Seq’, a hard task involving long-horizon navigation and object interactions (Chevalier-Boisvert et al., 2018).
-A Markovian and a Non-Markovian reward model were tested: (1) Pixl2R by Goyal et al.
-(2020), which uses only the current video frame to determine if the goal state specified in the instruction has been reached; and (2) ELLM-, a variant of ELLM by Du et al.
-(2023).
-Unlike ELLM, which queries instructions from LLMs in real-time, ELLM- directly uses preset expert instructions and compares them with the transition differences of the agent’s trajectory.
-The VLM backbones used are: (1) CLIP (Radford et al., 2021), pretrained by image-text pairs; and (2) X-CLIP (Ma et al., 2022), pretrained by video-text pairs.
-To ensure high-quality finetuning data, we used internal information from the game engine to annotate expert trajectories from expert agents.
-To demonstrate how noisy reward signals hinder learning, we selected a strong intrinsic reward model DEIR (Zhang et al., 2021) for comparison.
-It provides auxiliary rewards based on observation novelty to encourage exploration.
-See Appendix A.4 for detailed implementation of the experiments.
-Evaluation Metric.
-We adopted the score metric from the Crafter benchmark (Hafner, 2021) for performance evaluation, as it effectively measures consistent performance across multiple subtasks in sparse reward environments.
-Unlike the maximum total rewards metric, which does not adequately reflect consistent performance, the score metric offers a more reliable indicator of learning progress.
-See Appendix A.5 for its formal definition.
-Reward Noise Issue.
-To investigate H1, we evaluated the models’ sensitivity by examining how cosine similarity scores change for manipulated trajectory-instruction pairs.
-The state entanglement test involved reversing trajectories and negating instructions (i.e., “do not do $l_{k}$”).
-The composition insensitivity test examined concatenated pairs of trajectory-instruction data.
-Given $(\tau_{1},l_{1})$ and $(\tau_{2},l_{2})$, we create a concatenated pair $(\tau_{1}+\tau_{2},l_{1}+l_{2})$.
-We then test two types of manipulations – (1) swapping the order within one modality: e.g., $(\tau_{2}+\tau_{1},l_{1}+l_{2})$; and (2) truncating one modality: e.g., $(\tau_{1},l_{1}+l_{2})$.
-Overall, three types are evaluated: (1) matched pairs; (2) not-matched pairs and (3) manipulated pairs which are derived from matched pairs by polluting either the trajectory or the instruction.
-Our results reveal a critical flaw in the reward model: despite manipulated pairs that fundamentally fail to fulfill the instruction, the model paradoxically assigns high similarity scores (see Figure 4 for overall results and also Appendix A.6 for individual environments).
-It’s worth noting that the poor performance in the negation case aligns with broader challenges in natural language processing.
-Recent studies (Hossain et al., 2022; Truong et al., 2023) have highlighted that negation is central to language understanding but is not properly captured by modern language models.
-This limitation extends to VLMs and directly leads to false positive rewards.
 Figure 5: The heatmap shows rewards received at various locations, with larger circle sizes indicating higher rewards.
@@
 Figure 6: The ratio of false positive rewards is significantly reduced after applying BiMI
+Experimental Setup We evaluate these hypotheses through various challenging sparse-reward environments: (1) Crafter, an open-ended 2D Minecraft (Hafner 2022); (2) Montezuma, a classic hard adventure game in Atari (Bellemare et al.
+2013); and (3) Minigrid ‘Go To Seq’, a hard task involving long-horizon navigation and object interactions (Chevalier-Boisvert et al.
+2018).
+Details on why these environments are preferred over other test benchmarks, along with the evaluation metrics, are provided in Appendix A.
+A Markovian and a Non-Markovian reward model were tested: (1) Pixl2R by Goyal, Niekum, and Mooney (2020), which uses only the current video frame to determine if the goal state specified in the instruction has been reached; and (2) ELLM-, a variant of ELLM by Du et al.
+(2023) that directly uses preset expert instructions and compares them with the transition differences of the agent’s trajectory.
+The VLM backbones used are: (1) CLIP (Radford et al.
+2021), pretrained by image-text pairs; and (2) X-CLIP (Ma et al.
+2022), pretrained by video-text pairs.
+To ensure high-quality finetuning data, we used internal information from the game engine to annotate expert trajectories from expert agents.
+To demonstrate how noisy reward signals hinder learning, we selected a strong intrinsic reward model DEIR (Wan et al.
+2023) for comparison.
+It provides auxiliary rewards based on observation novelty to encourage exploration.
+The detailed VLM-based reward model procedure, finetuning process and training hyperparameters are in Appendix H.
+Reward Noise Issue To investigate H1 — the existence of two issues, state entanglement and composition insensitivity, in VLM-based reward models — we evaluated the models’ sensitivity by examining how cosine similarity scores change for manipulated trajectory-instruction pairs.
+These manipulations, designed to probe to robustness against noise, included the following:
+  1. 1.
+    Trajectory Reversal: We inverted the sequence of frames within each trajectory (i.e., frames = frame[::-1]) to assess the model’s ability to detect reversed state transitions.
+This manipulation tests whether the model can distinguish between forward and backward progression in the state transition.
+  2. 2.
+    Instruction Negation: We modified the original instructions by adding negation (e.g., changing “do $l_{k}$” to “do not do $l_{k}$” or “avoid $l_{k}$”).
+This tests the model’s sensitivity to semantic changes in the instruction that fundamentally alter the goal.
+  3. 3.
+    Concatenation and Order Swapping: Given two trajectory-instruction pairs $(\tau_{1},l_{1})$ and $(\tau_{2},l_{2})$, we created concatenated pairs and then swapped the order in one modality.
+For example: a) Original concatenation: $(\tau_{1}+\tau_{2},l_{1}+l_{2})$; b) Swapped trajectory: $(\tau_{2}+\tau_{1},l_{1}+l_{2})$; c) Swapped instruction: $(\tau_{1}+\tau_{2},l_{2}+l_{1})$.
+This tests the model’s sensitivity to the order of components in multi-step tasks.
+  4. 4.
+    Concatenation with Partial Content: We concatenated pairs but truncated one modality.
+For instance: a) Truncated trajectory: $(\tau_{1},l_{1}+l_{2})$; b) Truncated instruction: $(\tau_{1}+\tau_{2},l_{1})$.
+This assesses the model’s ability to detect partial mismatches in longer sequences.
+Our results reveal a critical flaw in the reward model: for manipulated pairs that fundamentally fail to fulfill the instruction, the model paradoxically assigns high similarity scores (see Figure 4 for overall results and also Appendix I for individual environments).
+Note that the poor performance in the negation case aligns with broader challenges in natural language processing.
+Recent studies (Hossain et al.
+2022; Truong et al.
+2023) have highlighted that negation is central to language understanding but is not properly captured by modern language models.
+This limitation extends to VLMs and directly leads to false positive rewards.
+Prevalence of False Positives To address H2, we analyzed reward distribution heatmap from VLM-based reward models during training.
+The heatmap (Figure 6 left) revealed a concerning trend: RL agents engage in reward hacking, receiving rewards across vast areas of the environment rather than just at goal states.
+For instance, in Montezuma where the goal is to grab the key and escape the room, we observed that agents received rewards even for falling off cliffs, which undoubtedly contribute to false positive rewards.
+For environments without fixed camera views, we calculated the step offset between the current rewarded state and the actual goal state.
+A positive offset indicates a false positive reward, as the reward was given before reaching the goal.
+Conversely, a negative offset indicates a false negative, where the agent reached the goal but the reward model failed to acknowledge it (see Figure 6).
+Notably, besides positive offsets, we observed a large amount of negative offsets in Minigrid environments.
+We attribute this to Minigrid’s abstract shape-based visual representations, which fall outside the VLM’s training distribution.
+Impact on Learning We trained agents using learned VLM reward models and compared
 Table 1: Score metric across environments (equivalent to total rewards, higher is better).
@@
 Models Type Monte. Minigrid Crafter % vs. DEIR
-PPO Pure $0.151$ $24.9$ $16.8$ $-28\%$
+[…]
 DEIR $\star$ Intrinsic $0.174$ $55.5$ $19.7$
-Pixl2R VLM $0.142$ $12.4$ $9.40$ $\mathbf{-49\%}$
-ELLM- VLM $0.150$ $19.4$ $10.8$ $-41\%$
-Pixl2R + DEIR VLM + intr. $0.176$ $17.3$ $10.4$ $-38\%$
-ELLM- + DEIR VLM + intr. $0.178$ $30.9$ $11.8$ $-27\%$
-Prevalence of False Positives.
-To address H2, we analyzed reward distribution heatmap from VLM-based reward models during training.
-The heatmap revealed a concerning trend: RL agents engage in reward hacking, receiving rewards across vast areas of the environment rather than just at goal states.
-For instance, in Montezuma where the goal is to grab the key and escape the room, we observed that agents received rewards even for falling off cliffs, which undoubtedly contribute to false positive rewards.
-For environments without fixed camera views, we calculated the step offset between the current rewarded state and the actual goal state.
-A positive offset indicates a false positive reward, as the reward was given before reaching the goal.
-Conversely, a negative offset indicates a false negative, where the agent reached the goal but the reward model failed to acknowledge it (see Figure 6).
-Interestingly, besides positive offsets, we observed a large amount of negative offsets in Minigrid environments.
-We attribute this to Minigrid’s abstract shape-based visual representations, which fall outside the VLM’s training distribution.
-Impact on Learning.
-We trained agents using learned VLM reward models and compared their learning efficacy against intrinsic reward models.
+[…]
+[…]
+[…]
+[…]
+their learning efficacy against intrinsic reward models.
 As shown in Table 1, our results confirmed H3: instruction-following RL agents using learned VLM reward models without noise handling consistently underperform compared to DEIR, the intrinsic reward-based RL agent.
-To investigate the impact of false negatives versus false positives (H4), we designed an oracle Pixl2R model with two variants: a false negative model and a false positive model.
-The false negative model only rewards the agent for reaching subgoal states described in the instruction, with a probability of xthat some rewarding states in the map are removed.
-In contrast, the false positive model rewards the agent for reaching every subgoal, but also introduces a small (0.1) one-off reward for certain locations, covering xof the map.
+In our efforts to assess the impact of false positive rewards from auxiliary reward model without the interference of other factors such as domain shift, poor data quality, and errors from other issues such as the choice of multimodal architectures, we devised a simulated reward model that access to internal state information from the game engine.
+Experiments with these simulated models provide initial evidence of the differential impact of false negatives versus false positives on training outcomes, as posited by H4, suggesting that false positives — particularly those tied to temporal insensitivity — more severely degrade final scores.
+Due to page limits, we move the details to Appendix J.
+To compare the effects of reward noise, we designed an oracle Pixl2R model with two variants: one simulating false negatives by randomly removing $x\%$ of true subgoal rewards, and one simulating false positives by adding small one-off rewards (0.1) to $x\%$ of irrelevant states.
+This design reflects a realistic failure mode of VLMs, where false positive rewards — despite being only 1/10 the magnitude of true subgoal rewards and sometimes located far from the optimal trajectory — can still be exploited by the agent through training iterations, a phenomenon known as reward hacking.
+In contrast, false negatives merely reduce the frequency of positive feedback without misleading the agent.
 The results indicate that false negatives were less detrimental to agent performance than false positives (see Figure 4).
-This performance difference can be explained through our theoretical analysis in Appendix A.2.2.
+This aligns with our theoretical analysis in Theorem 4.10.
 7 Addressing the False Positive Issue
-Our proposed solution is not specific to cosine similarity’s approximation error, which serves as a case study.
-Rather, it’s a broad strategy for mitigating false positive rewards across various reward models, including noisy VLM reward sources.
-6.1 Binary Signal and Conformal Prediction
-Our experiments have demonstrated that false positives are more detrimental to learning than false negatives.
-Based on these findings, we propose a reward function that issues a one-time binary reward only when the similarity between the agent’s current trajectory and the instruction exceeds a high confidence threshold.
+Binary Signal and Conformal Prediction  Having established the detrimental effects of false positive rewards, our solution is to strategically manage the trade-off: reducing false positives, even at the cost of a slight increase in false negatives.
+First, we propose a reward function that issues a one-time binary reward only when the similarity between the agent’s current trajectory and the instruction exceeds a high confidence threshold.
 This approach contrasts with previous methods, which provide continuous rewards whenever the reward score exceeds a predefined threshold, and continue to do so until reaching a maximum cap.
 Our method, however, delivers this reward only once.
-This approach minimizes the likelihood of accumulating false positive rewards while maintaining adherence to Assumption 4.3.
+This approach minimizes the likelihood of accumulating false positive rewards while maintaining adherence to Proposition 4.4.
 To achieve this, we introduce a thresholding mechanism using a calibration set of true positive trajectory-instruction pairs.
@@
 Pairs whose similarity scores fall below this threshold $\hat{q}$ receive no reward.
-Conversely, pairs exceeding $\hat{q}$ receive a one-time $+1$ reward, i.e., $r^{v}_{\textsc{Bi}}(\tau,l_{k})=\mathbf{1}_{\{p(l_{k}\mid\tau)\geq\hat{q}\}}$.
-This thresholding approach leverages statistical properties studied by Sadinle et al.
-(2019), which ensures a high probability (at least $1-\alpha$) that true positive pairs are recognized while minimizing frequency of false positives errors.
-See Appendix A.4.2 for the pseudo-code of the instruction-following RL algorithm and Appendix A.8 for detailed threshold calculation.
-6.2 Mutual Information Maximization
-Intuitively, when we observe rewards coming from a particular signal source too frequently, we tend to downplay the significance of that signal to avoid over-reliance.
+Conversely, pairs exceeding $\hat{q}$ receive a one-time +1+1 reward:
+$r^{v}_{\textsc{Bi}}(\tau,l_{k})=\mathbf{1}_{\{p(l_{k}\mid\tau)\geq\hat{q}\}}$ (1)
+It statistically guarantees a high probability (at least $1-\alpha$) that the true positive pairs are recognized (i.e., above the threshold) while minimizing the average number of mistakes predicting false positives (Sadinle et al.
+2019).
+The threshold calculation is detailed in Algorithm 2.
+Table 2: Model score across various environments.
+$\star$ is the baseline agents with a learned VLM-based reward model to compare with.
+BiMI significantly improves performance in Montezuma and Minigrid, while showing mixed results in Crafter
+Methods Montezuma % vs. $\star$ Minigrid % vs. $\star$ Crafter % vs. $\star$
+Pixl2R $\star$ $0.142\pm 0.003$ $12.4\pm 2.43$ $9.40\pm 0.022$
+[…]
+[…]
+Pixl2R + DEIR $0.176\pm 0.009$ $23\%$ $17.3\pm 0.51$ $39\%$ $4\pm 1.015 +10+10\%$
+[…]
+ELLM- $\star$ $0.150\pm 0.004$ $19.4\pm 10.06$ $10.8\pm 1.017$
+[…]
+[…]
+ELLM + DEIR $0.178\pm 0.029$ $20\%$ $30.9\pm 3.50$ $59\%$ $11.8\pm 1.152$ $9.5\%$
+[…]
+Figure 7: BiMI reward showed faster and higher success rates on difficult tasks in Montezuma
+Mutual Information Maximization Intuitively, when we observe rewards coming from a particular signal source too frequently, we tend to downplay the significance of that signal to avoid over-reliance.
 This intuition is effectively captured by incorporating a mutual information maximization term into the reward function.
 Specifically, the updated reward function $r^{v}_{\textsc{MI}}(\tau,l_{k})$ measures the mutual information between the agent’s trajectory and the instruction.
-Mathematically, it can be expressed as:
+Formally, it can be expressed as:
 $r^{v}_{\textsc{MI}}(\tau,l_{k})$ $=I(l_{k};\tau)=D_{KL}(p(l_{k},\tau)\mid\mid p(l_{k})p(\tau))$
-$=\mathbb{E}_{\tau\sim\pi_{\theta},l_{k}\sim L}[\log p(l_{k}\mid\tau)-\log p(l_{k})]$
+$=\mathbb{E}_{\tau\sim\pi_{\theta},l_{k}\sim L}[\log p(l_{k}\mid\tau)-\log p(l_{k})]$ (2)
 where $\tau=\langle s_{t-W},a_{t-W},\ldots,s_{t}\rangle$ is the agent’s trajectory up to current time step $t$, and $W$ is the memory size of the agent for its past trajectory.
-$p(l_{k}\mid\tau)$ comes from the similarity score provided by VLMs, referring to the likelihood of the instruction $l_{k}$ being fulfilled by a trajectory $\tau$.
+$p(l_{k}\mid\tau)$ comes from the similarity score provided by VLMs, referring to the likelihood of the instruction $l_{k}$ being fulfilled by trajectory $\tau$.
 $p(l_{k})$ is overall likelihood of encountering the instruction $l_{k}$ in the learning environment.
@@
 For instance, if a VLM frequently detects that the agent’s actions are fulfilling the “climbing the ladder” instruction, even when the agent is performing unrelated tasks, any reward signal from this instruction will be downplayed.
-$p(l_{k})$ is calculated as follows:
-$p(l_{k})=\mathbb{E}_{\tau\sim\pi_{\theta_{-1}}}\left[\left(\sum\nolimits_{t=1}^{T_{\tau}}\mathbf{1}_{\{p(l_{k}\mid\tau_{t})\geq\hat{q}\}}\right)/T_{\tau}\right]$
-Here, $\tau_{t}$ is the agent’s trajectory up to time $t$, and if the VLM deems the trajectory as fulfilling the instruction (i.e., $p(l_{k}\mid\tau_{t})\geq\hat{q}$), we increment the count.
-Dividing the count by the total trajectory length $T_{\tau}$ gives the empirical frequency of the instruction being fulfilled.
+We treat the set of instructions $\{l_{k}\}$ as a finite discrete support and define an empirical frequency:
+$s(l_{k})\;=\;\mathbb{E}_{\tau\sim\pi_{\theta_{-1}}}\!\Bigl[\tfrac{1}{T_{\tau}}\sum_{t=1}^{T_{\tau}}\mathbf{1}\bigl\{p(l_{k}\mid\tau_{t})\geq\hat{q}\bigr\}\Bigr$], (3)
+where $\tau_{t}$ is the agent’s trajectory up to time $t$, and we count one whenever the VLM scores $\tau_{t}$ as fulfilling $l_{k}$ over the total trajectory length $T_{\tau}$.
+We then normalize to obtain a probability mass function (PMF) over instructions: $p(l_{k})\;=\;\frac{s(l_{k})}{\sum_{j}s(l_{j})}$, and $\sum_{j}s(l_{j})=1$.
+Because $\{l_{k}\}$ is discrete, $p(l_{k})$ is a valid PMF (not a continuous density) and can be plugged directly into the usual discrete mutual information in Section 7.
 The subscript $\theta_{-1}$ in $\pi_{\theta_{-1}}$ indicates that the trajectories are sourced from rollouts in the previous policy iteration, acknowledging the impracticality of real-time computation of $p(l_{k})$ during an ongoing episode.
-Table 2: Model score across various environments.
-$\star$ is the baseline agents with a learned VLM-based reward model to compare with.
-BiMI significantly improves performance in Montezuma and Minigrid, while showing mixed results in Crafter due to task-specific characteristics
-Methods Montezuma % vs. $\star$ Minigrid % vs. $\star$ Crafter % vs. $\star$
-Pixl2R $\star$ $0.142\pm 0.003$ $12.4\pm 2.43$ $9.40\pm 0.022$
-Pixl2R + Bi $0.137\pm 0.009$ $-3.5\%$ $31.2\pm 2.04$ $+151\%$ $10.7\pm 0.784$ $+14\%$
-Pixl2R + BiMI $0.162\pm 0.022$ $+14\%$ $37.5\pm 7.83$ $+199\%$ $7.95\pm 0.351$ $-15\%$
-Pixl2R + DEIR $0.176\pm 0.009$ $+23\%$ $17.3\pm 0.51$ $+39\%$ $10.4\pm 1.015$ $+10\%$
-Pixl2R + BiMI + DEIR $0.267\pm 0.016$ $+88\%$ $57.7\pm 2.15$ $+364\%$ $11.0\pm 0.190$ $+17\%$
-ELLM- $\star$ $0.150\pm 0.004$ $19.4\pm 10.06$ $10.8\pm 1.017$
-ELLM- + Bi $0.151\pm 0.016$ $+0.6\%$ $29.7\pm 1.29$ $+53\%$ $11.1\pm 0.601$ $+3.2\%$
-ELLM- + BiMI $0.156\pm 0.014$ $+4.0\%$ $33.6\pm 3.99$ $+74\%$ $9.42\pm 0.267$ $-12\%$
-ELLM + DEIR $0.178\pm 0.029$ $+20\%$ $30.9\pm 3.50$ $+59\%$ $11.8\pm 1.152$ $+9.5\%$
-ELLM- + BiMI + DEIR $0.279\pm 0.078$ $+86\%$ $56.2\pm 6.19$ $+190\%$ $13.1\pm 0.393$ $+22\%$
-Figure 7: BiMI reward showed faster and higher success rates on difficult tasks in Montezuma
 To enhance the stability of the training process, we adopt a linearized version of the mutual information maximization approach, as proposed by Li et al.
@@
 $r^{v}_{\textsc{BiMI}}(\tau,l_{k})=\max(\mathbf{1}_{\{p(l_{k}\mid\tau)\geq\hat{q}\}}-p(l_{k}),0)$ (4)
-It’s important to note that the BiMI approach primarily mitigates false positives (FP) rather than false negatives (FN).
+Note that the BiMI approach primarily mitigates false positives (FPs) rather than false negatives (FN).
 Both Bi and MI aim to reduce the likelihood of rewarding unintended trajectories, thus addressing the FP issue.
 While this conservative approach may increase false negatives by pruning some correct trajectories, this trade-off is beneficial, as demonstrated by both our empirical and theoretical results.
-7 Experiments
-We continue to evaluate using Markovian Pixl2R and Non-Markovian ELLM- and their BiMI-enhanced counterparts, while also exploring potential synergies with the intrinsic reward model DEIR.
-We follow the same experimental setup as in Section 5.1, with additional details in Appendix A.9.
-7.1 Main Results
-In Montezuma, Pixl2R$+$BiMI demonstrated 14% performance increase compared to the original models (see Table 2), which is slightly below our expectations.
+8 Experiments and Results
+We follow the same experimental setup as in Section 6 to ensure a fair comparison.
+Furthermore, we set confidence level for empirical quantile calculation to be $1-\alpha=0.9$.
+We adhered to the standard requirement of limiting the training budget to 1 million frames (Hafner 2022).
+Montezuma  Pixl2R++BiMI demonstrated 14% performance increase compared to the original models (see Table 2), which is slightly below our expectations.
 We attribute this result to BiMI’s intentional strategy of providing less frequent discrete rewards.
-While this strategy effectively reduces false positives, it does not substantially mitigate the inherent reward sparsity issue in Montezuma.
+While this strategy effectively reduces FPs, it does not substantially mitigate the inherent reward sparsity issue in Montezuma.
 However, we discovered a remarkable synergy between BiMI and intrinsic reward models.
@@
 Meanwhile, DEIR’s intrinsic rewards fill the gaps between these anchor points, encouraging the agent to explore efficiently in the interim.
-See Figure 6 for a quantitative analysis: BiMI rewards are now concentrated on key locations.
-A significant improvement is the minimal rewards given for falling off cliffs, which was a common source of false positives in the original model.
+See Figure 6 left and Figure 6 for a quantitative analysis: BiMI rewards are now concentrated on key locations.
+A significant improvement is the minimal rewards given for falling off cliffs, which was a common source of FPs in the original model.
 Figure 7 demonstrates a higher success rate in grabbing the key in the first room, one of the most difficult tasks in Montezuma, highlighting the effectiveness of the proposed reward function and its synergy with intrinsic reward models in guiding agents to solve difficult sparse-reward tasks.
-See results for other environments in Appendix A.10.
-7.2 Overall Performance and Ablation Study
-The overall improvements were substantial.
-As shown in Table 2, BiMI led to a 67% improvement for Pixl2R (Markovian) and a 22% improvement for ELLM- (Non-Markovian).
-These results are also illustrated in Figure 17 in Appendix A.10.
-Our ablation study highlights the distinct contributions of the binary reward (Bi) and Mutual Information (MI) components within the BiMI framework.
-The binary reward mechanism alone accounted for a substantial 36.5% improvement in performance.
-When excluding the results from Crafter, MI component further contributes a 23% improvement over the binary reward alone.
-8 Conclusion
-We reveal two findings in VLM-based reward functions for RL agents: (1) false positive rewards, rather than false negatives, are more detrimental to policy learning; and (2) our proposed BiMI reward function advocates for pessimistic rewarding, significantly mitigating the slowdown in learning caused by false positives.
-Our results are supported by both theoretical analysis and empirical validation across three challenging embodied tasks.
-9 Limitations
-Our study primarily focused on linear sequences of language instructions, excluding more complex cases.
+Minigrid: ELLM-++BiMI achieved a remarkable 74% improvement in performance compared to the original models.This substantial gain is particularly noteworthy given the unique challenges presented by Minigrid.
+The abstract, shape-based visuals diverge drastically from the natural images used in VLMs’ pretraining, preventing the models from effectively utilizing their prior pretraining knowledge.
+Consequently, VLMs struggled to accurately assess similarities between Minigrid’s abstract visuals and textual instructions, resulting in highly noisy reward signals.
+The significant improvement demonstrated by BiMI underscores its effectiveness in handling noisy signals, directly addressing our primary research challenge.
+This capability is crucial for deploying instruction-following agents in real-world, unfamiliar scenarios, where visual inputs often deviate from the VLMs’ training distribution, leading to noisy reward signals.
+Crafter: We observed an intriguing pattern of results.
+The Bi component alone led to 14% and 3.2% improvement in performance over the original models.
+However, contrary to our observations in other environments, the addition of the MI component actually decreased this improvement.
+This unexpected outcome can be attributed to the unique nature of Crafter task, where agents must repeatedly achieve the same subtasks (e.g., drinking water) for survival.
+The MI component, designed to discourage over-reliance on frequently occurring signals, inadvertently penalized the necessary repetition of survival-critical actions.
+Nevertheless, Bi alone still managed to improve performance over vanilla VLM-based reward models, suggesting that reducing FPs is still beneficial across all testing environments.
+The combination of BiMI with DEIR (the intrinsic reward model) also showed promising results, indicating a productive balance between exploration (driven by DEIR) and exploitation (guided by BiMI instruction reward).
+Overall Performance and Ablation Study The BiMI reward function yielded significant performance gains, as detailed in Table 2.
+For the Markovian Pixl2R model, BiMI improved scores by 67%, while the non-Markovian ELLM- model improved by 22%.
+These improvements, alongside synergy with intrinsic rewards, are depicted in Figure 16, which illustrates how combining BiMI with intrinsic rewards (i.e., DEIR) boosts agent performance across different environments.
+Our ablation study further demonstrates the distinct contributions of the binary reward (Bi) and mutual information (MI) components within the BiMI framework with results shown in Figure 17.
+The Bi mechanism alone drove a 36.5% performance increase over baseline models.
+Excluding the Crafter environment, the MI component added a further 23% improvement over Bi alone.
+Together, these results underscore the importance of both components in mitigating false positive rewards and enhancing agent performance.
+9 Conclusion and Discussion
+Most existing work on VLM-based reward models for embodied instruction-following agents remains at the proof-of-concept stage in simplified environments (Du et al.
+2023; Rocamonde et al.
+2024; Wang et al.
+2024), overlooking the challenge of noisy rewards in complex, long-horizon tasks.
+Our work addresses this gap through theoretical analysis grounded in heuristic admissibility — a classical concept from the automated planning domain but rarely applied to embodied learning agents.
+This perspective enables us to rigorously explain why VLM-based reward models are prone to failure without explicit noise handling.
+Building on this foundation, we identify two key insights: (1) false positive rewards are substantially more harmful to policy learning than false negatives, and (2) the proposed BiMI reward function, which incorporates pessimistic rewarding, effectively mitigates this problem.
+Our findings are supported across three challenging embodied tasks, spanning Markovian and non-Markovian reward models.
+Limitations  We primarily focused on linear sequences of language instructions, excluding more complex cases.
 Future research should investigate conditional and ambiguous instructions, which likely introduce additional challenges for VLM-based reward models.
-We also did not explore finetuning the VLM during agent training, a useful strategy as discussed by Fu et al.
-(2024).
-However, we believe that finetuning large VLMs simultaneously during reinforcement learning (RL) is computationally expensive and may not be practical in real-world scenarios.
-We acknowledge that mutual information (MI) has been previously explored as a reward objective in RL.
-However, our work uncovers a novel insight: MI is particularly well-suited for preventing over-reliance on excessively frequent reward sources, which in turn helps to mitigate the issue of too many false positive rewards.
-This research is innovative in explicitly addressing the practical challenges posed by false positive rewards in instruction-following learning agents that utilize VLM rewards.
-We believe that our proposed BiMI reward function represents a novel application of MI specifically tailored for VLM-based rewards.
-There is a gap in providing a rigorous theoretical foundation for why our theoretical findings extend to non-Markovian reward models like ELLM-.
-Our Assumption 4.3 facilitates the integration of non-Markovian reward models into the HuRL framework, though some might argue this assumption is strong.
-However, with advancements in deep RL, the distinction between non-Markovian and Markovian models has become increasingly blurred.
-For instance, the pioneering DQN (Mnih, 2013) utilized sequences of past observations as input, ostensibly making it non-Markovian.
-Yet, with deep learning’s capacity to encapsulate complex state histories into a representation that can be treated as Markovian at a higher abstraction level, some have argued that non-Markovian elements can be effectively reconsidered as Markovian within this new context (Hausknecht and Stone, 2015).
-Our experiments were conducted in extremely sparse reward settings, with tasks intentionally designed to be challenging by requiring long-horizon planning and sequential execution.
-This setup may overemphasize the impact of false positive rewards.
-We acknowledge that in certain cases, rewarding partially correct trajectories may be beneficial for exploration.
-However, to prevent the agent from overvaluing these partial solutions, it is essential that the reward signal remains conservative, providing a pessimistic estimate of the value of partially correct trajectories.
-Appendix A Technical Appendix
-Continued from the main text of Understanding and Mitigating Noise in VLM Rewards, the technical appendix consists of the following:
-  • §A.1 Detailed Problem Setting, which provides a detailed problem setting of the instruction-following reinforcement learning. This is referred to in Section 3 and Section 4.
-  • §A.2 Theoretical Analysis and Proofs, which is referred to in Section 4
-  • §A.3 Additional Notes on Theoretical Setting
-  • §A.4 Implementation Details of the Experiments, which specify the implementation details for both the first and second stage experiments in Section 5.1 and Section 7 respectively.
-  • §A.5 Additional Details of the Experiments of False Positive Rewards, which is referred in Section 5.1.
-  • §A.6 Details of Showing the Prevalence of False Positives in VLM Cosine Similarity Scores, which provides extra figures that cannot be in the main body due to page limit.
-  • §A.7 Impact of Reward Model Quality on Policy Learning Efficiency: Supplementary Experiments and Quantitative Analysis in Montezuma, which provides extra evaluation results to further illustrate false positive rewards hinders agent learning.
-  • §A.8 Pseudo-code for Empirical Quantile Calculation for Binary Signal Threshold, which is referred in Section 6.1.
-  • §A.9 Additional Implementation Details of the Experiments of BiMI Reward Function, which is referred in Section 7.
-  • §A.10 Detailed Experiment Results of BiMI Reward Function, which provides detailed experiment results of BiMI reward function, which is referred in Section 7.
-  • §A.11 Deprecated Convergence Analysis, It is retained here solely for reference and backtracking purposes.
-As the more rigorous theoretical analysis is constructed, the content of this section is now deprecated.
-A.1 Detailed Problem Setting
-A.1.1 The Pointer Mechanism in Reward Machine
-In our problem setting, $r^{v}(\tau_{t},l_{m(t)})$ is the VLM based reward at time step $t$ evaluated using the sub-trajectory $\tau_{t}$ is the sub-trajectory up to time $t$.
-Existing VLM-reward implementations applied a pointer mechanism that decide which instruction should the agent consider at the current step.
-We follow this implementation and denote the pointer as $m(t)$.
-It indicates the current instruction of the agent is trying to complete at time step $t$.
-$m(t)$ is updated according to the following rule:
-$m(t+1)=\begin{cases}1&\text{if $t=0$}\\ m(t)+1&\text{if instr. $l_{m(t)}$ completed at $t$}\\ m(t)&\text{otherwise}\end{cases}$ (2)
-In both the ELLM and the Pixl2R framework that uses VLM-based rewards, the assessment of whether instruction $l_{m(t)}$ is completed at time $t$ is based on the cumulative reward for that instruction.
-The pointer remains on the current instruction until the accumulated reward reaches a predetermined threshold.
-For a detailed pseudo-code of the existing VLM reward + RL algorithm, please refer to Appendix A.4.2.
-Note that our proposed BiMI reward function uses a different mechanism to determine if instruction $l_{m(t)}$ is completed at time $t$.
-In logic, an atomic sentence is a type of statement which cannot be broken down into other simpler sentences.
-While the pointer mechanism effectively enforces the sequential ordering of instructions at a higher level, each individual instruction within the expert walkthrough is not necessarily atomic.
-This means that a single instruction can encapsulate multiple finer-grained requirements for the agent’s behavior.
-Therefore, even though the pointer mechanism is implemented, the internal composition of each instruction is not strictly enforced.
-Consequently, the issue of “composition insensitivity” can arise when the VLM tries to align agents’ trajectories with non-atomic instruction sentences.
-Just as potential-based reward shaping (PBRS) modifies the original MDP into a reshaped MDP through the use of auxiliary rewards, the HuRL framework similarly transforms the original MDP $\mathcal{M}$ into a new MDP ~$\widetilde{\mathcal{M}}=\langle\mathcal{S},\mathcal{A},\mathcal{P},s_{0},\widetilde{r},\widetilde{\gamma}\rangle$, employing a heuristic-based auxiliary reward $h$, along with a coefficient $\beta\in[0,1]$ that scales the auxiliary rewards.
-Thus, the updated reward function can be written as:
-$\widetilde{r}(s,a)=r^{e}(s,a,s^{\prime})+(1-\beta)\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a}[h(s^{\prime})]\quad\text{and $\widetilde{\gamma}=\beta\gamma$}$
-We also provide the formal definition of false negative reward as below.
-Definition A.1 (False Negative Rewards).
-A false negative reward occurs when:
-Instruction-Following Perspective: The VLM-based reward $r^{v}(\tau_{t},l_{m(t)})\approx 0$ for a trajectory $\tau_{t}$ that does satisfy instruction $l_{m(t)}$.
-Heuristic Perspective: The heuristic $h(s_{t})\approx 0<V^{*}(s_{t})$, underestimating the optimal value of $s_{t}$.
-A.1.2 Detailed Terminology
-This section contains detailed terminology from both the standard RL and the HuRL framework.
-We start by setting clear the terminology.
-The original MDP is defined by a tuple $\langle\mathcal{S},\mathcal{A},\mathcal{P},s_{0},r^{e},\gamma\rangle$ as set out in the main text.
-Let $G^{e}(\tau)$ denote the cumulative environment reward of a single trajectory $\tau=\langle s_{0},a_{0},\ldots,s_{T}\rangle$ where $T$ is the total time of the trajectory, we have
-$G^{e}(\tau)=\sum_{t=0}^{T}\gamma^{t}r^{e}(s_{t},a_{t}).$ (3)
-A policy $\pi_{\theta}$ is a conditional distribution $\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})$ parameterized by $\theta$, where $\Delta(\cdot)$ denotes the space of probability distribution.
-The probability of a trajectory $\tau$ being generated, given that actions come from $\pi_{\theta}$ and the starting state is $s_{0}$, is expressed as:
-$P(\tau\mid\pi_{\theta})=\prod_{t=0}^{T}\mathcal{P}(s_{t+1}\mid s_{t},a_{t})\pi_{\theta}(a_{t}\mid s_{t})$ (4)
-Because a policy $\pi$ is stochastic, different trajectories could be generated from the same policy for a given initial state.
-The objective function to be optimized is the expected cumulative environment reward of trajectories generated by $\pi$ starting from state $s_{0}$.
-It is also called the state value function of $s_{0}$ and denoted as:
-$V^{\pi}(s_{0}):=\mathbb{E}_{\rho_{s}^{\pi}}\left[G^{e}(\tau)\right]=\sum_{\textnormal{all $\tau_{i}$ generated by $\pi_{\theta}$}}G^{e}(\tau_{i})\cdot P(\tau_{i}|\pi)$ (5)
-where $\rho_{s}^{\pi}$ denotes the trajectory distribution of $s_{0},a_{0},s_{1},...$ induced by running $\pi$ starting from $s_{0}=s$. We defined an acceptable policy $\pi_{A}$ as
-$\pi_{A}:V^{\pi_{A}}\geq V^{\pi^{*}}-\epsilon$ (6)
-where $\pi^{*}$ is the optimal policy and $\epsilon$ is a control parameter that quantifies the acceptance of other policies.
-Acceptable policies defined in this way are also referred to as $\epsilon$-optimal policies.
-We set $V^{*}=V^{\pi^{*}}$ for easy reference.
-Any policy that is not an acceptable policy is an unacceptable policy $\pi_{U}$.
-We define a good trajectory $\tau_{G}$ as
-$\tau_{G}:G^{e}(\tau_{G})\geq V^{*}-\epsilon.$ (7)
-Any trajectory that is not a good trajectory is a bad trajectory $\tau_{B}$.
-Note that this way of distinguishing good/bad trajectories and acceptable/unacceptable policies is explicitly based on the environment reward $r^{e}$, and introducing additional or alternative rewards into the system does not change this definition.
-For easy reference, we call any trajectory that reaches a goal state a goal trajectory, and we call any policy that can generate at least one goal trajectory a goal policy.
-We denote the state distribution of a policy $\pi$ at time $t$ as $d^{\pi}_{t}$.
-Thus, the discounted average state distribution of a policy $\pi$ can be expressed as $d^{\pi}=(1-\gamma)\sum^{\infty}_{t=0}\gamma^{t}d^{\pi}_{t}$, where $(1-\gamma)$ is the normalization factor to ensure the result is a proper probability distribution.
-We also define a shorthand notation – for a state distribution $d\in\Delta(\mathcal{S})$, we define $V(d)=\mathbb{E}_{s\sim d}\left[V(s)\right]$.
-When the Heuristic-Guided Reinforcement Learning (HuRL) framework is applied, the reshaped MDP $\widetilde{\mathcal{M}}$ has separated reshaped reward $\widetilde{r}$ and separated discount factor $\widetilde{\gamma}=\beta\gamma$, where $\beta\in[0,1]$ is the hyperparameter that scales the auxiliary heuristic rewards.
-We denote the value function of the reshaped MDP as $\widetilde{V}$, and the optimal policy and its value function under $\widetilde{\mathcal{M}}$ as $\widetilde{\pi}^{*}$ and $\widetilde{V}^{*}(\widetilde{V}^{\widetilde{\pi}^{*}})$.
-The discussion regarding the convergence guarantee of learning $\pi$ in $\widetilde{\mathcal{M}}$ resulting in the optimal policy in the original MDP $\mathcal{M}$ can be found in Appendix A.3.1.
-For the purposes of this paper, it is sufficient to note that the authors of HuRL have provided a trick to ensure convergence.
-Moreover, empirical experiments conducted by the authors demonstrate that HuRL converges even without this trick.
-It is also important to distinguish the Bellman backup equation under the two different MDPs: the original $\mathcal{M}$ and the reshaped $\widetilde{\mathcal{M}}$.
-By definition, $(\mathcal{B}h)(s,a)=r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[h(s^{\prime})]$.
-In contrast, (~$(\widetilde{\mathcal{B}}h)(s,a)=\widetilde{r}(s,a)+\widetilde{\gamma}\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[h(s^{\prime})]$.
-Nevertheless, the two Bellman backup equations possess a remarkable property – they are essentially equivalent to each other:
-(~$(\widetilde{\mathcal{B}}h)(s,a)$ $=\widetilde{r}(s,a)+\widetilde{\gamma}\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[h(s^{\prime})]$
-$=\left(r(s,a)+(1-\beta)\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[h(s^{\prime})]\right)+\beta\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[h(s^{\prime})]$
-$=r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[h(s^{\prime})]$
-$=(\mathcal{B}h)(s,a)$ (8)
-A.1.3 Detailed Construction of Markovian Heuristic using Non-Markovian Rewards from VLM
-Specifically,
-$h(s_{t}):=\begin{cases}A(s_{t})V^{*}(s_{t})&\text{if $s_{t}\in S^{m(t)}$}\\ 0&\text{otherwise}\end{cases}$ (9)
-where
-$A(s_{t}):=\frac{\Pi_{t^{\prime}=0}^{t}r^{v}(\tau_{t^{\prime}},l_{m(t^{\prime})})}{\Pi_{t^{\prime}=0}^{t}r^{v}(\tau^{*}_{t^{\prime}},l_{m(t^{\prime})})}\in[0,1]$ (10)
-is a scaling factor that quantifies how closely the agent is following previous instructions.
-A.2 Theoretical Analysis and Proofs
-A.2.1 Convergence time on a sparse-reward landscape
-We can pin down a characteristic convergence time on a sparse-reward landscape. The sparse-reward setting enforces that
-$r^{e}(s_{t})=\begin{cases}1&\text{if $s_{t}\in\mathcal{S}_{G}$}\\ 0&\text{otherwise}\end{cases}.$ (11)
-It is clear from the definition that a good trajectory must always be a goal trajectory, and the cumulative reward is simply the reward at the final goal state so $G^{e}(\tau)=\gamma^{T}$ in this setting.
-For a randomly initialized policy, it is highly unlikely that the initial distribution of trajectories contains any goal trajectory due to the sparsity of goal states.
-The optimization of $V^{e}_{\pi_{\theta}}$ thus consists of two parts.
-The first part is to search for a goal trajectory.
-The gradient landscape is almost 0 everywhere, except for cases where a trajectory is $\delta$-close to a goal trajectory.
-Here $\delta$ is the differential unit in the numerical differentiation used in the gradient calculation
-$\theta=\theta+\alpha\nabla_{\theta}V_{\pi_{\theta}}$ (12)
-such that $\delta$-close means being numerically accessible within a distance of $|\delta|$ in parameter space.
-And the second part is to reduce $T$ so that goal trajectories become good trajectories and consequently achieving acceptable policies.
-For the first part, searching for a trajectory for the target is effectively a random walk in the $d$-dimensional parameter space due to the flat gradient landscape.
-Lemma A.2.
-For a random walk in $n$-dimensional space, the expected number of steps $T_{D}$ needed to travel a distance of $D$ scales with $D^{2}$.
-Proof: let $\vec{X_{1}}$, $\vec{X_{2}}$, … $\vec{X_{T}}$ be IID random unit vectors uniformly distributed on a $(d-1)$-dimensional sphere $\mathcal{S}^{d-1}\subset\mathbb{R}^{d}$, where $\vec{X_{i}}=(X_{i1},...,X_{id})$ and $|\vec{X_{i}}|^{2}=\sum_{j=1}^{d}X^{2}_{ij}=1$.
-Let $\vec{S_{T}}:=\sum_{i=1}^{T}\vec{X_{i}}$.
-By a $n$-dimensional cosine rule we have
-$|\vec{S}_{T}|^{2}$ $=|\vec{S}_{T-1}|^{2}+2\vec{S}_{T-1}\cdot\vec{X}_{T}+|\vec{X}_{T}|^{2},$ (13)
-and because $\mathbb{E}[\vec{X}_{T}]=\vec{0}$
-$\mathbb{E}[|\vec{S}_{T}|^{2}]$ $=\mathbb{E}[|\vec{S}_{T-1}|^{2}]+\mathbb{E}[2\vec{S}_{T-1}\cdot\vec{X}_{T}]+1$ (14)
-$=\mathbb{E}[|\vec{S}_{T-1}|^{2})]+2\vec{S}_{T-1}\mathbb{E}[\vec{X}_{T}]+1$ (15)
-$=\mathbb{E}[|\vec{S}_{T-1}|^{2}]+1$ (16)
-$=\mathbb{E}[|\vec{S}_{T-2}|^{2}]+1+1$ (17)
-$\vdots$ (18)
-$=T$ (19)
-i.e.
-$\mathbb{E}[|\vec{S}_{T}|]\sim\sqrt{T}$.
-When a policy is randomly initialized with $\theta_{0}$, the distance $D$ to a goal policy, $D:=||\theta_{goal}-\theta_{0}||$ is fixed and is the distance the random walk needs to travel ($\mathbb{E}[|\vec{S}_{T}|]=D\sim\sqrt{T_{D}}$), so the characteristic time needed to travel this distance $T_{D}\sim D^{2}$ as we have shown above.
-See 4.1 Auxiliary rewards essentially open the path for a divide-and-conquer approach by introducing intermediate rewards in the learning process. We introduce BiMI rewards as
-$r^{v}_{\textsc{BiMI}}(\tau,l_{k})=\max(\mathbf{1}_{\{p(l_{k}\mid\tau)\geq\hat{q}\}}-p(l_{k}),0)$ (20)
-and the cumulative reward of a single trajectory in the presence of BiMI rewards becomes
-$G^{v}_{\textsc{BiMI}}(\tau)=\sum_{t=0}^{T}\gamma^{t}r^{v}_{\textsc{BiMI}}(\tau,l_{m(t)}).$ (21)
-[…]
-See 4.2 Proof: the expected time taken for each of the sub-tasks then scales with ${}d_{i}^{2}$ respectively (Lemma A.2), and we have
-$d_{1}^{2}+d_{2}^{2}+...+d_{n-1}^{2}$ $<(d_{1}+d_{2}+...+d_{n-1})^{2}$ (22)
-$\mathbb{E}\left[\sum_{i=1}^{n-1}T_{d_{i}}\right]$ $<\mathbb{E}[T_{D}]$ (23)
-because $d_{i}>0\forall i$.
-We can also work out that the upper bound for this improvement is a factor of $n-1$ by invoking the Cauchy–Schwarz inequality $\left(\sum_{i=1}^{n}u_{i}^{2}\right)\left(\sum_{i=1}^{n}v_{i}^{2}\right)\geq\left(\sum_{i=1}^{n}u_{i}v_{i}\right)^{2}$:
-$(d_{1}^{2}+d_{2}^{2}+...+d_{n-1}^{2})(1^{2}+1^{2}+...+1^{2})$ $\geq(d_{1}+d_{2}+...+d_{n-1})^{2}$ (24)
-$(d_{1}^{2}+d_{2}^{2}+...+d_{n-1}^{2})$ $\geq\frac{1}{n-1}D^{2}$ (25)
-$\mathbb{E}\left[\sum_{i=1}^{n-1}T_{d_{i}}\right]\geq\frac{1}{n-1}\mathbb{E}[T_{D}]$ (26)
-and the equality sign holds (indicating maximal improvement) when $d_{1}=d_{2}=...=d_{n-1}$.
-The intuitive interpretation is that the divide-and-conquer approach is the most effective when the task is divided evenly into subtasks.
-Additionally, the presence of $p(l_{k})$ introduces some lexical-level tolerance in the learning process and results in a considerably larger $\delta$-close radius which could result in further improvement in the convergence rate.
-However, as we will discuss in the next section, introducing too much tolerance could harm the convergence rate rather than improve it.
-A.2.2 False Positive and the Violation of Pessimistic Property of Heuristics
-See 4.7
-Proof.
-We want to show that given:
+There is also a gap in providing a rigorous theoretical foundation for why our theoretical findings extend to non-Markovian reward models.
+However, with advancements in deep RL, the distinction between non-Markovian and Markovian models has become increasingly blurred (Hausknecht and Stone 2015).
+Appendix Table of Contents
   1. 1.
-    for all successor states $s^{\prime}$: $h(s^{\prime})\leq V^{*}(s^{\prime})$ for an arbitrary $s$.
+    Anticipatory defense on the choice of test environments and metrics (Appendix A)
+    Justification for chosen test environments and evaluation metrics.
   2. 2.
-    false positive at current arbitrary state $s$: $V^{*}(s)<h(s)$.
-The objective is to show that under the above conditions, we obtain $\max_{a}(\mathcal{B}h)(s,a)<h(s)$.
-  1. 1.
-    Express the Bellman backup for $h$: $(\mathcal{B}h)(s,a)=R(s,a)+\gamma\sum_{s^{\prime}}P(s^{\prime}|s,a)h(s^{\prime})$
-  2. 2.
-    Express the Bellman version of $V^{*}$: $V^{*}(s)=\max_{a\in\mathcal{A}}\left[R(s,a)+\gamma\sum_{s^{\prime}}P(s^{\prime}|s,a)V^{*}(s^{\prime})\right]$.
-    Given that $h(s^{\prime})\leq V^{*}(s^{\prime})$ for all $s^{\prime}$, we can expand it to $\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)h(s^{\prime})\leq\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)V^{*}(s^{\prime})$.
-Therefore, we have:
-    $R(s,a)+\gamma\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)h(s^{\prime})\leq R(s,a)+\gamma\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)V^{*}(s^{\prime})=Q^{*}(s,a)$ (27)
+    Detailed Related Work (Appendix B)
+    An expanded discussion of related research in the domain and research gaps.
   3. 3.
-    Taking the maximum over actions to both sides, we have:
-    $\max_{a}\left(R(s,a)+\gamma\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)h(s^{\prime})\right)$ $\leq max_{a}Q^{*}(s,a)$ (28)
-    $\max_{a}(\mathcal{B}h)(s,a)$ $\leq V^{*}(s)$ (29)
+    The Complete Procedure of the VLM-based Reward Model (Appendix C)
+    The complete procedure of involving VLM-based reward model in the RL training loop.
   4. 4.
-    Apply false positive condition: Given $V^{*}(s)<h(s)$, substituting into the above inequality, we have $\max_{a}(\mathcal{B}h)(s,a)<h(s)$
-Implication: Maintaining a pessimistic heuristic is inherently fragile because the introduction of a false positive in any state disrupts the pessimistic condition. ∎
-See 4.8 We analyze the convergence speed of policy learning by examining the performance gap, which is defined as the difference between the optimal value of the initial state $s_{0}$, $V^{*}(s_{0})$, and the value of the initial state under an arbitrary policy $\pi$, $V^{\pi}(s_{0})$.
-Specifically, we focus on deriving an upper bound for this performance gap.
-The key intuition is that a smaller upper bound implies faster convergence to the optimal policy, as fewer iterations of policy updates will be required to reach the optimum.
-We begin by stating the theorem made by HuRL authors:
-Theorem A.3 (Performance Gap Decomposition (Cheng et al., 2021)).
-For any policy $\pi$, heuristic $h:\mathcal{S}\rightarrow\mathbb{R}$, and mixing coefficient $\beta\in[0,1]$,
-$V^{*}(s_{0})-V^{\pi}(s_{0})=\mathrm{Regret}(h,\beta,\pi)+\mathrm{Bias}(h,\beta,\pi)$ (30)
-where the regret and the bias term are expressed as follows:
-$\mathrm{Regret}(h,\beta,\pi)$ $\coloneqq\beta\left(\widetilde{V}^{*}(s_{0})-\widetilde{V}^{\pi}(s_{0})\right)+\frac{1-\beta}{1-\gamma}\left(\widetilde{V}^{*}(d^{\pi})-\widetilde{V}^{\pi}(d^{\pi})\right)$ (31)
-$\mathrm{Bias}(h,\beta,\pi)$ $\coloneqq\left(V^{*}(s_{0})-\widetilde{V}^{*}(s_{0})\right)+\frac{\gamma(1-\beta)}{1-\gamma}\mathbb{E}_{s,a\sim d^{\pi}}\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}\left[h(s^{\prime})-\widetilde{V}^{*}(s^{\prime})\right]$ (32)
-We will not provide a proof for Theorem A.3 in this paper.
-Please refer to (Cheng et al., 2021) for details.
-The theorem demonstrates that the performance gap can be elegantly decomposed into two components: a regret term and a bias term such that:
-  1. 1.
-    The regret term quantifies the difference between $\widetilde{V^{*}}$ and $\widetilde{V^{\pi}}$, representing the error caused by $\pi$ being suboptimal in the reshaped MDP $\widetilde{\mathcal{M}}$.
-Since $\pi$ is trained by our selected RL algorithm directly on the reshaped MDP $\widetilde{\mathcal{M}}$, the primary responsibility for minimizing this regret term falls to the RL algorithm itself, not to the design of the auxiliary reward signal.
-Thus, when evaluating the effects of false positive or false negative rewards, we choose not to focus on bounding the regret term.
-  2. 2.
-    The bias term captures two key discrepancies: first, between the true optimal value function $V^{*}$ of the original MDP $\mathcal{M}$ and the optimal value function $\widetilde{V}^{*}$ of the reshaped MDP $\widetilde{\mathcal{M}}$; second, between $\widetilde{V}^{*}$ and the heuristic $h(s)$.
-This term, therefore, reflects the error introduced by addressing the reshaped MDP instead of the original one, alongside how well the heuristic $h$ approximates the optimal value function in the reshaped MDP.
-Consequently, this bias term directly relates to the quality of the heuristic reward signal $h$.
-Hence, we focus on analyzing its upper bound to assess the impact of false positive or false negative rewards on this heuristic.
-Cheng et al. (2021) have further derived the upper bound for the bias term, which we present here as a lemma. We omit the proof for brevity; for detailed proof, please refer to (Cheng et al., 2021).
-Lemma A.4 (Upper Bound of the Bias Term).
-$\mathrm{Bias}(h,\beta,\pi)\leq(1-\beta)\gamma\left(\mathbb{E}_{\rho^{\pi^{*}}}\left[\sum_{t=1}^{\infty}(\beta\gamma)^{t-1}(V^{*}(s_{t})-h(s_{t}))\right]+\mathbb{E}_{\rho^{\pi}}\left[\sum_{t=1}^{\infty}\gamma^{t-1}(h(s_{t})-\widetilde{V}^{*}(s_{t}))\right]\right)$ (33)
-where $\rho^{\pi}$ denotes the trajectory distribution of $s_{0},a_{0},s_{1},...$ induced by running $\pi$ starting from $s_{0}$.
-This upper bound elegantly illustrates a trade-off: if the heuristic $h$ is set too high (overestimation error), it reduces the first term but increases the second term.
-Conversely, if $h$ is set too low (underestimation error), it increases the first term but reduces the second term.
-Just by inspection, it is not immediately clear whether overestimation or underestimation results in a better upper bound, as both impact the bias term in opposing ways.
-We now prove that underestimation of $h$ (pessimistic $h$) is better than overestimation (i.e., $\exists s_{t}\in\mathcal{S},h(s_{t})>V^{*}(s_{t})$) for minimizing bias.
-Proof.
-  1. 1.
-    Define Terms:
-    • Let $B_{1}=\mathbb{E}_{\rho^{\pi^{*}}}\left[\sum_{t=1}^{\infty}(\beta\gamma)^{t-1}(V^{*}(s_{t})-h(s_{t}))\right]$
-    • Let $B_{2}=\mathbb{E}_{\rho^{\pi}}\left[\sum_{t=1}^{\infty}\gamma^{t-1}(h(s_{t})-\widetilde{V}^{*}(s_{t}))\right]$
-  2. 2.
-    Underestimation Case (Pessimistic $h$):
-    Lemma A.5.
-    If $h$ is pessimistic with respect to $\mathcal{M}$, $\forall\beta\in[0,1],s\in\mathcal{S},\widetilde{V}^{*}(s)\geq h(s)$
-    Proof.
-    To begin with, we need to use another lemma from Cheng et al. (2021), shown as follows:
-    Lemma A.6 (Bellman backup of reshaped MDP (Cheng et al., 2021)).
-    For any policy $\pi$, we have
-    $\widetilde{V}^{\pi}(s_{0})-h(s_{0})=\frac{1}{1-\beta\gamma}\mathbb{E}_{\widetilde{d}_{s_{0}}^{\pi}}[(\widetilde{\mathcal{B}}h)(s,a)-h(s)]$ (34)
-    where $\widetilde{d}_{s_{0}}^{\pi}$ refers to the discounted average state distribution of policy $\pi$ in reshaped MDP $\widetilde{\mathcal{M}}$.
-It can be expressed as $\widetilde{d}_{s_{0}}^{\pi}=(1-\widetilde{\gamma})\sum_{t=0}^{\infty}\widetilde{\gamma}^{t}d_{t}^{\pi}$, and $d_{t}^{\pi}$ is the state distribution of policy $\pi$ at time $t$ with $d_{0}^{\pi}=\mathbf{1}\{s=s_{0}\}$.
-    For brevity, we do not include the proof for Lemma A.6.
-    First of all, due to the definition of optimal value $V^{*}$, we have
-    $\widetilde{V}^{*}(s_{0})\geq\widetilde{V}^{\pi}(s_{0})$ (35)
-    Then, according to Lemma A.6, we can get
-    $\widetilde{V}^{*}(s_{0})\geq\widetilde{V}^{\pi}(s_{0})=h(s_{0})+\frac{1}{1-\beta\gamma}\mathbb{E}_{\widetilde{d}_{s_{0}}^{\pi}}[(\widetilde{\mathcal{B}}h)(s,a)-h(s)]$ (36)
-    Let $\pi$ denote the greedy policy of $\arg\max_{a}(\mathcal{B}h)(s,a)$ and then trace back to Equation 8, we have (~$(\widetilde{\mathcal{B}}h)(s,a)=(\mathcal{B}h)(s,a)$, that means
-    $\widetilde{V}^{*}(s_{0})\geq\widetilde{V}^{\pi}(s_{0})$ $=h(s_{0})+\frac{1}{1-\beta\gamma}\mathbb{E}_{\widetilde{d}_{s_{0}}^{\pi}}[(\widetilde{\mathcal{B}}h)(s,a)-h(s)]$ (37)
-    $=h(s_{0})+\frac{1}{1-\beta\gamma}\mathbb{E}_{\widetilde{d}_{s_{0}}^{\pi}}[(\mathcal{B}h)(s,a)-h(s)]$ (38)
-    $\geq h(s_{0})\quad\text{[direct result of $h$ being pessimistic]}$] (39)
-    Generalization to Any State: The lemma from Cheng et al.
-(2021) (Lemma A.6) is stated in terms of any starting state $s_{0}$.
-Therefore, we can replace $s_{0}$ with any state $s\in\mathcal{S}$ in our analysis.
-Thus we get $\forall s\in\mathcal{S},\widetilde{V}^{*}(s)\geq h(s)$
-    Lemma A.5 implies that when $h$ is pessimistic, $B_{2}=\mathbb{E}_{\rho^{\pi}}\left[\sum_{t=1}^{\infty}\gamma^{t-1}(h(s_{t})-\widetilde{V}^{*}(s_{t}))\right]\leq 0$ because $\widetilde{V}^{*}(s)\geq h(s)$ for all $s$.
-Therefore, the bias error is only bounded by $B_{1}$ when $h$ is pessimistic, i.e., $\mathrm{Bias}(h,\beta,\pi)\leq B_{1}$.
-  3. 3.
-    Overestimation Case ($\exists s_{t}\in\mathcal{S},h(s_{t})>V^{*}(s_{t})$):
-    We show briefly that, unlike the underestimation error, the overestimation error does not have a closed-form upper bound expression.
-The difficulty, as pointed out by (Cheng et al., 2021), originates from the trajectory-dependence on $\mathbb{E}_{\rho^{\pi}}[\cdot]$ as the $l_{\infty}$ approximation error here can be difficult to control in large state spaces.
-It is possible that, upon picking up falsely high $h$ states, $\rho^{\pi}$ gets further distorted away from $\rho^{\pi^{*}}$ and accumulates more falsely high $h$ states.
-In other words, this trajectory dependence makes $B_{2}$ prone to a feedback loop of accumulating overestimation errors and resulting in a much larger upper bound for the bias.
-A.3 Additional Notes on Theoretical Setting
-A.3.1 Convergence Guarantee of HuRL Compared With Potential-Based Reward Shaping
-Previous work on potential-based reward shaping (PBRS) has established that learning a policy on a reshaped MDP can converge to the optimum of the original MDP, as proved by Ng et al.
-(1999).
-They proved that the optimal $Q$ value of the reshaped MDP is equivalent to the optimal $Q$ value of the original MDP minus a state-dependent function.
-Consequently, for an optimal policy defined as $\pi^{*}=\arg\max_{a\in\mathcal{A}}Q^{*}(s,a)$, this additional state-dependent value does not alter the policy.
-Thus, the optimal policy of the original MDP $\mathcal{M}$ is also the optimal policy for the reshaped $\widetilde{\mathcal{M}}$.
-However, the theoretical convergence for the reshaped MDP within the framework of Heuristic-Guided Reinforcement Learning (HuRL) has not been proven with the same rigor.
-The authors of HuRL made a trick which involves manipulating the coefficient $\beta$, which scales the heuristic-based auxiliary rewards.
-If $\beta$ increases gradually from 0 to 1 over the training process, the agent effectively transitions from interacting with the modified MDP back to the original MDP.
-This method ensures that, at the end of training, the agent is exactly solving the original problem.
-Moreover, according to the Blackwell optimal property (Chen et al., 2018), convergence can occur before $\beta$ reaches 1, thereby allowing the trained policy to maintain optimality in the original MDP under HuRL.
-However, in the original HuRL paper, among the five test environments, the $\beta$ hyperparameter was actually fixed for four of them.
-Surprisingly, HuRL still converged to the optimal policy faster than the original MDP without updating $\beta$.
-This observation poses an enigma within the HuRL framework, as it is unclear why this convergence to the optimum of the original MDP occurs without dynamically adjusting $\beta$.
-Thus, while the practical success of HuRL in these environments is evident, the underlying reasons for such convergence remain to be fully understood.
-This discrepancy between theoretical expectation and empirical results leaves room for future research to explore why HuRL converges under fixed $\beta$.
-Nevertheless, proving the convergence properties goes beyond the scope of our paper, which aims to highlight the prevalence of false positive rewards and their impact.
-An alternative perspective on this issue is to consider the learning process of maximizing auxiliary rewards as akin to performing behavior cloning from experts.
-In this sense, we do not need to focus on convergence within the current MDP but can view it as a form of pretraining.
-A.4 Implementation Details of the Experiments
-A.4.1 Environment Details
+    Convergence time on a sparse-reward landscape (Appendix D)
+    Proof of convergence time on a sparse-reward landscape.
+  5. 5.
+    VLM-based reward model is a heuristic function (Appendix E)
+    Proof that VLM-based reward model is a heuristic function.
+  6. 6.
+    False Positive and the Violation of Pessimistic Property of Heuristics (Appendix F)
+    Proof that false positive rewards violate the pessimistic property of heuristics.
+  7. 7.
+    False Positive gives unbounded bias in the optimality gap (Appendix G)
+    Proof that false positive rewards give unbounded bias in the optimality gap.
+  8. 8.
+    Additional Experimental Setup Details (Appendix H)
+    Supplementary information about experimental configurations.
+  9. 9.
+    Details of Showing the Prevalence of False Positives in VLM Cosine Similarity Scores (Appendix I)
+    Extra figures for stage 1 experiments.
+  10. 10.
+    Extra Experiment on Simulated Oracle Reward Model to demonstrate H3 and H4 (Appendix J)
+    Supporting experiments for hypothesis H3 and H4 using simulated oracle reward models.
+  11. 11.
+    Algorithm for Empirical Quantile Calculation (Appendix K)
+    Algorithm for calculating empirical quantile of cosine similarity scores.
+  12. 12.
+    Ablation Study Lineplots for BiMI reward (Appendix L)
+    Lineplots for ablation study of BiMI reward function.
+  13. 13.
+    Encoder-based Vision-Language Models Explanation (Appendix M)
+    A brief overview of two types of VLMs, encoder-based and generative VLMs.
+  14. 14.
+    Sparse Reward and Random Walk (Appendix N)
+    Proposition and proof of the relationship between RL convergence in sparse reward setting and random walk.
+Appendix A Anticipatory defense on the choice of test environments and metrics
+As discussed in Appendix B, we need test environments that clearly demonstrate the negative impact of noisy reward signals from VLM-based reward models.
+This requires complex state spaces (both large and diverse) where the approximation errors of VLMs become evident — specifically, where they struggle to provide accurate alignment scores between an agent’s trajectory and given instructions due to the complexity of visual observations.
+Additionally, we need longer-horizon tasks where the impact of noisy reward signals can accumulate over time.
+In contrast, many embodied environments used in previous studies are not suitable due to their short-horizon tasks and simple state spaces.
+Table 3 lists common RL embodied environments used in the literature, explaining why certain environments are inadequate for our evaluation purposes while highlighting suitable environments we aim to use.
+Table 3: Common RL embodied environments and their suitability for evaluating noisy reward signals from VLM-based reward models.
+Env Type State Space Complexity Horizon Length Instruction Source
+Not Suitable Envs
+CartPole (Brockman et al. 2016) Classic Control Simple Short Manual annotation
+MountainCar (Brockman et al. 2016) Classic Control Simple Short Manual annotation
+Humanoid (Todorov, Erez, and Tassa 2012) Bipedal robot control Complex Short Manual annotation
+MetaWorld (Yu et al. 2019) Robotic arm Moderate Short Predefined task descriptions
+SoftGym (Lin et al. 2020) Robotic arm Moderate Short Predefined task descriptions
+Suitable Envs
+Montezuma (Bellemare et al. 2013) Platformer Complex Long Manual annotation
+Crafter (Hafner 2022) Open world survival Complex Long Programmable (via engine data)
+Minigrid (Chevalier-Boisvert et al. 2023) Grid-world maze Controllable Controllable Predefined task descriptions
 We describe each testing environment used in our experiments.
-More details introduction can be found in on the official project homepage of each benchmark (Chevalier-Boisvert et al., 2018; Bellemare et al., 2013; Hafner, 2021).
+More details introduction can be found in on the official project homepage of each benchmark (Hafner 2022; Bellemare et al.
+2013; Chevalier-Boisvert et al.
+2023).
   • Crafter features randomly generated 2D worlds where the player needs to forage for food and water, find shelter to sleep, defend against monsters, collect materials, and build tools.
@@
 The instruction complexity is set to 3, meaning there are at least 3 target objects to interact with in a specific order.
+Figure 8: Illustration of the Crafter task. The agent must survive as long as possible and explore for new crafting recipes.
 Figure 9: Illustration of the Montezuma’s Revenge task. The agent must navigate through a series of rooms to collect treasures and keys.
 Figure 10: Illustration of the Minigrid ‘Go to seq’ task. The agent must navigate through a sequence of rooms and touch target objects in the correct order.
-Figure 10: Illustration of the Crafter task. The agent must survive as long as possible and explore for new crafting recipes.
-A.4.2 Instruction-Guide Procedure Details
-The VLM-based reward model will have a pointer to the sequence of the instruction sentence, starting at the first sentence.
-For original models Pixl2R and ELLM-, we follow the setting in there original work where for each instruction sentence (yes, the full instruction essay will be split into multiple sentences and treat each sentence as atomic instruction $l_{k}$), the reward model will have a maximum cap of rewards (2.0) it can assign to the agent in one episode.
-When the cap is reached, the reward model will move its pointer to the next instruction sentence.
-For the BiMI reward model, the reward model will move its pointer to the next instruction sentence when the binary signal is triggered.
-Below is the pseudo-code for the instruction-following RL training procedure for both Pixl2R and ELLM- models.
-Algorithm 1 ELLM and Pixl2R pseudo-code for instruction-following RL
-1:Initialize policy network $\pi_{\theta}$
-2:Initialize value network $V_{\phi}$
-3:Setup VLM-based reward model $E(\cdot)$
-4:Split instruction essay into sentences $\{l_{1},l_{2},...,l_{K}\}$
-5:Initialize instruction pointer $p=1$
-6:Initialize cumulative VLM reward $r_{\text{cum}}=0$
-7:Initialize cumulative VLM reward threshold $q$
-8:Initialize replay buffer $\mathcal{D}$
-9:Initialize agent trajectory memory queue $\tau$ with length $W$Foreach episode
-[…]
-11:Select action $a_{t}\sim\pi_{\theta}(a_{t}|s_{t})$
-12:Execute $a_{t}$, observe next state $s_{t+1}$ and extrinsic reward $r^{e}_{t}$
-13:Enqueue $(s_{t},a_{t},r_{t},s_{t+1})$ in $\tau$
-14:Compute VLM reward:
-15:$r^{v}_{t}=\frac{E(\tau)\cdot E(l_{p})}{\|E(\tau)\|\|E(l_{p})\|}$triangleright Apply BiMI here if needed
-16:Combine rewards: $r_{t}=r^{e}_{t}+(1-\beta)\gamma r^{v}_{t}$triangleright $\beta$ is a scaling factor
-17:Store $(s_{t},a_{t},r_{t},s_{t+1})$ in $\mathcal{D}$
-[…]
-19:$p\leftarrow\min(p+1,K)$
-[…]
-21:Sample mini-batch $\{(s_{j},a_{j},r_{j},s_{j+1})\}$ from $\mathcal{D}$
-22:Compute TD errors:
-23:$\delta_{j}=r_{j}+\gamma V_{\phi}(s_{j+1})-V_{\phi}(s_{j})$
-24:Update value network:
-25:$\phi\leftarrow\phi+\alpha_{v}\sum_{j}\delta_{j}\nabla_{\phi}V_{\phi}(s_{j})$
-26:Update policy network:
-[…]
-A.4.3 Finetuning VLM-based Reward Models
-In contrast to previous work on instructing following RL where they rely on hand-crafted oracle multimodal reward models, we use actual pretrained VLMs to generate reward signals.
-2 VLM backbone models are used in our experiments: 1) CLIP (Radford et al., 2021), pretrained by image-text pairs; and (2) X-CLIP (Ma et al., 2022), pretrained by video-text pairs.
-In particular, Pixl2R uses CLIP because it only uses the single latest frame as input.
-In contrast, ELLM- takes a slice of trajectory (i.e., multiple frames) as input, and thus uses either X-CLIP or CLIP with additional RNN encoder as the reward model.
-Due to the cartoonish and abstract visuals of the testing environments, we need to further fine-tune the VLMs to adapt to this new visual domain.
-We use well-trained expert agents based on Moon et al.
-(2023) to generate expert trajectories for the Crafter environments and annotate them with instructions using internal information from the game engine.
-For Minigrid environments, we use classical search-based planning robots to generate expert trajectories and annotate them with the corresponding task instructions.
-For Montezuma’s Revenge, we manually annotate the expert trajectories.
-For Minigrid and Crafter, we have 80,000 training pairs, while for Montezuma’s Revenge, we have around 300 training pairs.
-These training data are of high quality, as we have made every effort to avoid false positive rewards due to poor training data quality.
-To enhance our models’ robustness, we also employed contrastive learning techniques during VLM training, utilizing similar manipulated data as hard negatives.
-However, despite the fine-tuning process, false positive rewards remain unavoidable.
-Figure 11: Example of training data for the Montezuma environment.
-We used the threshold $\hat{q}$ introduced in Section 6.1 to make binary classification on the testing pairs to evaluate the performance of the fine-tuned VLM-based reward models.
-We found that VLM models had difficulty achieving high accuracy on Minigrid environment, which is likely due to the too abstract and cartoonish nature of the environment, causing the VLMs to struggle to learn the visual-textual correspondence.
-We also found that X-CLIP did not perform better than CLIP in our experiments.
-We hypothesize that the cartoonish nature of the testing environments may have caused the X-CLIP model to struggle to learn the visual-textual correspondence.
-Thus, we used CLIP as the backbone model throughout our following experiments.
-The performance of the fine-tuned VLM-based reward models is shown in Table 3.
-Even when the precision score reaches 0.98, indicating that only 2% of the rewards are false positives in the validation set, the agent can still significantly underperform in the testing environments.
-The core issue is that in out-of-distribution (O.O.D.) testing environments, false positive rewards are prevalent and inevitable.
-Therefore, it is crucial to design a reward function that is robust to reward noise.
-Table 3: Performance of fine-tuned VLM reward model on the testing dataset using the 90th percentile empirical quantile as threshold
-Environment Precision Accuracy F1 Score Recall Model
-Crafter 0.9847 0.9466 0.8538 0.9702 CLIP ELLM-
-Crafter 0.9799 0.9028 0.7618 0.9842 CLIP Pixl2R
-Crafter 0.2095 0.2514 0.2868 0.9657 XCLIP ELLM-
-Minigrid 0.7260 0.9200 0.7849 0.9763 CLIP ELLM-
-Minigrid 0.6992 0.9086 0.7592 0.9616 CLIP Pixl2R
-Minigrid 0.1716 0.2310 0.2642 0.9704 XCLIP ELLM-
-Montezuma 0.8838 0.9638 0.8825 0.9478 CLIP ELLM-
-Montezuma 0.8343 0.9108 0.7652 0.9842 CLIP Pixl2R
-Montezuma 0.8044 0.9259 0.8045 0.9657 XCLIP ELLM-
-A.4.4 Hyperparameters for Instruction-Following RL Agents
-In the experiments, all methods are implemented based on PPO with same model architecture.
-The Minigrid and Crafter environments use the same training hyperparameters as the Achievement Distillation paper (Moon et al., 2023).
-For Montezuma’s Revenge, we found that the performance of the agent was sensitive to the gamma and GAE lambda parameters.
-To improve the performance of agents in Montezuma’s Revenge, we took two additional steps: (1) normalizing the observation inputs when computing the rewards, and (2) not normalizing the advantage during the GAE calculation.
-The hyperparameters are shown in the following tables.
-Table 4: Model Parameters
-Parameter Value
-model_cls “PPORNNModel”
-hidsize 1024
-gru_layers 1
-impala_kwargs
-- chans [64, 128, 128]
-- outsize 256
-- nblock 2
-- post_pool_groups 1
-- init_norm_kwargs
-- batch_norm false
-- group_norm_groups 1
-dense_init_norm_kwargs
-- layer_norm true
-Table 5: Crafter and Minigrid RL Parameters
-Parameter Value
-gamma 0.95
-gae_lambda 0.65
-algorithm_cls “PPOAlgorithm”
-algorithm_kwargs
-- ppo_nepoch 3
-- ppo_nbatch 8
-- clip_param 0.2
-- vf_loss_coef 0.5
-- ent_coef 0.01
-- lr 3.0e-4
-- max_grad_norm 0.5
-- aux_freq 8
-- aux_nepoch 6
-- pi_dist_coef 1.0
-- vf_dist_coef 1.0
-Table 6: Montezuma RL Training Parameters
-Parameter Value
-gamma 0.99
-gae_lambda 0.95
-int_rew_type “rnd”
-pre_obs_norm_steps 50
-algorithm_cls “PPOAlgorithm”
-algorithm_kwargs
-- update_proportion 0.25
-- ppo_nepoch 3
-- ppo_batch_size 256
-- clip_param 0.1
-- vf_loss_coef 0.5
-- ent_coef 0.001
-- lr 1.0e-4
-A.5 Additional Details of the Experiments on the Impact of Noisy Rewards
-Evaluation Metric Details In our experiments, we used a score metric adapted from the Crafter paper to evaluate agent performance across different environments.
+A.1 Metrics
+In our experiments, we used a score metric adapted from the Crafter benchmark (Hafner 2022) to evaluate agent performance across different environments.
 This score metric aggregates success rates for individual subtasks using a geometric mean.
@@
 For Minigrid and Montezuma environments, we use the internal information from the game engine to detect whether the subtasks are completed, thus facilitating the calculation of the score metric.
-A.5.1 Details of Manipulated Trajectory-Instruction Pairs to Evaluate Robustness
-We evaluated the models’ sensitivity by examining how cosine similarity scores change for manipulated trajectory-instruction pairs.
-These manipulations were designed to test the robustness of the models against various types of noise.
-Here’s a detailed breakdown of the manipulations:
+To evaluate the learning speed of our RL algorithms, we utilize the Area Under the Learning Curve (AUC) metric.
+This metric, previously employed in the literature (Goyal, Niekum, and Mooney 2019, 2020), quantifies the cumulative performance of the agent throughout the training process, effectively indicating how rapidly the agent improves its policy.
+We implement a win cap, limiting the maximum number of wins an agent can achieve during training to a certain number (e.g., 5000).
+Training automatically terminates once this limit is reached, under the assumption that policy learning has stabilized by this point.
+The AUC is therefore number of wins normalized by the total number of training episodes.
+Formally, this is expressed as:
+$\text{AUC}=\frac{\sum_{i=0}^{T}\text{wins}_{i}}{T}$
+where $\text{wins}_{i}$ is the number of wins at episode $i$, and $T$ is the total number of training episodes.
+Appendix B Detailed Related Work
+The approach of converting natural language instructions into reward signals for RL agents has been a longstanding area of exploration.
+Early studies (Kaplan et al.
+2017; Goyal, Niekum, and Mooney 2019; Wang et al.
+2019; Goyal, Niekum, and Mooney 2020) trained separate text and vision encoders from scratch so as to convert multimodal data into continuous feature vectors, then used a discriminator to assess alignment between trajectories and instructions.
+The cosine similarity metric, found by Frome et al.
+(2013) as an effective measure of semantic alignment, became foundational to these methods.
+Strikingly, the core architecture of VLM-based reward models, as illustrated in Figure 1, has remained largely unchanged throughout the years.
+Both pioneering efforts (Kaplan et al.
+2017; Goyal, Niekum, and Mooney 2019; Wang et al.
+2019; Goyal, Niekum, and Mooney 2020) and modern implementations continue relying on computing cosine similarity scores between semantic embeddings of instructions and agent trajectories to generate rewards.
+Formally, given two embedding vectors, $E(l)$ and $E(\tau)$, which denote the semantic embeddings of an instruction $l$ and a trajectory $\tau$, respectively, the cosine similarity is defined as:
+$\text{Cosine Similarity}(E(l),E(\tau))=\frac{E(l)\cdot E(\tau)}{\|E(l)\|\|E(\tau)\|}$ (6)
+However, before the advent of CLIP (Radford et al.
+2021) in 2021, these efforts were hampered by the absence of powerful pre-trained encoder-based VLMs.
+Consequently, researchers focused on proof-of-concept studies, testing their approaches on simpler, short-horizon tasks or sub-tasks — for example, Goyal, Niekum, and Mooney (2019) evaluated their method on sub-tasks rather than the full Montezuma’s Revenge game333available at https://www.retrogames.cz/play˙124-Atari2600.php.
+This likely reflected an understanding that complex, long-horizon tasks were impractical at the time.
+As such, the field had yet to grapple with noisy reward signals, which remained minimal in these simpler environments.
+CLIP (Radford et al.
+2021), an encoder-based VLM, marked a turning point by providing a high-quality joint semantic embeddings for vision and language data.
+Leveraging CLIP, researchers started addressing more complex environments.
+For example, Mahmoudieh, Pathak, and Darrell (2022) used a CLIP-based reward model to guide a robotic arm in manipulation tasks, demonstrating its feasibility for real-world applications.
+In addition to the topics mentioned in the main body of the paper, we continue to highlight related work on reward signals and mitigation strategies below.
+Reward Signal from Human Preference.
+Recent work on RL from Human Feedback (RLHF) (Ouyang et al.
+2022) also leverage expert preference as a reward signal.
+However, our work differs in key aspects.
+Unlike RLHF’s focus on textual outputs, our approach involves evaluating cross-modal similarities between visual and textual data in environments requiring long-horizon decision-making and frequent embodied interactions, a domain not typically covered by RLHF.
+Mitigating Misspecified Rewards.
+Prior works proposed mitigating false positive rewards by training a parallel exploration policy to escape local optima caused by misspecified rewards (Ghosal et al.
+2022; Fu et al.
+2024).
+In contrast, we propose a novel reward function that directly penalizes likely false positive reward signals during training.
+We further show that our method complements exploration-based methods and achieves superior performance when combined.
+Research Gap Successes with CLIP-based reward models have been particularly pronounced in environments with short-horizon tasks.
+For instance, Rocamonde et al.
+(2024) achieved promising results in classic control tasks such as CartPole, MountainCar (Brockman et al.
+2016), and Humanoid (Todorov, Erez, and Tassa 2012), environments that lack long-horizon dependencies.
+Similarly, Wang et al.
+(2024) used the SoftGym environment (Lin et al.
+2020), which includes a suite of short-horizon robotic arm manipulation tasks, as well as MetaWorld environment (Yu et al.
+2019), which lacks strict ordering constraints on subgoals, eventually leading to a lack of long-horizon dependencies.
+Importantly, VLM reward models are learned-based reward models and, as such, inherently suffer from noisy reward signals.
+Sadly, the tendency to focus on simpler environments, or to sidestep the challenges of noisy reward signals, keeps appear in recent work.
+For example, Chan et al.
+(2023) explored CLIP-based rewards to reduce dependence on human-engineered reward functions, but their experiments were confined to simple housekeeping tasks with small action and state spaces — far less complex than pixel-based environments like Montezuma’s Revenge.
+Similarly, ELLM (Du et al.
+2023) relied on a hard-coded oracle reward model (labeled as a “VLM reward model” for conceptual demonstration) to maintain training stability.
+Notably, without this oracle, their agent performed worse than a pure RL agent without any reward shaping — highlighting the limitations imposed by noisy VLM-based rewards in complex environments.
+Despite the growing interest of using VLM-based reward models in embodied RL, the literature has largely overlooked the detrimental effects of noisy reward signals on learning efficiency.
+Key questions remain underexplored: how reliable are these models in complex, long-horizon environments, how do false positive and false negative reward noise affect learning speed, and how to maintain effective learning under noisy reward conditions.
+These gaps highlight a critical area for progress.
+In this work, we aim to address this gap by examining these issues in depth and proposing solutions to enhance the applicability of language-guided RL in real-world scenarios.
+Appendix C The Complete Procedure of the VLM-based Reward Model
+Language instructions in real-world tasks are rarely singular; they typically form a sequence that guides an agent step-by-step.
+A VLM-based reward model must therefore include a mechanism to transition between instructions as the agent progresses.
+Existing VLM-reward implementations applied a pointer mechanism to decide which instruction the agent should consider at the current step.
+We follow this implementation and denote the pointer as $m(t)$.
+It indicates the instruction that the agent is trying to complete at time step $t$.
+$m(t)$ is updated according to the following rule:
+$m(t+1)=\begin{cases}1&\text{if $t=0$}\\ m(t)+1&\text{if instr. $l_{m(t)}$ completed at $t$}\\ m(t)&\text{otherwise}\end{cases}$ (7)
+The pointer remains on the current instruction until the accumulated reward for completing $l_{m(t)}$ reaches a predetermined threshold.
+Given a sequence of instruction sentences $L=\{l_{1},l_{2},\dots,l_{n}\}$, a typical VLM-based reward model, as seen in Pixl2R (Goyal, Niekum, and Mooney 2020) and ELLM (Du et al.
+2023), maintains this pointer to the current instruction, starting with $l_{1}$.
+To track progress, it imposes a maximum reward cap — for example, 2.0 — on each instruction.
+Once the cumulative reward reaches this cap, the pointer advances to the next instruction in the sequence.
+This approach ensures the agent is incentivized to complete one subtask before moving on to the next.
+The complete procedure for training an RL agent with a VLM-based reward model is detailed in algorithm 1.
+This algorithm builds on concepts including MDP formulation, RL algorithms, and the VLM-based reward model (Appendix B).
+Algorithm 1 Instruction-following RL training with VLM-based reward model
+1: Initialize policy network $\pi_{\theta}$
+2: Initialize value network $Q_{\phi}$
+3: Setup VLM-based reward model $E(\cdot)$
+4: Split instruction essay into sentences $\{l_{1},l_{2},...,l_{K}\}$
+5: Initialize instruction pointer $p=1$
+6: Initialize cumulative VLM reward $r_{\text{cum}}=0$
+7: Initialize cumulative VLM reward threshold $q$
+8: Initialize replay buffer $\mathcal{D}$
+9: Initialize agent trajectory memory queue $\tau$ with length $W$
+10: for each episode do
+11:   Initialize state $s_{0}$
+12:   for $t=0$ to T1T-1 do
+13:    Select action $a_{t}\sim\pi_{\theta}(a_{t}|s_{t})$
+14:    Execute $a_{t}$, observe next state $s_{t+1}$ and environmental reward $r^{e}_{t}$
+15:    Enqueue $(s_{t},a_{t},r_{t},s_{t+1})$ in $\tau$
+16:    $\triangleright$  Compute VLM reward   $\triangleleft$:
+17:    $r^{v}_{t}=p(l_{p}|\tau)=\frac{E(\tau)\cdot E(l_{p})}{\|E(\tau)\|\|E(l_{p})\|}$
+18:    Combine rewards: $r_{t}=r^{e}_{t}+(1-\beta)\gamma r^{v}_{t}$ $\triangleright$ $\beta$ is a scaling factor
+19:    Store $(s_{t},a_{t},r_{t},s_{t+1})$ in $\mathcal{D}$
+20:    $r_{\text{cum}}\leftarrow r_{\text{cum}}+r^{v}_{t}$
+21:    if $r_{\text{cum}}\geq q$ then
+22:      $p\leftarrow\min(p+1,K)$
+23:      $r_{\text{cum}}\leftarrow 0$
+24:    if Reach Update Frequency then
+25:      Sample mini-batch $\{(s_{j},a_{j},r_{j},s_{j+1})\}$ from $\mathcal{D}$
+26:      $\triangleright$  Compute TD errors   $\triangleleft$:
+27:      $\delta_{j}=r_{j}+\gamma Q_{\phi}(s_{j+1},a_{j+1})-Q_{\phi}(s_{j},a_{j})$
+28:      $\triangleright$  Update value network   $\triangleleft$:
+29:      $\phi\leftarrow\phi+\alpha_{v}\sum_{j}\delta_{j}\nabla_{\phi}Q_{\phi}(s_{j},a_{j})$
+30:      $\triangleright$  Update policy network   $\triangleleft$:
+31:      $\theta\leftarrow\theta+\alpha_{p}\sum_{j}Q(s_{j},a_{j})\nabla_{\theta}\log\pi_{\theta}(a_{j}|s_{j})$
+Appendix D Convergence time on a sparse-reward landscape
+In this work, Section 4, we have shown that the convergence time of a policy learning algorithm on a sparse-reward landscape can be characterized by the distance in the parameter space.
+Here, we provide a detailed proof of the convergence time on a sparse-reward landscape.
+We can pin down a characteristic convergence time on a sparse-reward landscape.
+The sparse-reward setting enforces that
+$r^{e}(s_{t})=\begin{cases}1&\text{if $s_{t}\in\mathcal{S}_{G}$}\\ 0&\text{otherwise}\end{cases}.$ (8)
+It is clear from the definition that a good trajectory must always be a goal trajectory, and the cumulative reward is simply the reward at the final goal state so $G^{e}(\tau)=\gamma^{T}$ in this setting.
+For a randomly initialized policy, it is highly unlikely that the initial distribution of trajectories contains any goal trajectory due to the sparsity of goal states.
+The optimization of $V^{e}_{\pi_{\theta}}$ thus consists of two parts.
+The first part is to search for a goal trajectory.
+The gradient landscape is almost 0 everywhere, except for cases where a trajectory is $\delta$-close to a goal trajectory.
+Here $\delta$ is the differential unit in the numerical differentiation used in the gradient calculation
+$\theta=\theta+\alpha\nabla_{\theta}V_{\pi_{\theta}}$ (9)
+such that $\delta$-close means being numerically accessible within a distance of $|\delta|$ in parameter space.
+And the second part is to reduce $T$ so that goal trajectories become good trajectories and consequently achieving acceptable policies.
+For the first part, searching for a trajectory for the target is effectively a random walk in the $d$-dimensional parameter space due to the flat gradient landscape.
+Lemma D.1.
+For a random walk in $n$-dimensional space, the expected number of steps $T_{D}$ needed to travel a distance of $D$ scales with $D^{2}$.
+Proof.
+Let $\vec{X_{1}}$, $\vec{X_{2}}$, … $\vec{X_{T}}$ be IID random unit vectors uniformly distributed on a (d1)(d-1)-dimensional sphere $\mathcal{S}^{d-1}\subset\mathbb{R}^{d}$, where $\vec{X_{i}}=(X_{i1},...,X_{id})$ and $|\vec{X_{i}}|^{2}=\sum_{j=1}^{d}X^{2}_{ij}=1$.
+Let $\vec{S_{T}}:=\sum_{i=1}^{T}\vec{X_{i}}$.
+By a $n$-dimensional cosine rule we have
+$|\vec{S}_{T}|^{2}$ $=|\vec{S}_{T-1}|^{2}+2\vec{S}_{T-1}\cdot\vec{X}_{T}+|\vec{X}_{T}|^{2},$ (10)
+and because $\mathbb{E}[\vec{X}_{T}]=\vec{0}$
+$\mathbb{E}[|\vec{S}_{T}|^{2}]$ $=\mathbb{E}[|\vec{S}_{T-1}|^{2}]+\mathbb{E}[2\vec{S}_{T-1}\cdot\vec{X}_{T}]+1$ (11)
+$=\mathbb{E}[|\vec{S}_{T-1}|^{2})]+2\vec{S}_{T-1}\mathbb{E}[\vec{X}_{T}]+1$ (12)
+$=\mathbb{E}[|\vec{S}_{T-1}|^{2}]+1$ (13)
+$=\mathbb{E}[|\vec{S}_{T-2}|^{2}]+1+1$ (14)
+$\vdots$ (15)
+$=T$ (16)
+[…]
+See 4.1 Auxiliary rewards essentially open the path for a divide-and-conquer approach by introducing intermediate rewards in the learning process. We introduce BiMI rewards as
+$r^{v}_{\textsc{BiMI}}(\tau,l_{k})=\max(\mathbf{1}_{\{p(l_{k}\mid\tau)\geq\hat{q}\}}-p(l_{k}),0)$ (17)
+and the cumulative reward of a single trajectory in the presence of BiMI rewards becomes
+$G^{v}_{\textsc{BiMI}}(\tau)=\sum_{t=0}^{T}\gamma^{t}r^{v}_{\textsc{BiMI}}(\tau,l_{m(t)}).$ (18)
+[…]
+See 4.2
+Proof.
+The expected time taken for each of the sub-tasks then scales with $d_{i}^{2}$ respectively (Lemma D.1), and we have
+$d_{1}^{2}+d_{2}^{2}+...+d_{n-1}^{2}$ $<(d_{1}+d_{2}+...+d_{n-1})^{2}$ (19)
+$\mathbb{E}\left[\sum_{i=1}^{n-1}T_{d_{i}}\right]$ $<\mathbb{E}[T_{D}]$ (20)
+because $d_{i}>0\forall i$.
+We can also work out that the upper bound for this improvement is a factor of n1n-1 by invoking the Cauchy–Schwarz inequality $\left(\sum_{i=1}^{n}u_{i}^{2}\right)\left(\sum_{i=1}^{n}v_{i}^{2}\right)\geq\left(\sum_{i=1}^{n}u_{i}v_{i}\right)^{2}$:
+$(d_{1}^{2}+d_{2}^{2}+...+d_{n-1}^{2})(1^{2}+1^{2}+...+1^{2})$ $\geq(d_{1}+d_{2}+...+d_{n-1})^{2}$ (21)
+$(d_{1}^{2}+d_{2}^{2}+...+d_{n-1}^{2})$ $\geq\frac{1}{n-1}D^{2}$ (22)
+$\mathbb{E}\left[\sum_{i=1}^{n-1}T_{d_{i}}\right]\geq\frac{1}{n-1}\mathbb{E}[T_{D}]$ (23)
+and the equality sign holds (indicating maximal improvement) when $d_{1}=d_{2}=...=d_{n-1}$.
+The intuitive interpretation is that the divide-and-conquer approach is the most effective when the task is divided evenly into subtasks.
+∎
+Appendix E VLM-based reward model is a heuristic function
+See 4.4
+Proof.
+Consider a sparse reward environment where the external reward is defined as:
+$r^{e}(s_{t})=\begin{cases}1&\text{if }s_{t}\in\mathcal{S}_{G},\\ 0&\text{otherwise},\end{cases}$
+with discount factor $\gamma\in(0,1)$. For an optimal policy $\pi^{*}$, the value function at state $s_{t}$ is:
+$V^{*}(s_{t})=\gamma^{\widetilde{T}-t},$
+where $\widetilde{T}$ is the time to reach a goal state $s_{g}\in\mathcal{S}_{G}$ along the shortest path from $s_{0}$, and $\widetilde{T}-t$ is the remaining steps from $s_{t}$.
+This follows because the agent receives a single reward of 1 at $s_{g}$, discounted back to $s_{t}$.
+In Algorithm 1, the reward $r^{v}(\tau_{t},l_{m(t)})\in[0,1]$ measures alignment between the agent’s transition $(\tau_{t-1},s_{t},a_{t})$ and the expert instruction $l_{m(t)}$ for sub-task $m(t)$.
+By design, $r^{v}=1$ when the transition fully matches $l_{m(t)}$, and $r^{v}<1$ otherwise, implying partial alignment.
+We assume expert instructions $\{l_{1},\dots,l_{n}\}$ define a trajectory $\tau^{*}$ that reaches $\mathcal{S}_{G}$ optimally or near-optimally.
+If the agent fully follows $l_{m(t)}$ at each step, its path aligns with $\tau^{*}$, and $\widetilde{T}$ is minimized (say, $\widetilde{T}^{*}$).
+When the agent perfectly follows the instruction at each step along $\tau^{*}$, the VLM reward is maximized, i.e., $r^{v}(\tau_{t},l_{m(t)})=1$.
+Crucially, moving along the optimal trajectory $\tau^{*}$ leads to the optimal value function $V^{*}(s_{t})=\gamma^{\widetilde{T}^{*}-t}$.
+Now, if the agent deviates from $l_{m(t)}$, it takes a suboptimal action, increasing the expected path length to $\mathcal{S}_{G}$.
+Let $\widetilde{T}^{\prime}>\widetilde{T}^{*}$ denote this detour.
+Then, $V^{*}(s_{t})=\gamma^{\widetilde{T}^{\prime}-t}<\gamma^{\widetilde{T}^{*}-t}$.
+Since $r^{v}$ decreases with misalignment (e.g., from 1 to some $0\leq r<1$), it exhibits a monotonic relationship with the value function: higher alignment (larger $r^{v}$) corresponds to movement towards a shorter path and thus a larger $V^{*}(s_{t})$, and vice versa.
+Thus, VLM-based rewards, under Algorithm 1, function as a heuristic $h(s_{t})$ that approximates the relative ordering of states in terms of their optimal value $V^{*}(s)$. ∎
+Appendix F False Positive and the Violation of Pessimistic Property of Heuristics
+See 4.9
+Proof.
+We want to show that given:
   1. 1.
-    Trajectory Reversal: We inverted the sequence of frames within each trajectory (i.e., frames = frame[::-1]) to assess the model’s ability to detect reversed state transitions.
-This manipulation tests whether the model can distinguish between forward and backward progression in the state transition.
+    for all successor states $s^{\prime}$: $h(s^{\prime})\leq V^{*}(s^{\prime})$ for an arbitrary $s$.
   2. 2.
-    Instruction Negation: We modified the original instructions by adding negation (e.g., changing “do $l_{k}$” to “do not do $l_{k}$” or “avoid $l_{k}$”).
-This tests the model’s sensitivity to semantic changes in the instruction that fundamentally alter the goal.
+    false positive at current arbitrary state $s$: $V^{*}(s)<h(s)$.
+The objective is to show that under the above conditions, we obtain $\max_{a}(\mathcal{B}h)(s,a)<h(s)$.
+  1. 1.
+    Express the Bellman equation for $h$: $(\mathcal{B}h)(s,a)=R(s,a)+\gamma\sum_{s^{\prime}}P(s^{\prime}|s,a)h(s^{\prime})$
+  2. 2.
+    Express the Bellman version of $V^{*}$: $V^{*}(s)=\max_{a\in\mathcal{A}}\left[R(s,a)+\gamma\sum_{s^{\prime}}P(s^{\prime}|s,a)V^{*}(s^{\prime})\right]$.
+    Given that $h(s^{\prime})\leq V^{*}(s^{\prime})$ for all $s^{\prime}$, we can expand it to $\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)h(s^{\prime})\leq\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)V^{*}(s^{\prime})$.
+Therefore, we have:
+    $R(s,a)+\gamma\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)h(s^{\prime})\leq R(s,a)+\gamma\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)V^{*}(s^{\prime})=Q^{*}(s,a)$ (24)
   3. 3.
-    Instruction Rephrasing: We rephrase the original instructions while maintaining their core meaning.
-This evaluates the model’s robustness to linguistic variations and its ability to capture the essential semantic content of instructions.
+    Taking the maximum over actions to both sides, we have:
+    $\max_{a}\left(R(s,a)+\gamma\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)h(s^{\prime})\right)$ $\leq max_{a}Q^{*}(s,a)$ (25)
+    $\max_{a}(\mathcal{B}h)(s,a)$ $\leq V^{*}(s)$ (26)
   4. 4.
-    Concatenation and Order Swapping: Given two trajectory-instruction pairs $(\tau_{1},l_{1})$ and $(\tau_{2},l_{2})$, we created concatenated pairs and then swapped the order in one modality.
-For example:
-    • Original concatenation: $(\tau_{1}+\tau_{2},l_{1}+l_{2})$
-    • Swapped trajectory: $(\tau_{2}+\tau_{1},l_{1}+l_{2})$
-    • Swapped instruction: $(\tau_{1}+\tau_{2},l_{2}+l_{1})$
-    This tests the model’s sensitivity to the order of components in multi-step tasks.
-  5. 5.
-    Concatenation with Partial Content: We concatenated pairs but truncated one modality. For instance:
-    • Truncated trajectory: $(\tau_{1},l_{1}+l_{2})$
-    • Truncated instruction: $(\tau_{1}+\tau_{2},l_{1})$
-    This assesses the model’s ability to detect partial mismatches in longer sequences.
-A.6 Details of Showing the Prevalence of False Positives in VLM Cosine Similarity Scores
-We list the figures of reward signals from learned VLMs for different types of trajectory-instruction pairs for each individual environment.
+    Apply false positive condition: Given $V^{*}(s)<h(s)$, substituting into the above inequality, we have $\max_{a}(\mathcal{B}h)(s,a)<h(s)$
+Implication: Maintaining a pessimistic heuristic is inherently fragile because the introduction of a false positive in any state disrupts the pessimistic condition. ∎
+Appendix G False Positive gives unbounded bias in the optimality gap
+In this work, Section 4, we analyze the impact of false positive rewards on the convergence speed of policy learning based on the optimality gap analysis by Cheng et al.
+(2021).
+Here, we provide a detailed proof of the impact of false positive rewards on the convergence speed of policy learning.
+We have defined the formal definition of false positive and false negative from both instruction-following (IF) and heuristic perspectives in Section 4.
+See 4.10 We analyze the convergence speed of policy learning by examining the optimality gap, which is defined as the difference between the optimal value of the initial state $s_{0}$, $V^{*}(s_{0})$, and the value of the initial state under an arbitrary policy $\pi$, $V^{\pi}(s_{0})$.
+Specifically, we focus on deriving an upper bound for this optimality gap.
+The key intuition is that a smaller upper bound implies faster convergence to the optimal policy, as fewer iterations of policy updates will be required to reach the optimum.
+We begin by stating the theorem made by HuRL authors:
+Theorem G.1 (Optimality Gap Decomposition (Cheng et al. 2021)).
+For any policy $\pi$, heuristic $h:\mathcal{S}\rightarrow\mathbb{R}$, and mixing coefficient $\beta\in[0,1]$,
+$V^{*}(s_{0})-V^{\pi}(s_{0})=\mathrm{Regret}(h,\beta,\pi)+\mathrm{Bias}(h,\beta,\pi)$ (27)
+where the regret and the bias term are expressed as follows:
+$\mathrm{Regret}(h,\beta,\pi)$ $\coloneqq\beta\left(\widetilde{V}^{*}(s_{0})-\widetilde{V}^{\pi}(s_{0})\right)+\frac{1-\beta}{1-\gamma}\left(\widetilde{V}^{*}(d^{\pi})-\widetilde{V}^{\pi}(d^{\pi})\right)$ (28)
+$\mathrm{Bias}(h,\beta,\pi)$ $\coloneqq\left(V^{*}(s_{0})-\widetilde{V}^{*}(s_{0})\right)+\frac{\gamma(1-\beta)}{1-\gamma}\mathbb{E}_{s,a\sim d^{\pi}}\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}\left[h(s^{\prime})-\widetilde{V}^{*}(s^{\prime})\right]$ (29)
+We will not provide a proof for Theorem G.1 in this paper.
+Please refer to (Cheng et al.
+2021) for details.
+The theorem demonstrates that the performance gap can be elegantly decomposed into two components: a regret term and a bias term such that:
+  1. 1.
+    The regret term quantifies the difference between $\widetilde{V^{*}}$ and $\widetilde{V^{\pi}}$, representing the error caused by $\pi$ being suboptimal in the reshaped MDP $\widetilde{\mathcal{M}}$.
+Since $\pi$ is trained by our selected RL algorithm directly on the reshaped MDP $\widetilde{\mathcal{M}}$, the primary responsibility for minimizing this regret term falls to the RL algorithm itself, not to the design of the auxiliary reward signal.
+Thus, when evaluating the effects of false positive or false negative rewards, we choose not to focus on bounding the regret term.
+  2. 2.
+    The bias term captures two key discrepancies: first, between the true optimal value function $V^{*}$ of the original MDP $\mathcal{M}$ and the optimal value function $\widetilde{V}^{*}$ of the reshaped MDP $\widetilde{\mathcal{M}}$; second, between $\widetilde{V}^{*}$ and the heuristic $h(s)$.
+This term, therefore, reflects the error introduced by addressing the reshaped MDP instead of the original one, alongside how well the heuristic $h$ approximates the optimal value function in the reshaped MDP.
+Consequently, this bias term directly relates to the quality of the heuristic reward signal $h$.
+Hence, we focus on analyzing its upper bound to assess the impact of false positive or false negative rewards on this heuristic.
+Cheng et al. (2021) have further derived the upper bound for the bias term, which we present here as a lemma. We omit the proof for brevity; for detailed proof, please refer to (Cheng et al. 2021).
+Lemma G.2 (Upper Bound of the Bias Term).
+$\mathrm{Bias}(h,\beta,\pi)\leq(1-\beta)\gamma\left(\mathbb{E}_{\rho^{\pi^{*}}}\left[\sum_{t=1}^{\infty}(\beta\gamma)^{t-1}(V^{*}(s_{t})-h(s_{t}))\right]+\mathbb{E}_{\rho^{\pi}}\left[\sum_{t=1}^{\infty}\gamma^{t-1}(h(s_{t})-\widetilde{V}^{*}(s_{t}))\right]\right)$ (30)
+where $\rho^{\pi}$ denotes the trajectory distribution of $s_{0},a_{0},s_{1},$... induced by running $\pi$ starting from $s_{0}$.
+This upper bound elegantly illustrates a trade-off: if the heuristic $h$ is set too high (overestimation error), it reduces the first term but increases the second term.
+Conversely, if $h$ is set too low (underestimation error), it increases the first term but reduces the second term.
+Just by inspection, it is not immediately clear whether overestimation or underestimation results in a better upper bound, as both impact the bias term in opposing ways.
+We now prove that underestimation of $h$ (pessimistic $h$) is better than overestimation (i.e., $\exists s_{t}\in\mathcal{S},h(s_{t})>V^{*}(s_{t})$) for minimizing bias.
+Proof.
+  1. 1.
+    Define Terms:
+    • Let $B_{1}=\mathbb{E}_{\rho^{\pi^{*}}}\left[\sum_{t=1}^{\infty}(\beta\gamma)^{t-1}(V^{*}(s_{t})-h(s_{t}))\right]$
+    • Let $B_{2}=\mathbb{E}_{\rho^{\pi}}\left[\sum_{t=1}^{\infty}\gamma^{t-1}(h(s_{t})-\widetilde{V}^{*}(s_{t}))\right]$
+  2. 2.
+    Underestimation Case (Pessimistic $h$):
+    Lemma G.3.
+    If $h$ is pessimistic with respect to $\mathcal{M}$, $\forall\beta\in[0,1],s\in\mathcal{S},\widetilde{V}^{*}(s)\geq h(s)$
+    Proof.
+    To begin with, we need to use another lemma from (Cheng et al. 2021), shown as follows:
+    Lemma G.4 (Bellman equation of reshaped MDP (Cheng et al. 2021)).
+    For any policy $\pi$, we have
+    $\widetilde{V}^{\pi}(s_{0})-h(s_{0})=\frac{1}{1-\beta\gamma}\mathbb{E}_{\widetilde{d}_{s_{0}}^{\pi}}[(\widetilde{\mathcal{B}}h)(s,a)-h(s)]$ (31)
+    where $\widetilde{d}_{s_{0}}^{\pi}$ refers to the discounted average state distribution of policy $\pi$ in reshaped MDP $\widetilde{\mathcal{M}}$.
+It can be expressed as $\widetilde{d}_{s_{0}}^{\pi}=(1-\widetilde{\gamma})\sum_{t=0}^{\infty}\widetilde{\gamma}^{t}d_{t}^{\pi}$, and $d_{t}^{\pi}$ is the state distribution of policy $\pi$ at time $t$ with $d_{0}^{\pi}=\mathbf{1}\{s=s_{0}\}$.
+    For brevity, we do not include the proof for Lemma G.4.
+    First of all, due to the definition of optimal value $V^{*}$, we have
+    $\widetilde{V}^{*}(s_{0})\geq\widetilde{V}^{\pi}(s_{0})$ (32)
+    Then, according to Lemma G.4, we can get
+    $\widetilde{V}^{*}(s_{0})\geq\widetilde{V}^{\pi}(s_{0})=h(s_{0})+\frac{1}{1-\beta\gamma}\mathbb{E}_{\widetilde{d}_{s_{0}}^{\pi}}[(\widetilde{\mathcal{B}}h)(s,a)-h(s)]$ (33)
+    Recall that the use of VLM-based reward model transform the original MDP into a shaped MDP $\widetilde{\mathcal{M}}=\langle\mathcal{S},\mathcal{A},\mathcal{P},s_{0},\tilde{r},\gamma\rangle$, where the reward function becomes $\tilde{r}=r^{e}(s,a)+r^{v}(\tau_{t},l_{m(t)})$.
+Also, it is important to distinguish the Bellman backup equation under the two different MDPs: the original $\mathcal{M}$ and the reshaped $\widetilde{\mathcal{M}}$.
+By definition, $(\mathcal{B}h)(s,a)=r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[h(s^{\prime})]$.
+In contrast, $(\widetilde{\mathcal{B}}h)(s,a)=\widetilde{r}(s,a)+\widetilde{\gamma}\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[h(s^{\prime})]$.
+Nevertheless, the two Bellman backup equations possess a remarkable property – they are essentially equivalent to each other:
+    $(\widetilde{\mathcal{B}}h)(s,a)$ $=\widetilde{r}(s,a)+\widetilde{\gamma}\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[h(s^{\prime})]$
+    $=\left(r(s,a)+(1-\beta)\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[h(s^{\prime})]\right)+\beta\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[h(s^{\prime})]$
+    $=r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim\mathcal{P}(\cdot|s,a)}[h(s^{\prime})]$
+    $=(\mathcal{B}h)(s,a)$ (34)
+    Let $\pi$ denote the greedy policy of $\arg\max_{a}(\mathcal{B}h)(s,a)$ and then trace back to Equation 34, we have $(\widetilde{\mathcal{B}}h)(s,a)=(\mathcal{B}h)(s,a)$, that means
+    $\widetilde{V}^{*}(s_{0})\geq\widetilde{V}^{\pi}(s_{0})$ $=h(s_{0})+\frac{1}{1-\beta\gamma}\mathbb{E}_{\widetilde{d}_{s_{0}}^{\pi}}[(\widetilde{\mathcal{B}}h)(s,a)-h(s)]$ (35)
+    $=h(s_{0})+\frac{1}{1-\beta\gamma}\mathbb{E}_{\widetilde{d}_{s_{0}}^{\pi}}[(\mathcal{B}h)(s,a)-h(s)]$ (36)
+    $\geq h(s_{0})\quad\text{[direct result of $h$ being pessimistic]}$ (37)
+    Generalization to Any State: The lemma from (Cheng et al.
+2021) (Lemma G.4) is stated in terms of any starting state $s_{0}$.
+Therefore, we can replace $s_{0}$ with any state $s\in\mathcal{S}$ in our analysis.
+Thus we get $\forall s\in\mathcal{S},\widetilde{V}^{*}(s)\geq h(s)$
+    Lemma G.3 implies that when $h$ is pessimistic, $B_{2}=\mathbb{E}_{\rho^{\pi}}\left[\sum_{t=1}^{\infty}\gamma^{t-1}(h(s_{t})-\widetilde{V}^{*}(s_{t}))\right]\leq 0$ because $\widetilde{V}^{*}(s)\geq h(s)$ for all $s$.
+Therefore, the bias error is only bounded by $B_{1}$ when $h$ is pessimistic, i.e., $\mathrm{Bias}(h,\beta,\pi)\leq B_{1}$.
+  3. 3.
+    Overestimation Case ($\exists s_{t}\in\mathcal{S},h(s_{t})>V^{*}(s_{t})$):
+    We show briefly that, unlike the underestimation error, the overestimation error does not have a closed-form upper bound expression.
+The difficulty, as pointed out by Cheng et al.
+(2021), originates from the trajectory-dependence on $\mathbb{E}_{\rho^{\pi}}[\cdot]$ as the $l_{\infty}$ approximation error here can be difficult to control in large state spaces.
+It is possible that, upon picking up falsely high $h$ states, $\rho^{\pi}$ gets further distorted away from $\rho^{\pi^{*}}$ and accumulates more falsely high $h$ states.
+In other words, this trajectory dependence makes $B_{2}$ prone to a feedback loop of accumulating overestimation errors and resulting in a much larger upper bound for the bias.
+G.1 Convergence Guarantee of HuRL Compared With Potential-Based Reward Shaping
+Previous work on potential-based reward shaping (PBRS) has established that learning a policy on a reshaped MDP can converge to the optimum of the original MDP, as proved by Ng, Harada, and Russell (1999).
+They proved that the optimal $Q$ value of the reshaped MDP is equivalent to the optimal $Q$ value of the original MDP minus a state-dependent function.
+Consequently, for an optimal policy defined as $\pi^{*}=\arg\max_{a\in\mathcal{A}}Q^{*}(s,a)$, this additional state-dependent value does not alter the policy.
+Thus, the optimal policy of the original MDP $\mathcal{M}$ is also the optimal policy for the reshaped $\widetilde{\mathcal{M}}$.
+However, the theoretical convergence for the reshaped MDP within the framework of Heuristic-Guided Reinforcement Learning (HuRL) has not been proven with the same rigor.
+The authors of HuRL made a trick which involves manipulating the coefficient $\beta$, which scales the heuristic-based auxiliary rewards.
+If $\beta$ increases gradually from 0 to 1 over the training process, the agent effectively transitions from interacting with the modified MDP back to the original MDP.
+This method ensures that, at the end of training, the agent is exactly solving the original problem.
+Moreover, according to the Blackwell optimal property (Chen, Kochenderfer, and Spaan 2018), convergence can occur before $\beta$ reaches 1, thereby allowing the trained policy to maintain optimality in the original MDP under HuRL.
+However, in the original HuRL paper, among the five test environments, the $\beta$ hyperparameter was actually fixed for four of them.
+Surprisingly, HuRL still converged to the optimal policy faster than the original MDP without updating $\beta$.
+This observation poses an enigma within the HuRL framework, as it is unclear why this convergence to the optimum of the original MDP occurs without dynamically adjusting $\beta$.
+Thus, while the practical success of HuRL in these environments is evident, the underlying reasons for such convergence remain to be fully understood.
+This discrepancy between theoretical expectation and empirical results leaves room for future research to explore why HuRL converges under fixed $\beta$.
+Nevertheless, proving the convergence properties goes beyond the scope of our paper, which aims to highlight the prevalence of false positive rewards and their impact.
+An alternative perspective on this issue is to consider the learning process of maximizing auxiliary rewards as akin to performing behavior cloning from experts.
+In this sense, we do not need to focus on convergence within the current MDP but can view it as a form of pretraining.
+Appendix H Additional Experimental Setup Details
+H.1 VLM-based Reward Model Procedure
+We provide a brief overview of the VLM-based reward model procedure.
+For a detailed procedure, please refer to Appendix C.
+The VLM-based reward model will have a pointer to the sequence of the instruction sentence, starting at the first sentence.
+For original models Pixl2R and ELLM-, we follow the setting in their original work where for each instruction sentence (the full instruction essay will be split into multiple sentences and treat each sentence as atomic instruction $l_{k}$), the reward model will have a maximum cap of rewards (2.0) it can assign to the agent in one episode.
+When the cap is reached, the reward model will move its pointer to the next instruction sentence.
+This cap value was selected as a hyperparameter in prior work, intended to serve as a heuristic threshold indicating instruction completion.
+However, we note that this approach does not guarantee semantic completion, as high cumulative similarity scores can still occur despite incomplete or incorrect execution.
+For the BiMI reward model, the reward model will move its pointer to the next instruction sentence when the binary signal is triggered.
+In terms of the compute resources, we mainly use one NVIDIA 3090 GPU (24GB VRAM) for training and running both the VLM-based reward model and the RL policy model.
+H.2 Finetuning VLM-based Reward Models
+In contrast to previous work where they rely on hand-crafted oracle multimodal reward models (e.g., Du et al.
+(2023)), we use actual pretrained VLMs to generate reward signals.
+2 VLM backbone models are used in our experiments: 1) CLIP (Radford et al.
+2021), pretrained by image-text pairs; and (2) X-CLIP (Ma et al.
+2022), pretrained by video-text pairs.
+In particular, Pixl2R uses CLIP because it only uses the single latest frame as input.
+In contrast, ELLM- takes a slice of trajectory (i.e., multiple frames) as input, and thus uses either X-CLIP or CLIP with additional RNN encoder as the reward model.
+Due to the cartoonish and abstract visuals of the testing environments, we further fine-tune the VLMs to adapt to this new visual domain.
+This fine-tuning is based on the contrastive learning method.
+To collect data, we use well-trained expert agents from (Moon et al.
+2023) to generate expert trajectories for the Crafter environments and annotate them with instructions using internal information from the game engine.
+For Minigrid environments, we use classical search-based planning robots to generate expert trajectories and annotate them with the corresponding task instructions.
+For Montezuma’s Revenge, we manually annotate the expert trajectories.
+For Minigrid and Crafter, we collected approximately 80,000 training pairs, each consisting of a natural language instruction and a corresponding visual observation (i.e.
+a short trajectory snippet).
+These annotations were generated using internal game state information with corresponding instruction templates.
+For Montezuma’s Revenge, we manually created around 300 high-quality instruction-observation pairs based on expert demonstrations.
+These training data are of high quality, as we have made every effort to avoid false positive rewards due to poor training data quality.
+However, despite the fine-tuning process, false positive rewards remain unavoidable.
+Figure 11 presents an example of instruction data employed for VLM finetuning in the Montezuma’s Revenge environment.
+Figure 11: Example of expert instruction data for VLM finetuning under in Montezuma’s Revenge environment.
+We determine the threshold $\hat{q}$ using conformal prediction to convert continuous cosine similarity scores into binary labels for evaluating the VLM’s ability to recognize correct instruction-observation pairs.
+This allows us to compute standard classification metrics such as precision, recall, and F1-score—providing a preliminary assessment of VLM performance independent of downstream RL interaction.
+The performance of the fine-tuned VLM-based reward models is shown in Table 4.
+However, this high precision does not guarantee robust agent performance.
+As we demonstrate later, agents guided by these models significantly underperform in out-of-distribution (O.O.D.) testing environments, where false positive rewards become prevalent due to distributional shifts, revealing a critical limitation in generalization.
+Table 4: Performance of fine-tuned VLM reward model on the testing dataset using the 90th percentile empirical quantile as threshold
+Environment Precision Accuracy F1 Score Recall Model
+Crafter 0.9847 0.9466 0.8538 0.9702 CLIP ELLM-
+Crafter 0.9799 0.9028 0.7618 0.9842 CLIP Pixl2R
+Crafter 0.2095 0.2514 0.2868 0.9657 XCLIP ELLM-
+Minigrid 0.7260 0.9200 0.7849 0.9763 CLIP ELLM-
+Minigrid 0.6992 0.9086 0.7592 0.9616 CLIP Pixl2R
+Minigrid 0.1716 0.2310 0.2642 0.9704 XCLIP ELLM-
+Montezuma 0.8838 0.9638 0.8825 0.9478 CLIP ELLM-
+Montezuma 0.8343 0.9108 0.7652 0.9842 CLIP Pixl2R
+Montezuma 0.8044 0.9259 0.8045 0.9657 XCLIP ELLM-
+H.3 Hyperparameters for VLM Reward Model + RL Agents
+In the experiments, all methods are implemented based on PPO with same model architecture.
+The Minigrid and Crafter environments use the same training hyperparameters as the Achievement Distillation paper (Moon et al.
+2023).
+For Montezuma’s Revenge, we found that the performance of the agent was sensitive to the gamma and GAE lambda parameters.
+To improve the performance of agents in Montezuma’s Revenge, we took two additional steps: (1) normalizing the observation inputs when computing the rewards, and (2) not normalizing the advantage during the GAE calculation.
+The hyperparameters are shown in Table 5, Table 6 and Table 7.
+Table 5: Agent Policy Model Parameters
+Parameter Value
+model_cls Recurrent PPO
+hidsize 1024
+gru_layers 1
+impala_kwargs
+   - chans [64, 128, 128]
+   - outsize 256
+   - nblock 2
+   - post_pool_groups 1
+   - init_norm_kwargs
+   - batch_norm false
+   - group_norm_groups 1
+dense_init_norm_kwargs
+   - layer_norm true
+Table 6: Crafter and Minigrid RL Training Parameters
+Parameter Value
+gamma 0.95
+gae_lambda 0.65
+algorithm_cls PPO Algorithm
+algorithm_kwargs
+   - ppo_nepoch 3
+   - ppo_nbatch 8
+   - clip_param 0.2
+   - vf_loss_coef 0.5
+   - ent_coef 0.01
+   - lr 3.0e-4
+   - max_grad_norm 0.5
+   - aux_freq 8
+   - aux_nepoch 6
+   - pi_dist_coef 1.0
+   - vf_dist_coef 1.0
+Table 7: Montezuma’s Revenge RL Training Parameters
+Parameter Value
+gamma 0.99
+gae_lambda 0.95
+int_rew_type “rnd”
+pre_obs_norm_steps 50
+algorithm_cls PPO Algorithm
+algorithm_kwargs
+   - update_proportion 0.25
+   - ppo_nepoch 3
+   - ppo_batch_size 256
+   - clip_param 0.1
+   - vf_loss_coef 0.5
+   - ent_coef 0.001
+   - lr 1.0e-4
+H.4 Evaluation Metrics
+As detailed in Appendix A.1, we adopted the score metric from the Crafter benchmark (Hafner 2022) for performance evaluation, as it effectively measures consistent performance across multiple subtasks in sparse reward environments.
+Unlike the maximum total rewards metric, which does not adequately reflect consistent performance, the score metric offers a more reliable indicator of learning progress.
+Appendix I Details of Showing the Prevalence of False Positives in VLM Cosine Similarity Scores
+In this work, we have shown that the VLM-based reward models assign high rewards to manipulated trajectory-instruction pairs, indicating the prevalence of false positive rewards.
+We list the figures of reward signals from learned VLMs for different types of trajectory-instruction pairs for each individual environment — the Montezuma, Minigrid, and Crafter environments.
 All figures show that the VLM-based reward models assign high rewards to manipulated trajectory-instruction pairs, indicating the prevalence of false positive rewards.
@@
 Figure 14: Cosine similarity scores for match, mismatch and manipulated trajectory-instruction pairs in Crafter.
-A.7 Impact of Reward Model Quality on Policy Learning Efficiency: Supplementary Experiments and Quantitative Analysis in Montezuma
+Appendix J Extra Experiment on Simulated Oracle Reward Model to demonstrate H3 and H4
 In our efforts to assess the impact of false positive rewards from auxiliary reward model without the interference of other factors such as domain shift, poor data quality, and errors from other issues such as the choice of multimodal architectures, we devised a simulated auxiliary reward model (also known as oracle auxiliary reward model) that access to internal state information from the game engine.
@@
 Metrics measured are the Area Under the Curve (AUC) for total reward (where higher is better) and success rate (SR) of reaching the goal state.
-The baseline is denoted as B, with $\star$ marking statistical significance ($p<0.05$).
-{tblr}
-width = 0.95colspec = Q[473]Q[346]Q[106],row1 = p,hline1-2,8 = -,hline4,6 = -dashed Model & AUC SR
-(1) PPO (Schulman et al., 2017) all failed 0%
-(2) PPO+RND (B) (Burda et al., 2018) 0.550$\pm$0.066 100%
-(3) PPO + Sim RM 1 (Perfect) 0.287$\pm$0.048 100%
-(4) PPO+RND + Sim RM 1 (Perfect) 0.608$\pm$0.073 100%
-(5) PPO+RND + Sim RM 2 (False Positive) 0.183$\pm$0.187 $\star$ 73.3%
-(6) PPO+RND + Sim RM 3 (Temp. Insen.) 0.051$\pm$0.116 $\star$ 16.7%
-  • Perfect auxiliary reward model have shown enhanced performance compared to weak RL baselines like PPO.
-While agents trained solely with PPO struggled to play Montezuma, incorporating Perfect auxiliary reward model into PPO (i.e., (3) in Table 7) did find the goal state, with the success rate increasing from 0% to 100%.
-However, we observed that the auxiliary reward model learns more slowly than the intrinsic reward model (see (2) in Table 7) when attempting to reach the goal state.
-This finding was initially surprising, but upon examining the agent movement heatmaps in Figure 15(a) and Figure 15(b), an explanation becomes clear: the PPO+RND model discovers shorter paths to the goal state.
-The heatmap reveals that the PPO+RND agent learns to directly jump to a rope, bypassing the use of a ladder and conveyor belt as suggested by the expert instructions.
-This observation raises a concern about instruction-following based reward signals; they might limit the agent’s exploration, confining it to states favored by the expert.
-If the quality of the expert’s knowledge about the task is not optimal, this can result in the agent failing to compete with an exploration-based reward model.
-  • Again, we observed a remarkable synergy between the instruction-following-based reward model and the intrinsic reward model, as discussed in §7.1.
-As shown in (4) in Table 7, the PPO+RND + Sim RM 1 (Perfect) agent achieves the highest AUC score, demonstrating that it learns to reach the goal state faster than any other model.
-  • The involvement of false positive rewards significantly slows down the learning process of the model, as evidenced by entries (5) and (6) in Table 7.
-Additionally, both (5) and (6) show lower AUC scores compared to (2), which is the vanilla PPO+RND model.
-If we consider the vanilla PPO+RND model to represent a scenario with full false negative instruction-following rewards (since no instruction-following reward model is implemented here), then comparing the scores of (2) with (5) and (6) further indicates that false positive rewards are more detrimental than false negative ones.
-  • The RL + VLM reward approach completely failed to converge when the reward model disregarded temporal ordering, as evidenced by entry (6) in Table 7.
-The success rate had a significant drop to 16.7%, and even the RND intrinsic reward was unable to recover the success rate within the time budget.
-This illustrates the severe impact of false positive rewards due to temporal ordering insensitivity, an example of composition insensitivity discussed in §5; they trap the agent in a cycle of ineffective actions.
-However, note that if every sub-task is fulfilled, the agent will still receive the maximum total rewards, suggesting that, in theory, the policy could eventually converge.
-Yet, the practical implications of such false positives emphasize how they can drastically extend the learning period or lead to non-convergence in real-world scenarios where time or computational resources are limited.
-(a) PPO+RND agent (same as + Sim RM (False Negative))
-(b) PPO+RND + Sim RM 1 (Perfect)
-(c) PPO+RND + Sim RM 2 (False Positive)
-(d) PPO+RND + Sim RM 3 (Temporal Insensitivity)
-Figure 15: Movement heatmap for PPO+RND agents when different simulated auxiliary instruction-following-based reward models are involved.
-We also provided movement heatmap of different models for further qualitative analysis.
-As shown in Figure 15(a), intrinsic reward model explored a shortcut towards the goal state by directly jump from the initial state to the rope, thereby learning to reach the goal state faster than other models.
-Figures 15(b) and 15(c) do not exhibit a significant contrast, but they do reveal that the presence of false positive rewards leads the agent to more frequently visit dead ends, such as falling off cliffs.
-This behavior is consistent with the mechanics of the simulated false positive reward model, which defines a radius $\sigma$ around target states.
-Within this radius, even if an agent falls off a cliff, it might receive rewards because its position could be close enough to the intended intermediate rewarding state.
-Consequently, agents had a chance of mistakenly perceiving falling off as advantageous.
-Figure 15(d) illustrates the consequences when temporal ordering restrictions are removed: the learning agent persistently pursues rewards associated with the final instruction, “walk left to the door.” However, simply executing this last instruction does not suffice for escaping the room because access to the door requires a key.
-This suggests that false positive rewards cause the agent to get trapped in local minima closer to $s_{0}$, indicating that deep RL can easily hack the reward system and effectively exploit shortcuts (Clark, 2016).
-A critical issue arises here: even with intrinsic rewards, the agent fails to escape these local optima, suggesting it has reached an equilibrium in the MDP where policy updates cease.
-This scenario is particularly relevant when the initial state distribution does not cover the entire state space.
-If the initial distribution does not have a positive probability for all states (i.e., $foralls\in\mathcal{S},d_{0}(s)>0$), it is not guaranteed that the RL algorithm will eventually reach the global optimum with maximum total rewards, as highlighted by (Agarwal et al., 2021).
-A.8 Pseudo-code for Empirical Quantile Calculation for Binary Signal Threshold
-Using empirical quantile as threshold guarantees a high probability (at least $1-\alpha$) that the true positive pairs are recognized while minimizing the average number of mistakes predicting false positives (Sadinle et al., 2019):
-Algorithm 2 Calculate Empirical Quantile ($\hat{q}$)
-Calibration set $\{\tau,l\}_{n}$, where $l$is the instruction sentence, $\tau$is the corresponding trajectory, and $n$is the number of samples; Significance level $\alpha$; VLM model reward model $v$Obtain the similarity-based score
-[…]
-2:$q_{\text{level}}\leftarrow\frac{\lceil(n-1)\times(1-\alpha)\rceil}{n}$LCommentCompute the empirical quantile
+The baseline is annotated using a $\star$ symbol.
+The results are averaged over 5 runs, with standard deviations reported.
+Model AUC SR
+(1) PPO (Schulman et al. 2017) all failed 0%
 […]
@@
 […]
-A.9 Implementation Details of the Experiments of BiMI Reward Function
-We set confidence level for empirical quantile calculation to be $0.91-\alpha=$ 0.9.
-We adhered to the standard requirement of limiting the training budget to 1 million frames (Hafner, 2021).
-This constraint poses a significant challenge, particularly in sparse reward settings, as it demands that agents both explore efficiently and exploit their knowledge effectively within this limited budget.
-To achieve the 1 million frame budget, we used the following configuration:
-  • nproc: 8 (Number of processes used for parallel environments)
-  • nstep: 512 (Length of the rollout stored in the buffer)
-  • nepoch: 250 (Number of epochs to train the RL policy)
-This configuration results in approximately 1 million steps: 250 epochs $\times$ 512 steps $\times$ 8 processes $=$ 1,024,000 frames.
-In the case of Montezuma’s Revenge, we found that the 1 million frame limit used in Crafter was insufficient due to the game’s complexity and sparse reward structure.
-To address this, we extended the training budget to 8 million frames.
-It’s important to note that even with this increased frame count, agents were still unable to fully solve the task.
-As Zhang et al.
-(2021) pointed out, about 1 billion frames are required to truly master Montezuma’s Revenge.
-This vast difference in required training time (8 million vs 1 billion frames) underscores the exceptional difficulty of Montezuma’s Revenge as a sparse reward task.
-The implementation details for the BiMI reward function are consistent with those outlined in the first stage of experiments.
-A.10 Detailed Experiment Results of BiMI Reward Function
+[…]
+  1. 1.
+    Perfect auxiliary reward model have shown enhanced performance compared to weak RL baselines like PPO.
+While agents trained solely with PPO struggled to play Montezuma, incorporating Perfect auxiliary reward model into PPO (i.e., entry (3) in Table 8) did find the goal state, with the success rate increasing from 0% to 100%.
+However, we observed that the auxiliary reward model learns more slowly than the intrinsic reward model.
+This is evidenced by its lower AUC score (0.287 for entry (3) vs.
+0.550 for entry (2) in Table 8), where AUC reflects how quickly the agent learn to reach the goal state.
+This finding was initially surprising, but upon examining the agent movement heatmaps in Figure 15(a) and Figure 15(b), an explanation becomes clear: the PPO+RND model discovers shorter paths to the goal state.
+The heatmap reveals that the PPO+RND agent learns to directly jump to a rope, bypassing the use of a ladder and conveyor belt as suggested by the expert instructions.
+This observation highlights a nuance in VLM-based reward signals, potentially overlooking more efficient routes when expert instructions do not serve as landmarks (a well-known concept in heuristic search in the field of classical planning).
+  2. 2.
+    We observed a remarkable synergy between the instruction-based reward model and the intrinsic reward model.
+As shown in entry (4) in Table 8, the PPO+RND + Sim RM 1 (Perfect) agent achieves the highest AUC score, demonstrating that it learns to reach the goal state faster than any other model.
+  3. 3.
+    False positive rewards significantly impeded learning, as shown by entries (5) and (6) in Table 8.
+Heatmaps (Figure 15(b), Figure 15(c)) reveal agents frequently visiting dead ends (e.g., falling off cliffs) under these conditions, suggesting that the agent is trapped in local minima.
+To assess their impact, we compare these to PPO+RND (entry (2)), a baseline lacking VLM-based instruction rewards.
+Since (2) does not even have a VLM-based reward model, we treat it as a baseline with full false negative instruction-based rewards.
+The AUC scores of (5) and (6) are lower than (2), indicating that false positives are more detrimental than false negatives, supporting our hypothesis H4, as detailed shortly.
+  4. 4.
+    When the VLM reward model ignored temporal ordering (entry (6)), convergence failed, with the success rate dropping to 16.7% and RND unable to recover within the time limit.
+Heatmap (Figure 15(d)) illustrates the agent fixating on the final instruction (“walk left to the door”) without securing the key, trapping it in local minima near $s_{0}$ due to false positives — a form of composition insensitivity (Section 5).
+Moreover, given that Montezuma’s Revenge environment has a fixed initial state distribution, (i.e., the agent always starts at a fixed $s_{0}$), it is not guaranteed that the RL algorithm will eventually reach the global optimum with maximum total rewards, as highlighted by Agarwal et al.
+(2019).
+(a) PPO+RND (same as + RM (False Negative))
+(b) PPO+RND + Sim RM 1 (Perfect Simulation)
+(c) PPO+RND + Sim RM 2 (False Positive)
+(d) PPO+RND + Sim RM 3 (Temporal Insensitivity)
+Figure 15: Movement heatmap for PPO+RND agents when different simulated auxiliary instruction-following-based reward models are involved.
+Figure 15 provides initial evidence of the differential impact of false negatives versus false positives on training outcomes, as posited by H4, suggesting that false positives — particularly those tied to temporal insensitivity — more severely degrade final scores.
+Appendix K Algorithm for Empirical Quantile Calculation
+The empirical quantile $\hat{q}$ is calculated using the conformal prediction method, shown as follows:
+Algorithm 2 Calculate Empirical Quantile ($\hat{q}$) in VLM RM
+1: Calibration set $\{\tau,l\}_{n}$, where $l$ is the instruction sentence, $\tau$ is the corresponding trajectory, and $n$ is the number of samples; User-defined error rate $\alpha$; VLM model reward model $v$
+[…]
+3: $r\}_{n}\leftarrow\{v(\tau,l)\}_{n}$
+[…]
+5: $q_{\text{level}}\leftarrow\frac{\lceil(n-1)\times(1-\alpha)\rceil}{n}$
+[…]
+7: $\hat{q}\leftarrow\texttt{np.quantile($\{r\}_{n}$, $q_{\text{level}}$, method=`lower')}$
+8: return $\hat{q}$
+Appendix L Ablation Study Lineplots for BiMI reward
 Figure 16: Besides the improvements of the score performance of agents across different environments with the BiMI reward function, it also collaborates well with intrinsic rewards.
@@
 Excluding Crafter, Mutual Information (MI) provided a 23% further improvement over Bi alone
-A.10.1 Minigrid
-ELLM-$+$BiMI achieved a remarkable 74% improvement in performance compared to the original models.This substantial gain is particularly noteworthy given the unique challenges presented by Minigrid.
-The abstract, shape-based visuals in Minigrid diverge drastically from the natural images used in VLMs’ pretraining, preventing the models from effectively utilizing their prior pretraining knowledge.
-Consequently, VLMs struggled to accurately assess similarities between Minigrid’s abstract visuals and textual instructions, resulting in highly noisy reward signals.
-The significant improvement demonstrated by BiMI underscores its effectiveness in handling noisy signals, directly addressing our primary research challenge.
-This capability is crucial for deploying instruction-following agents in real-world, unfamiliar scenarios, where visual inputs often deviate from the VLMs’ training distribution, leading to noisy reward signals.
-A.10.2 Crafter
-We observed an intriguing pattern of results.
-The Bi component alone led to 14% and 3.2% improvement in performance over the original models.
-However, contrary to our observations in other environments, the addition of the MI component actually decreased this improvement.
-This unexpected outcome can be attributed to the unique nature of Crafter task, where agents must repeatedly achieve the same subtasks (e.g., drinking water) for survival.
-The MI component, designed to discourage over-reliance on frequently occurring signals, inadvertently penalized the necessary repetition of survival-critical actions.
-Furthermore, note that instruction-following RL agents, regardless of the reward model employed, were unable to outperform pure RL agents in Crafter.
-This discrepancy is due to the open-world nature of Crafter, which requires dynamic strategies and real-time decision-making that our testing setups did not fully capture.
-Despite these challenges, it is noteworthy that Bi alone still managed to improve performance over vanilla VLM-based reward models, suggesting that reducing false positives is still beneficial across all testing environments.
-The combination of BiMI with DEIR (the intrinsic reward model) also showed promising results, indicating a productive balance between exploration (driven by DEIR) and exploitation (guided by BiMI instruction reward).
-A.11 DEPRECATED Proof of the Reduction of Convergence Rate
-WARNING: The following content is deprecated.
-This section contains the original convergence analysis for the ARR October Version, which has been identified as problematic by Reviewer nSDu.
-It is retained here solely for reference and backtracking purposes.
-Formally, Abel et al. (2021) defined that:
-Theorem A.7 (Abel et al. (2021)).
-A reward function realizes a Range Set of Acceptable Policies (Range-SOAP) $\Pi_{G}$ when there exists an $\epsilon\geq 0$ such that every $\pi_{g}\in\Pi_{G}$ is $\epsilon$-optimal in start-state value, $V^{*}(s_{0})-V^{\pi_{g}}(s_{0})\leq\epsilon$, while all other policies are worse.
-When the reward signal is sparse, the agent only receives a reward upon reaching some goal states, and the reward function does not provide any feedback during the intermediate steps.
-We argue that the sparse reward function realizes a Range-SOAP, leading to the categorization of policies into acceptable and unacceptable policies.
-Proposition A.8.
-Sparse reward function “realizes” Range-SOAP (i.e., Range Set of Acceptable Policies).
-Justification:
-A sparse reward function, where the agent only receives a reward upon completing the task, cannot prefer policies that lead to shorter task completion times.
-This is because either the agent completes the goal very quickly or slowly, they will receive nearly the same amount of cumulative rewards, and the reward function will not show a strong preference.
-Since the sparse reward function does not induce a strict partial ordering on the policies, we say this reward function cannot realize a Partial Ordering (PO) task.
-Specifically, a PO on policies is a generalization of a Set of Acceptable Policies (SOAP) task.
-In a PO, the agent specifies a partial ordering on the policy space, where some policies are identified as “great”, some as “good”, and some as “bad” to strictly avoid, while remaining indifferent to the rest.
-Therefore, the sparse reward function can realize a Set of Acceptable Policies (SOAP), where there is a set of policies that are all considered “good” or near-optimal, while all other policies are worse.
-Furthermore, the sparse reward function will lead to a Range-SOAP, rather than an Equal-SOAP.
-Specifically, Equal-SOAP is a SOAP where all the acceptable policies are equally optimal in start-state value.
-This is because the good policies in the SOAP may differ slightly in their start-state values, as some may reach multiple goal states in the environment and thereby receiving different cumulative rewards.
-Therefore, the sparse reward function will realize a Range-SOAP, where there is a range of acceptable policies that are all near-optimal in start-state value.
-This proposition forms the basis of our theoretical analysis in Section A.11.
-Below is the convergence analysis when the total return can be classified by returns of trajectories from acceptable policies $P(\tau\in\mathcal{T}_{G}\mid\theta)\cdot\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{G}]$ versus returns of trajectories from unacceptable policies $P(\tau\in\mathcal{T}_{B}\mid\theta)\cdot\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{B}]$, where $\mathcal{T}_{G}\sim\Pi_{G}$, $\mathcal{T}_{B}\sim\Pi_{B}$ and $\Pi_{B}=\Pi\setminus\Pi_{G}$.
-Note on Derivation: The proof presented here can be viewed as a reformulation of the Q-value function’s gradient in the context of our specific problem setup.
-While this derivation may appear straightforward to readers well-versed in reinforcement learning, we include it to provide a clear mathematical foundation for our analysis of false positive rewards in the VLM-RL context.
-This formulation helps bridge the gap between standard RL theory and our specific problem domain on instruction-following RL with VLM-based rewards.
-Specifically, the update rule of Actor-Critic algorithm is:
-  • Critic:
-    $\phi\leftarrow\phi-\alpha_{\phi}\nabla_{\phi}(\delta)^{2}$ (41)
-    where
-    $\delta=\mathbb{E}_{\pi_{\theta}}[G^{\pi_{\theta}}-Q_{\phi}(s,a)]$ is the Monte Carlo (MC) estimation error, $Q_{\phi}(s,a)$ is the $Q$-value function which measures the expected discounted cumulative reward given the state $s$ and action $a$, and $\pi_{\theta}$ is the policy.
-    $G^{\pi_{\theta}}$ is the rollout cumulative rewards from the trajectory $\tau^{\pi_{\theta}}$ generated from $\pi_{\theta}$.
-  • Actor:
-    $\theta\leftarrow\theta+\alpha_{\theta}\frac{Q_{\phi}(s,a)\nabla_{\theta}\pi_{\theta}(a|s)}{\pi_{\theta}(a|s)}$ (42)
-We need to make the following assumptions to simplify the theoretical analysis:
-  1. 1.
-    We use $\mathcal{T}_{G}$ to represent the set of acceptable trajectories and use $\mathcal{T}_{B}$ as a non-overlapping set of “not acceptable” trajectories, we can say:
-    $\mathcal{T}_{G}\cap\mathcal{T}_{B}=\emptyset$ (43)
-    where $\mathcal{T}$ is the trajectory distribution space of the agent policy $\pi_{\theta}$.
-  2. 2.
-    When we define $\mathcal{T}_{B}$ and $\mathcal{T}_{G}$ within the universe $\mathcal{T}$, we must account for the possibility of a remaining set to ensure a complete partition of $\mathcal{T}$.
-The remaining set $\mathcal{T}\setminus(\mathcal{T}_{B}\cup\mathcal{T}_{G})\}$ can be interpreted as a trajectories that partially fulfill the instruction.
-Given our focus on the influence of false positive rewards within challenging sequential decision-making tasks, we assume that partially correct trajectories are also unacceptable.
-This stance is justified by the fact that partially correct trajectories can be viewed as false positives: they appear to be plausible behaviors but ultimately fail to fulfill the instructions.
-Rewarding such trajectories could reinforce incomplete or potentially dangerous behaviors.
-The errors discussed in Section 5 can be seen as instances of partially correct trajectories that are ultimately unacceptable.
-For example, incompletely following a safety-critical task such as “turn off the machinery before performing maintenance” could pose significant risks, and rewarding such behavior would be undesirable.
-Consequently, for the purposes of our analysis, we can assume:
-    $\mathcal{T}_{G}\cup\mathcal{T}_{B}=\mathcal{T}$ (44)
-  3. 3.
-    Assume the policy class parameterized by $\theta$ should be expressive enough to capture optimal or near-optimal policies, and the policy is initialized randomly from uniform distribution, i.e., $\pi_{\theta_{0}}\sim\mathcal{U}(\Pi)$.
-Meanwhile, the Q-value is initialized as $Q_{\phi_{0}}(s,a)=0,\forall_{s\in S}\forall_{a\in A}$.
-  4. 4.
-    We assume that $|\mathcal{T}_{G}|$ and $|\mathcal{T}_{B}|$ is a fixed number predefined by the task environment while $\mathbb{E}_{\tau\in\pi_{\theta}}[G(\tau)\mid\tau\in\mathcal{T}_{B}]$ is non-zero as false positive rewards are unavoidable in real-world VLMs.
-Since the update rule for $Q$-value is a gradient descent on $\|\mathbb{E}_{\tau\in\pi_{\theta}}[G(\tau)-Q_{\phi}(s,a)]\|^{2}$, the updated $Q$-value will approach as follows:
-$Q_{\phi}(s,a)\rightarrow\mathbb{E}[G(\tau)]$ (45)
-$=$ $P(\tau\in\mathcal{T}_{B}\mid\theta)\cdot\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{B}]+P(\tau\in\mathcal{T}_{G}\mid\theta)\cdot\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{G}]$
-$+P(\tau\in\mathcal{T}\setminus(\mathcal{T}_{B}\cup\mathcal{T}_{G})\mid\theta)\cdot\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}\setminus(\mathcal{T}_{B}\cup\mathcal{T}_{G})]$ (46)
-[…]
-Given that the update rule for policy $\pi$ is the gradient ascent on $Q_{\phi}(s,a)\pi_{\theta}(a\mid s)$, we have the following:
-$\nabla_{\theta}Q_{\phi}(s,a)\pi_{\theta}(a\mid s)$ (49)
-$=$ $\big{(}\nabla_{\theta}P(\tau\in\mathcal{T}_{B}\mid\theta_{\text{old}})\cdot\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{B}]\pi_{\theta}(a\mid s)\big{)}$
-$+\big{(}\nabla_{\theta}P(\tau\in\mathcal{T}_{G}\mid\theta_{\text{old}})\cdot\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{G}]\pi_{\theta}(a\mid s)\big{)}$ (50)
-$=$ $P(\tau\in\mathcal{T}_{B}\mid\theta_{\text{old}})\cdot\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{B}]\nabla_{\theta}\pi_{\theta}(a\mid s)$
-$+P(\tau\in\mathcal{T}_{G}\mid\theta_{\text{old}})\cdot\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{G}]\nabla_{\theta}\pi_{\theta}(a\mid s)\quad\text{$\theta_{\text{old}}$ terms are constants w.r.t.
-$\theta$}$ (51)
-$=$ $\big{(}1-P(\tau\in\mathcal{T}_{G}\mid\theta_{\text{old}})\big{)}\cdot\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{B}]\nabla_{\theta}\pi_{\theta}(a\mid s)$
-$+P(\tau\in\mathcal{T}_{G}\mid\theta_{\text{old}})\cdot\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{G}]\nabla_{\theta}\pi_{\theta}(a\mid s)$ (52)
-$=$ $\textit{const}\cdot\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{B}]\nabla_{\theta}\pi_{\theta}(a\mid s)$
-[…]
-Note the highlighted term in green ($P(\tau\in\mathcal{T}_{G}\mid\theta_{\text{old}})$) is the direction of the gradient ascent on parameter $\theta$.
-Justification of the assumptions.
-  • Regarding the non-zero probability of recovering the optimal policy at initialization, it is standard in theoretical analyses to assume a uniform distribution of a random variable at initialization (see Agarwal et al.
-(2021)).
-  • In reference to the realizability condition implied by Assumption 3, the expressiveness of the policy class parameterized by $\theta$ is an underlying assumption for deep learning models, supported by the The Universal Approximation Theorem (Hornik et al., 1989).
-Observation.
-Since the goal of the learning agent is to maximize $P(\tau\in\mathcal{T}_{G}\mid\theta)$ (i.e., to converge the agent policy to the distribution of acceptable trajectories), we can see that the second term provides the target direction with rate $(\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{G}]-\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{B}])$.
-Therefore, the ascent rate will decrease when the cumulative rewards from unacceptable trajectories (i.e., the false positive rewards) gets higher.
-In addition, the first term $\textit{const}\cdot\mathbb{E}[G(\tau)\mid\tau\in\mathcal{T}_{B}]\nabla_{\theta}\pi_{\theta}(a\mid s)$ can be regarded as the deviation term of target direction.
-This formalization follows the intuition that the presence of false positive rewards can slow down the convergence rate of the learning agent.
+Appendix M Encoder-based Vision-Language Models Explanation
+Similar to LLMs, VLMs can be broadly classified into two categories: encoder-based models (or dual-encoder models) and generative-based models (often referred to as Vision LLMs).
+Encoder-based models focus on mapping vision and language data into a shared semantic space where embeddings of corresponding visual and textual content are positioned close to each other.
+CLIP (Radford et al.
+2021) and SigLIP (Tschannen et al.
+2025) exemplify this approach.
+They are often referred to as VLM dual-encoders as such VLMs consists of both vision encoder and text encoder which can be operated separately to encode vision or language data.
+Dual-encoder models are particularly well-suited for cross-modal retrieval tasks, such as finding the most relevant image for a given text query or vice versa.
+This is achieved by computing similarity scores (e.g., cosine similarity) or other distance metrics between the embeddings of the query and the images in the dataset.
+Dual-encoders are therefore being explored to serve as learned-based reward models for embodied RL agents (Rocamonde et al.
+2024; Du et al.
+2023), which we discuss in Appendix B.
+In contrast, generative-based models such as LLaVA (Liu et al.
+2023) and Sa2VA (Yuan et al.
+2025) are typically built by connecting a vision encoder (e.g., CLIP-based vision encoder) to an LLM via a projection layer (e.g., a Multilayer Perceptron (MLP)).
+The backbone LLM takes both vision feature vectors and text tokens as input and generates textual outputs, enabling these models to perform tasks such as Vision Question Answering (VQA).
+Although both categories are commonly classified as VLMs, they serve fundamentally different purposes and exhibit distinct behaviors.
+Appendix N Sparse Reward and Random Walk
+Below, we will discuss the characteristics of RL training under sparse rewards
+Proposition N.1.
+With sparse rewards, the gradient landscape is nearly flat, making gradient-ascent updates indistinguishable from a random walk in parameter space.
+While this may seem obvious, no existing textbook has yet explained this concept clearly.
+Here, we aim to formalize and justify this proposition, as it will be crucial for analyzing the performance of VLM-based reward models in this work.
+Recall the gradient of the policy gradient update in online Actor Critic methods: $\theta\leftarrow\theta+\alpha\mathbb{E}_{\pi_{\theta}}\left[Q_{\phi}(s,a)\nabla_{\theta}\log\pi_{\theta}(a|s)\right]$.
+Using chain rule, the gradient of the policy gradient update can be decomposed as:
+$\mathbb{E}_{\pi_{\theta}}\left[\frac{Q_{\phi}(s,a)}{\pi_{\theta}(a|s)}\nabla_{\theta}\pi_{\theta}(a|s)\right]$
+In sparse reward settings, the Q-function $Q^{\pi_{\theta}}(s,a)$, often approximated by a neural network in actor-critic methods, exhibits significant randomness.
+Initially, the critic’s weights are randomly initialized (e.g., via Xavier initialization (Glorot and Bengio 2010)), producing arbitrary Q-values (Mnih et al.
+2015).
+With most transitions yielding $r=0$, the TD error $\delta=\gamma Q(s^{\prime},a^{\prime})-Q(s,a)$ is typically non-zero due to the stochastic variation between $Q(s,a)$ and $Q(s^{\prime},a^{\prime})$ resulting from random initialization.
+This leads to random updates to the Q-network, as the noisy TD error lacks a consistent direction, further increasing the stochasticity of Q-value estimates.
+Such stochasticity propagates to the policy gradient update, where it even amplifies this effect: At initialization, $\pi_{\theta}(a|s)\approx 1/K$ for a discrete action space with $K$ actions, as random weights yield a near-uniform softmax distribution.
+The gradient term $\frac{\nabla_{\theta}\pi_{\theta}(a|s)}{\pi_{\theta}(a|s)}$ scales $Q_{\phi}(s,a)$ by the inverse of the policy probability.
+Therefore, the update step $\theta\leftarrow\theta+\alpha_{\theta}\frac{Q_{\phi}(s,a)\nabla_{\theta}\pi_{\theta}(a|s)}{\pi_{\theta}(a|s)}$ has a high variance due to noisy Q-value estimates and non-trivial gradient magnitudes from the scaling effect of $\pi_{\theta}(a|s)$.
+Consequently, the update exhibits $\text{Var}(\hat{\nabla}_{\theta}J(\theta))>0$ and $\mathbb{E}[\hat{\nabla}_{\theta}J(\theta)]\approx 0$.
+This behavior renders policy updates indistinguishable from a random walk in parameter space.
+The parameters $\theta$ drift stochastically, as each step’s direction is dominated by noise rather than progress toward a better policy.
+This random walk behavior is a key reason why training under sparse rewards is challenging.